### v3

# ARMAr-LASSO: Mitigating the Impact of Predictor Serial Correlation on the LASSO

Abstract
We explore estimation and forecast accuracy for sparse linear models, focusing on scenarios where both predictors and errors carry serial correlations. We establish a clear link between predictor serial correlation and the performance of the LASSO, showing that even orthogonal or weakly correlated stationary AR processes can lead to significant spurious correlations due to their serial correlations. To address this challenge, we propose a novel approach named ARMAr-LASSO (ARMA residuals LASSO), which applies the LASSO to predictors that have been pre-whitened with ARMA filters and lags of dependent variable. We derive both asymptotic results and oracle inequalities for the ARMAr-LASSO, demonstrating that it effectively reduces estimation errors while also providing an effective forecasting and feature selection strategy. Our findings are supported by extensive simulations and an application to real-world macroeconomic data, which highlight the superior performance of the ARMAr-LASSO for handling sparse linear models in the context of time series.
Keywords: LASSO, Time Series, Serial Correlation, Spurious Correlation.
1 Introduction
The LASSO (tibshirani96) is perhaps the most commonly employed approach to handle regressions with a large number of predictors. From a theoretical standpoint, its effectiveness in terms of estimation, prediction, and feature selection is contingent upon either orthogonality or reasonably weak correlation among predictors (see zhaoyu2006; Bickel2009; Negahban2012; Hastie2015). This hinders the use of the LASSO for the analysis of economic time series data, which are notoriously characterized by intrinsic multicollinearity; that is, by predictor correlations at the population level (Lippi2000; SeW2002a; demol2008; Medeiros2012). A common procedure to address this issue is to model multicollinearity and remove it, as proposed, e.g., by Fan2020, who filter time series using common factors and then apply the LASSO to the filtered residuals. However, mitigating or even eliminating multicollinearity is not the end of the story, as effectiveness of the LASSO can also be affected by spurious correlations. These occur when predictors are orthogonal or weakly correlated at the population level, but a lack of sufficient independent replication (lack of degrees of freedom) introduces correlations at the sample level, potentially leading to false scientific discoveries and incorrect statistical inferences (Fan16). This issue has been broadly explored in ultra-high dimensional settings, where the number of predictors can vastly exceed the available sample size (Fan13). We argue that in time series data, a shortage of independent replication can be due not only to a shortage of available observations but also to serial correlation.
This article introduces two elements of novelty. First, we establish an explicit link between serial correlations and spurious correlations. At a theoretical level, we derive the density of the sample correlation between two orthogonal stationary Gaussian AR(1) processes, and show how such density depends not only on the sample size but also on the degree of serial correlation; an increase in serial correlation results in a larger probability of sizeable spurious correlations. Then we use extensive simulations to show how this dependence holds in much more general settings (e.g., when the underlying processes are not orthogonal, or non-Gaussian ARMA).
Second, we propose an approach that, using a filter similar to that proposed by Fan2020, rescues the performance of the LASSO in the presence of serially correlated predictors. Our approach, which we name ARMAr-LASSO (ARMA residuals LASSO), relies upon a working model where, instead of the observed predictor time series, we use as regressors the residuals of ARMA processes fitted on such series, augmented with lags of the dependent variable. We motivate our choice of working model and provide some asymptotic arguments concerning limiting distribution and feature selection consistency. Next, we employ the mixingale and near-epoch dependence framework (davidson1994; ADAMEK2023) to prove oracle inequalities for the estimation and forecast error bounds of the ARMAr-LASSO, while simultaneously addressing the issue of estimating ARMA residuals. To complete the analysis, we use simulations to validate and generalize theoretical results. Furthermore, we apply our methodology to a high-dimensional dataset for forecasting the consumer price index in the Euro Area. Simulations and empirical exercises demonstrate that the ARMAr-LASSO produces more parsimonious models, better coefficient estimates, and more accurate forecasts than LASSO-based benchmarks. Notably, both theoretical and numerical results concerning our approach hold even in the presence of factor-induced multicollinearity, provided that the idiosyncratic components are orthogonal or weakly correlated processes exhibiting serial correlation.
On the serial correlation front, most of the theoretical econometric literature has focused on its impact in the error terms, particularly regarding post-LASSO inference (see, e.g., Chernozhukov2021; chronopoulos2023; Babii2024; ADAMEK2023). The present study shows that serial correlation in the predictors deserves similar attention, as it can adversely affect both the estimation and forecast accuracy of the LASSO. Our work complements the vast literature on error bounds for LASSO-based methods in time series analysis, which addresses estimation and forecast consistency in scenarios with autocorrelated errors and autoregressive processes (see, e.g., nardi2011; Uematsu2019). Such scenarios are ubiquitous, e.g., they are easily found in US and Euro Area monthly macroeconomic data (see McCracken2016 and Ale2021). Moreover, our methodology is consistent with the existing literature on pre-whitening filters, which aim to mitigate autocorrelation and multicollinearity by applying LASSO or related methods to filtered residuals (see, e.g., Robinson88; Belloni2013; Hansen2019; Fan2020). In particular, the Generalized Least Squares LASSO (GLS-LASSO; chronopoulos2023) and the AutoRegressive Distributed Lag LASSO (ARDL-LASSO; medeiros2017ADA) provide two natural benchmarks for our method. GLS-LASSO improves efficiency by filtering both the dependent variable and the predictors using autoregressive coefficients estimated from residuals of a preliminary LASSO fit, while ARDL-LASSO addresses serial correlation by including lags of both the predictors and the dependent variable. In this paper, we argue that ARMAr-LASSO is preferable to these benchmarks, as it more effectively removes serial correlation in the predictors and thereby enhances both estimation and forecasting performance.
The remainder of the article is organized as follows. Section 2 introduces the problem setup and our results concerning the link between serial correlations and spurious correlations. Section 3 introduces the ARMAr-LASSO and explores its theoretical properties. Section 4 presents simulations and real data analyses to evaluate the proposed methodology. Section 5 provides some final remarks. Appendix A and B contain the proofs of theoretical results and technical details. The Supplement encompasses additional studies and simulations excluded from the main manuscript.
[…]
Code and replicability materials are at https://zenodo.org/records/15089775
2 Problem Setup
Consider the linear regression model
$y_{t}=\mathbf{x}_{t}^{\prime}\boldsymbol{\alpha}^{*}+\varepsilon_{t}\ \ ,\ \ \ \ \ \ t=1,\ldots,T\ \ ,$ (1)
where $\mathbf{x}_{t}=(x_{1,t},\ldots,x_{n,t})^{\prime}$ is a $n\times 1$ vector of predictors, $\boldsymbol{\alpha}^{*}$ is a $n\times 1$ unknown $s$-sparse vector of regression coefficients, i.e. $\boldsymbol{\alpha}^{*}||_{0}=s<n$, and $\varepsilon_{t}$ is an error term. We impose the following assumptions on the processes $\mathopen{}\mathclose{{\left\{\mathbf{x}_{t}}}\right\}$ and $\mathopen{}\mathclose{{\left\{\varepsilon_{t}}}\right\}$.
Assumption 1:
(a) $\mathopen{}\mathclose{{\left\{\mathbf{x}_{t}}}\right\}$ and $\mathopen{}\mathclose{{\left\{\varepsilon_{t}}}\right\}$ are non-deterministic second-order stationary processes of the form
$x_{i,t}$ $=$ $\sum_{l=1}^{p_{i}}\phi_{i,l}x_{i,t-l}+\sum_{k=1}^{q_{i}}\theta_{i,k}u_{i,t-k}+u_{i,t}\ \ ,\ \ \ \ i=1,\dots,n\ \ ,\ \ \ \ p_{i},q_{i}<\infty\ \ ,$ (2)
$\varepsilon_{t}$ $=$ $\sum_{l=1}^{p_{\varepsilon}}\phi_{\varepsilon,l}\varepsilon_{t-l}+\sum_{k=1}^{q_{\varepsilon}}\theta_{\varepsilon,k}\omega_{t-k}+\omega_{t}\ \ ,\ \ \ \ \ \ p_{\varepsilon},q_{\varepsilon}<\infty\ \ .$ (3)
(b) The innovation processes $u_{i,t}\sim i.i.d.(0,\sigma_{i}^{2})$, $\omega_{t}\sim i.i.d.(0,\sigma_{\omega}^{2})$, where $u_{i,t}\perp u_{j,t-l}$ for any $i\neq j,\ t$ and $l\neq 0$; and $u_{i,t-l}\perp\omega_{t}$ for any $i,t$ and $l$.
There are several approaches to estimate a sparse $\boldsymbol{\alpha}^{*}$ (zhang2012; James2013); here we focus on the LASSO estimator (tibshirani96) given by $\widehat{\boldsymbol{\alpha}}=\underset{\boldsymbol{\alpha}\in\mathbb{R}^{n}}{\operatornamewithlimits{argmin\ }}\>\mathopen{}\mathclose{{\left\{\frac{1}{2T}||\mathbf{y}-\mathbf{X}^{\prime}\boldsymbol{\alpha}||_{2}^{2}+\ddot{\lambda}||\boldsymbol{\alpha}||_{1}}}\right\},$ where $\mathbf{y}=(y_{1},\dots,y_{T})^{\prime}$ is the $T\times 1$ response vector, $\mathbf{X}=(\mathbf{x}_{1},\dots,\mathbf{x}_{T})$ is the $n\times T$ design matrix, and $\ddot{\lambda}>0$ is the weight of the $\ell_{1}$ penalty and must be “tuned” to guarantee that regression coefficient estimates are effectively shrunk to zero – thus ensuring predictor, or feature, selection.
However, linear associations among predictors are well known to affect LASSO performance. Bickel2009; Buhlmann2011 and Negahban2012 have shown that the LASSO estimation and prediction accuracy are inversely proportional to the minimum eigenvalue of the predictor sample covariance matrix. Thus, highly correlated predictors deteriorate estimation and prediction performance. Moreover, zhaoyu2006 proved that the LASSO struggles to differentiate between relevant (i.e., $\{i\in\{1,\dots,n\}:\alpha_{i}^{*}\neq 0\}$) and irrelevant (i.e., $\{i\in\{1,\dots,n\}:\alpha_{i}^{*}=0\}$) predictors when they are closely correlated, leading to false positives. Thus, highly correlated predictors may also deteriorate feature selection performance. The irrepresentable condition addresses this issue ensuring both estimation and feature selection consistency through bounds on the sample correlations between relevant and irrelevant predictors (zhaoyu2006, see also Buhlmann2011). Nevertheless, orthogonality or weak correlation seldom hold in the context of economic and financial data. For instance, decades of literature provide evidence for co-movements of macroeconomic variables (Lippi2000; Lippi2005; SeW2002a; SeW2002b). Special methods have been proposed to mitigate the negative effects of these linear associations, such as Factor-Adjusted Regularized Model Selection (FarmSelect) (Fan2020), which applies the LASSO to the idiosyncratic components of economic variables, obtained by filtering the variables through a factor model. Although approaches such as FarmSelect can be very effective in addressing multicollinearity, strong spurious correlations can emerge at the sample level and affect the LASSO even when predictors are orthogonal or weakly correlated at the population level. Sample-level spurious correlations can be particularly prominent in regressions with many predictors, especially if the sample sizes are relatively small, and the problem can be yet more serious for time series data, where independent replication can be further hindered by serial correlations (Bartlett35; Mcgregor1965). This is exactly the focus of this article; in the next section, we introduce a theoretical result linking serial correlations within time series to the sample correlations between them.
2.1 Serial and Sample Correlations for Time Series
Consider a first order $n$-variate autoregressive process $\mathbf{x}_{t}=\boldsymbol{\phi}\mathbf{x}_{t-1}+\mathbf{u}_{t}$, $t=1,\dots,T$, where $\boldsymbol{\phi}$ is the $n\times n$ diagonal matrix with $\text{diag}(\boldsymbol{\phi})=(\phi_{1},\dots,\phi_{n})$, $|\phi_{i}|<1$ for each $i=1,\dots,n$, and $\mathbf{u}_{t}\sim N\mathopen{}\mathclose{{\left(\boldsymbol{0}_{n},\boldsymbol{I}_{n}}}\right)$. Here $\mathbf{x}_{0}\sim$ $N\mathopen{}\mathclose{{\left(\boldsymbol{0}_{n},\mathbf{C}_{x}}}\right)$ and $\mathbf{x}_{t}\sim$ $N\mathopen{}\mathclose{{\left(\boldsymbol{0}_{n},\mathbf{C}_{x}}}\right)$ with $\mathopen{}\mathclose{{\left(\mathbf{C}_{x}}}\right)_{ii}=\frac{1}{1-\phi_{i}^{2}}$, and $\mathopen{}\mathclose{{\left(\mathbf{C}_{x}}}\right)_{ij}={c}_{ij}^{x}=0$, for $i\neq j$. Let $\widehat{\mathbf{C}}_{x}=\frac{1}{T}\mathbf{X}\mathbf{X}^{\prime}$ be the sample covariance, or equivalently, correlation matrix – with generic off-diagonal element $\widehat{c}_{ij}^{x}$ and eigenvalues $\widehat{\psi}_{max}^{x}\geq\ldots\geq\widehat{\psi}_{min}^{x}$. Our next task is to link $\Pr(|\widehat{c}_{ij}^{x}|\geq\tau)$, $\tau\in[0,1)$, to serial correlations. To this end, the following proposition provides an approximation to the probability density of the sample correlation, yielding a formulation that is simpler than that of Mcgregor1965 and builds upon the results of Anderson03 for i.i.d. random variables.
Proposition 1:
Let $\{\mathbf{x}_{t}\}$ be a stationary $n$-variate Gaussian AR(1) process with autoregressive residuals $\mathbf{u}_{t}\sim N\mathopen{}\mathclose{{\left(\boldsymbol{0}_{n},\boldsymbol{I}_{n}}}\right)$. Let $\ddot{\phi}=\phi_{i}\phi_{j}$, where $\phi_{i}$ and $\phi_{j}$ are the autoregressive coefficients of the $i$-th and $j$-th processes, respectively. For some $\nu\in\mathbb{Z}^{+}$, and for all sample sizes $T\geq\mathopen{}\mathclose{{\left\lfloor\nu\mathopen{}\mathclose{{\left(\frac{1+\ddot{\phi}}{1-\ddot{\phi}}}}\right)}}\right\rceil$, the density of $\widehat{c}_{ij}^{x}$ is approximated by
$\mathcal{D}(r)=\frac{\Gamma\mathopen{}\mathclose{{\left(k_{v}+\frac{1}{2}}}\right)(1-\ddot{\phi})\sqrt{\xi_{v}}}{\Gamma\mathopen{}\mathclose{{\left(k_{v}}}\right)\sqrt{\pi}}\frac{\mathopen{}\mathclose{{\left[1-r^{2}}}\right]^{k_{v}-1}\mathopen{}\mathclose{{\left[2T_{v}(1-\ddot{\phi}^{2})}}\right]^{k_{v}}}{\mathopen{}\mathclose{{\left[(1-r^{2})2T_{v}(1-\ddot{\phi}^{2})+r^{2}\xi_{v}(1-\ddot{\phi})^{2}}}\right]^{k_{v}+\frac{1}{2}}}\ \ ,\ \ r\in[-1,1]\ \ ,$
where $T_{v}=\mathopen{}\mathclose{{\left\lfloor\frac{(T-1)(1-\ddot{\phi})^{2}-(1-\ddot{\phi}^{2})}{(1-\ddot{\phi})^{2}}}}\right\rceil$, $\xi_{v}=3T_{v}-T_{v}^{2}+2\sum_{t=1}^{T_{v}-1}(T_{v}-t)(1+2\phi_{j}^{2t})$, and $k_{v}=\frac{T_{v}}{\xi_{v}}$.
Remark 1:
Proposition 1 establishes a lower bound on the sample size at which $\Pr(|\widehat{c}_{ij}^{x}|\geq\tau)\approx\int_{-1}^{-\tau}\mathcal{D}(r)dr+\int_{\tau}^{1}\mathcal{D}(r)dr$. The bound depends on two quantities: $\nu$, which represents the degrees of freedom under independence (henceforth effective degrees of freedom); and the factor $\mathopen{}\mathclose{{\left(\frac{1+\ddot{\phi}}{1-\ddot{\phi}}}}\right)$, which corresponds to the Bartlett correction for AR(1) processes. The dependence on $\ddot{\phi}$ indicates that, due to serial correlation, the effective degrees of freedom—and thus the effective sample size—is smaller than the nominal sample size (see, e.g., Bartlett35). Note that, for any $|\ddot{\phi}|<1$, $\mathcal{D}(r)$ converges to the Normal distribution as $T\to\infty$.
Figure 1 reports the densities of $\widehat{c}_{ij}^{x}$, indicated as $d(r)$, obtained through 5000 Monte Carlo simulations considering $\phi_{i}=\phi_{j}=0.3,0.6,0.9,0.95$ and $T=50,100,250$. For any $T$ value, an increase in $\ddot{\phi}=\phi_{i}\phi_{j}=\phi^{2}$ results in a density with thicker tails, and thus in a higher $\Pr(|\widehat{c}_{ij}^{x}|\geq\tau)$. The results confirm that serial correlation increases the probability of spurious correlations. This, in turn, leads to a higher probability of a small minimum eigenvalue (because $\Pr(\widehat{\psi}_{min}^{x}\leq 1-\tau)\geq\Pr(|\widehat{c}_{ij}^{x}|\geq\tau)$; see Appendix B), and to a higher chance of breaking the irrepresentable condition if, say, one of the processes is relevant for the response and the other is not ($\alpha_{i}^{*}\neq 0$ and $\alpha_{j}^{*}=0$, or vice versa). Note that this happens when $Sign(\phi_{i})=Sign(\phi_{j})$. In contrast, when $Sign(\phi_{i})\neq Sign(\phi_{j})$, an increase in $|\ddot{\phi}|$ results in a density more concentrated around the origin. In Supplement B, we report a detailed analysis of the results in Figure 1. Furthermore, we investigate the impact of $Sign(\ddot{\phi})$, and more scenarios with correlated, non-Gaussian, and/or ARMA processes, through multiple simulation experiments.
Figure 2 compares $d(r)$ (blue histograms and dashed lines) with $\mathcal{D}(r)$ (red line) considering $\nu=20$ and $\phi_{i}=\phi_{j}=0.3,0.6,0.9$. We observe that $d(r)$ is well approximated by $\mathcal{D}(r)$, indicating that Proposition 1 allows us to explicitly link the probability of sizeable spurious correlations to serial correlations. Note that in the proof of Lemma 1 we provide a theoretical justification for considering $\nu=20$.
Figure 1: Monte Carlo densities $d(r)$ of $\widehat{c}_{ij}^{x}$ for different values of $T$ and $\phi$.
Figure 2: Monte Carlo densities $d(r)$ (blue histograms and dashed lines) and $\mathcal{D}(r)$ (red lines) for $\nu=20$ and different values of $\phi$. The $p$-values correspond to the Shapiro test for Gaussianity.
We conclude this Section with a simple “toy experiment”. We generate data for $t=1,\dots,T$ from a $10$-variate process $\mathbf{x}_{t}=\boldsymbol{\phi}\mathbf{x}_{t-1}+\mathbf{u}_{t}$, where all components share the same autoregressive coefficient $\phi_{i}=\phi$, $i=1,\ldots,10$, and $\mathbf{u}_{t}\sim N\mathopen{}\mathclose{{\left(\boldsymbol{0}_{10},\boldsymbol{I}_{10}}}\right)$. Because of orthogonality, for the population correlation matrix $\mathbf{C}_{x}$ we have $\underset{i\neq j}{\text{max}}|c_{ij}^{x}|=0$ and $\psi_{min}^{x}=1$. We consider $\phi=0.0,0.3,0.6,0.9,0.95$, and $T=50,100,250$. For each scenario we calculate the average and standard deviation of $\underset{i\neq j}{\text{max}}|\widehat{c}_{ij}^{x}|$ and $\widehat{\psi}_{min}^{x}$ over 5000 Monte Carlo simulations. Results are shown in Figure 3; a stronger persistence (higher $\phi$) increases the largest spurious sample correlations and decreases the smallest eigenvalue. However, as expected, an increase in the sample size from $T=50$ (panel (a)) to $T=250$ (panel (c)), reduces the impact of $\phi$. For example, the values of $\underset{i\neq j}{\text{max}}|\widehat{c}_{ij}^{x}|$ and $\widehat{\psi}_{min}^{x}$ in the case of $T=50$ and $\phi=0.3$ are quite similar to those obtained for $T=100$ and $\phi=0.6$, and for $T=250$ and $\phi=0.9$. Note that these results are valid for any orthogonal or weakly correlated predictors, as long as they carry serial correlations. These predictors can be either directly observed variables or, for example, factor model residuals.
Figure 3: Numerical “toy example”. Panel (a) $T=50$, Panel (b) $T=100$, Panel (c) $T=250$. Orange circles/bars and blue triangles/bars represent, respectively, means/standard deviations of $\max_{i\neq j}|\widehat{c}_{ij}^{x}|$ and $\widehat{\psi}_{min}^{x}$, for various values of $\phi$, as obtained from 5000 Monte Carlo simulations.
2.2 LASSO Oracle Inequalities for Orthogonal AR(1) Gaussian Processes
In this section, we establish a connection between LASSO performance and serial correlation, building upon the results derived in Section 2.1. For consistency of exposition, we assume that both the predictors and the error terms are independent Gaussian AR(1) processes.
Assume that each row of $\mathbf{X}$ is standardized to have mean 0 and variance 1. Let $\widehat{\mathbf{C}}_{x}\hskip 2.84544pt\stackrel{{\scriptstyle a.s.}}{{\rightarrow}}\mathbf{C}_{x}$, where $\mathbf{C}_{x}$ is a non-negative definite matrix. In settings where $n$ may be larger than $T$, we usually make the following assumption.
Assumption 2:
[…]
Remark 2:
Since the compatibility constant $\gamma_{x}^{2}$ is directly linked to the smallest eigenvalue of the sample covariance matrix, an increase in predictors’ sample correlations drives the minimum eigenvalue toward zero (see Figure 3 and Appendix B), thereby weakening the RE condition.
Proposition 2:
[…]
Corollary 1:
As a consequence of Proposition 2 we have that: ($a$) $\frac{1}{T}\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left|\mathbf{X}^{\prime}(\widehat{\boldsymbol{\alpha}}-\boldsymbol{\alpha}^{*})}}\right|}}\right|_{2}^{2}=O_{P}\mathopen{}\mathclose{{\left(s{\sigma_{x\varepsilon}^{2}\frac{log(n)}{T}}}}\right)$; and ($b$) $\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left|\widehat{\boldsymbol{\alpha}}-\boldsymbol{\alpha}^{*}}}\right|}}\right|_{1}=O_{P}\mathopen{}\mathclose{{\left(s\sigma_{x\varepsilon}\sqrt{\frac{log(n)}{T}}}}\right)$, where $\sigma_{x\varepsilon}^{2}=\frac{1-\phi^{2}\phi_{\varepsilon}^{2}}{(1-\phi_{\varepsilon}^{2})(1-\phi\phi_{\varepsilon})^{2}}$.
Proposition 2 and Corollary 1 show that oracle inequalities and convergence rates for the LASSO critically depend on $\sigma_{x\varepsilon}^{2}$ and $\gamma_{x}^{2}$. In turn, the results in Section 2.1 show that these key quantities are affected by serial correlation. Hence, as $\phi$ and $\phi_{\varepsilon}$ increase, the bounds in Proposition 2 become larger and the convergence rates in Corollary 1 become slower. Moreover, an increase in $\phi$ leads to an increase in sample correlations (see Proposition 1), which in turn leads to a reduction of $\gamma_{x}^{2}$ (see Remark 2), further amplifying the bounds in Proposition 2.
3 The ARMAr-LASSO
We now switch to describing ARMAr-LASSO (ARMA residuals LASSO), the approach that we propose to rescue LASSO performance in the presence of serially correlated predictors. ARMAr-LASSO is formulated as a two-step procedure. In the first step we estimate a univariate ARMA model on each predictor. In the second step, we run the LASSO using, instead of the original predictors, the residuals from the ARMA model, i.e. estimates of the $u$’s in equation (2), plus lags of the response. We start by introducing the “working model” on which our proposal relies; that is, the model that contains the true, non-observable ARMA residuals (their estimation will be addressed later)
$y_{t}=\mathbf{w}_{t}^{\prime}\boldsymbol{\beta}^{*}+v_{t}\ \ .$ (1)
Model (1) is the linear projection of $y_{t}$ on $\mathbf{w}_{t}=(u_{1,t},\dots,u_{n,t},y_{t-1},\dots,y_{t-p_{y}})^{\prime}$, which contains $n$ ARMA residuals and $p_{y}$ lagged values of the response. $\boldsymbol{\beta}^{*}=(\boldsymbol{\alpha}^{*^{\prime}},\phi_{y,1},\dots,\phi_{y,p_{y}})^{\prime}$ represents the corresponding best linear projection coefficients and $v_{t}$ is the error term, which is unlikely to be $i.i.d.$. It should be noted that the choice of $p_{y}$ is arbitrary and that some lags will be relevant while others will not. The relevant lags will be directly selected using LASSO. Model (1) is misspecified, in the sense that it does not correspond to the true data generating process (DGP) for the response, but it is similar in spirit to the factor filter used in the literature to mitigate multicollinearity (Fan2020). The idea behind model (1) is to leverage the serial independence of the $u$ terms, thereby avoiding the risk of sizeable spurious correlation. However, the $u$ terms alone may explain only a small portion of the variance of $y_{t}$, particularly in situations with high persistence. This is why we introduce the response lags as additional predictors; these amplify the signal in our model and consequently improve the forecast of $y_{t}$. Furthermore, the inclusion of lagged terms also helps mitigating serial correlation in the residuals.
Remark 1:
When $p_{y}=0$, ARMAr-LASSO uses only the $u_{t}$ as predictors, leaving $v_{t}$ to capture all serial correlation. Including lags of $y_{t}$ ($p_{y}>0$) allows one to model and exploit serial correlation. Under the common AR($p$) restriction, $s_{y}=p$ ensures that $v_{t}$ is a white noise; using fewer lags ($s_{y}<p$) reduces forecasting performance. Without this restriction, ARMAr-LASSO automatically selects $s_{y}$ based on the maximum AR/ARMA order of the predictors and errors. Although a formal theoretical proof of this result is beyond the scope of the present study, our numerical evidence shows that when $p_{y}$ exceeds the maximum order of the predictors and/or error terms, ARMAr-LASSO consistently outperforms the benchmark methods (see Section 4.1.1).
We list some important facts that capture how misspecification affects coefficient estimation and feature selection.
Fact 1:
(on the ARMA residuals) ($a$) $E(v_{t}|\mathbf{u}_{t})=0$; ($b$) $E(\mathbf{u}_{t}y_{t-l})=\boldsymbol{0}$, $\forall$ $l\geq 1$, and $E(u_{it}y_{t-l}|u_{it-1}y_{t-l-1},u_{it-2}y_{t-l-2},\dots)=0$, $\forall$ $i,j\geq 1$.
Fact 1 follows from Assumption 1. Fact 1 ($a$) ensures that the least square estimator of $\boldsymbol{\alpha}^{*}$ is unbiased and consistent. Fact 1 ($b$) is crucial for feature selection among the $u$’s. In particular, $E(\mathbf{u}_{t}y_{t-l})=\boldsymbol{0}$ removes population level multicollinearity, while $E(u_{it}y_{t-l}|u_{it-1}y_{t-l-1},u_{it-2}y_{t-l-2},\dots)=0$ removes the risk of spurious correlation due to serial correlation (see Section 2.1).
Fact 2:
(on the lags of $y_{t}$) ($a$) $E(v_{t}|y_{t-1},y_{t-2},\dots)$ can be $\neq 0$; ($b$) $E(y_{t-l}|y_{t-l-1},y_{t-l-2},\dots)\neq 0$, $\forall$ $l\geq 0$.
Fact 2 ($a$) relates to the possible misspecification of the working model (1), which leads to an endogeneity problem between $v_{t}$ and the lags of $y_{t}$. However, as previously said, the lags of $y_{t}$ and the corresponding parameters $\phi_{y,1},\dots,\phi_{y,p_{y}}$ are introduced to enhance the variance explained, and thus the ability to forecast the response – tolerating a potential endogenous variable bias. Fact 2 ($b$) relates to potential correlations between the lags of $y_{t}$, which is serial in nature. This implies that relevant lags may be represented by irrelevant ones. However, selection of relevant lags of $y_{t}$ is not of interest in this context.
Next, we provide three illustrative examples. In the first, and simplest, predictors and error terms have an AR(1) representation with a common coefficient; we refer to this as the common AR(1) restriction case. In the second, the AR(1) processes have different autoregressive coefficients. In the third, predictors admit a common factor representation with AR(1) idiosyncratic components. Note that in all the examples $p_{y}=1$.
Example 1:
(common AR(1) restriction). Suppose both predictors and error terms in model (1) admit a common AR(1) representation; that is, $x_{i,t}=\phi x_{i,t-1}+u_{i,t}$ and $\varepsilon_{t}=\phi\varepsilon_{t-1}+\omega_{t}$. In this case $y_{t}=\sum_{i=1}^{n}\alpha_{i}^{*}x_{i,t}+\varepsilon_{t}=\sum_{i=1}^{n}\alpha_{i}^{*}(\phi x_{i,t-1}+u_{i,t})+\phi\varepsilon_{t-1}+\omega_{t}=\sum_{i=1}^{n}\alpha_{i}^{*}u_{i,t}+\phi y_{t-1}+\omega_{t}.$ Thus, under the common AR(1) restriction (also known as common factor restriction, Mizon1995), the working model (1) is equivalent to the true model (1) because of the decomposition of the AR(1) processes $\{\mathbf{x}_{t}\}$ and $\{\varepsilon_{t}\}$.
Remark 2:
The working model (1) coincides with the true model (1) under a common AR($p$) restriction; that is, when $x_{i,t}=\sum_{l=1}^{p}\phi_{l}x_{i,t-l}+u_{i,t}$ and $\varepsilon_{t}=\sum_{l=1}^{p}\phi_{l}\varepsilon_{t-l}+\omega_{t}$. In fact, it is easy to show that $y_{t}=\sum_{i=1}^{n}\alpha_{i}^{*}x_{i,t}+\varepsilon_{t}=\sum_{i=1}^{n}\alpha_{i}^{*}u_{i,t}+\sum_{l=1}^{p}\phi_{l}y_{t-l}+\omega_{t}$ for any autoregressive order pp. Moreover, in this case $v_{t}=\omega_{t}$ and $E(v_{t}|\mathbf{w}_{t})=0$ – so we have unbiasedness and consistency also for the coefficients of the lags of $y_{t}$.
Example 2:
(different AR(1) coefficients). Suppose $x_{i,t}=\phi_{i}x_{i,t-1}+u_{i,t}$ and $\varepsilon_{t}=\phi_{\varepsilon}\varepsilon_{t-1}+\omega_{t}$, where $u_{i,t},\ \omega_{t}\sim i.i.d.\ N(0,1)$. Then the working model (1) has $v_{t}=\sum_{i=1}^{n}(\phi_{i}-\phi_{y})x_{i,t-1}+(\phi_{\varepsilon}-\phi_{y})\varepsilon_{t-1}+\omega_{t},$ where $\phi_{y}=\frac{E(y_{t}y_{t-1})}{E(y_{t}^{2})}=\mathopen{}\mathclose{{\left(\sum_{i=1}^{n}\frac{\phi_{i}\alpha_{i}^{*2}}{1-\phi_{i}^{2}}+\frac{\phi_{\varepsilon}}{1-\phi_{\varepsilon}^{2}}}}\right)/\mathopen{}\mathclose{{\left(\sum_{i=1}^{n}\frac{\alpha_{i}^{*2}}{1-\phi_{i}^{2}}+\frac{1}{1-\phi_{\varepsilon}^{2}}}}\right)$. Therefore, $E(v_{t}|\mathbf{u}_{t})=0$ and $E(v_{t}|y_{t-1})=\sum_{i=1}^{n}(\phi_{i}-\phi_{y})x_{i,t-1}+(\phi_{\varepsilon}-\phi_{y})\varepsilon_{t-1}\neq 0$.
Example 3:
(common factor). Suppose $x_{i,t}=\lambda_{i}f_{t}+z_{t},\ f_{t}=\phi_{f}f_{t-1}+\delta_{t},\ z_{i,t}=\phi_{i}z_{i,t-1}+\eta_{i,t}$ and $\varepsilon_{t}=\phi_{\varepsilon}\varepsilon_{t-1}+\omega_{t},$ where $\delta_{t},\eta_{it},\omega_{t}\sim i.i.d\ N(0,1).$ In this case, any $x_{it}$ is a sum of two independent AR(1) processes and, therefore, $x_{it}\sim ARMA(2,1)$ (Granger1976). Again, by Assumption 1, we have $E(v_{t}|\mathbf{u}_{t})=0$ and $E(v_{t}|y_{t-1})\neq 0$.
In the next section, we will provide some theoretical results concerning the use of the LASSO estimator of $\boldsymbol{\beta}^{*}$ in working model (1), which is obtained as
$\widehat{\boldsymbol{\beta}}=\underset{\boldsymbol{\beta}\in\mathbb{R}^{n+p_{y}}}{\operatornamewithlimits{argmin\ }}\>\mathopen{}\mathclose{{\left\{\>\frac{1}{2T}\>\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left|\mathbf{y}-\mathbf{W}^{\prime}\boldsymbol{\beta}}}\right|}}\right|_{2}^{2}+\lambda||\boldsymbol{\beta}||_{1}\>}}\right\}\ \ ,$ (2)
where $\lambda>0$ is a tuning parameter. In particular, in Section 3.3, we will provide the limiting distribution and feature selection consistency of (2) in the classical framework with $n$ fixed and $T\rightarrow\infty$. Next, in Section 3.4, we will establish oracle inequalities for the estimation and forecast error bounds of the ARMAr-LASSO, allowing $n$ to grow as a function of $T$ (i.e., $n=n_{T}$). We will also tackle the problem of estimating the $u$’s. Henceforth, we assume that each row of the $(n+p_{y})\times T$ design matrix $\mathbf{W}=\{\mathbf{w}_{t}\}_{t=1}^{T}$ is standardized to have mean 0 and variance 1, which implies $\frac{1}{T}\underset{1\leq t\leq T}{\text{max}}\mathbf{w}_{t}^{\prime}\mathbf{w}_{t}\stackrel{{\scriptstyle p}}{{\rightarrow}}0$. Moreover, $\widehat{\mathbf{C}}_{w}=\frac{1}{T}\sum_{t=1}^{T}\mathbf{w}_{t}\mathbf{w}_{t}^{\prime}\hskip 2.84544pt\stackrel{{\scriptstyle a.s.}}{{\rightarrow}}\mathbf{C}_{w},$ where $\mathbf{C}_{w}=E(\mathbf{w}_{t}\mathbf{w}_{t}^{\prime})$ is a non-negative definite matrix.
Let $\mathbf{q}_{t}=(\mathbf{w}_{t}^{\prime},v_{t})^{\prime}$. To derive theoretical results for ARMAr-LASSO, we rely on the fact that, due to Assumption 1, $\mathbf{q}_{t}$ depends almost entirely on the “near epoch” of its shock. In particular, it is characterized as near-epoch dependent (NED) (refer to davidson1994 ch. 17 and ADAMEK2023 for details). NED is a very popular tool for modelling dependence in econometrics. It allows for cases where a variable’s behaviour is primarily governed by the recent history of explanatory variables or shock processes, potentially assumed to be mixing. davidson1994 shows that even if a variable is not mixing, its reliance on the near epoch of its shocks makes it suitable for applying limit theorems, particularly the mixingale property (see Supplement A.5 for details). The NED framework accommodates a wide range of models, including those that are misspecified as our working model (1). For instance, in Examples 2 and 3, $(\mathbf{w}_{t}^{\prime},v_{t})$ have a moving average representation with geometrically decaying coefficients, and are thus NED on $(\mathbf{u}_{t}^{\prime},\omega_{t})$ and $(\delta_{t},\boldsymbol{\eta}_{t}^{\prime},\omega_{t})$, respectively.
3.1 Estimation of ARMA residuals
In this section, we summarize the asymptotic properties of the Bayesian Information Criterion (BIC) and the Maximum Likelihood (ML) estimator for identifying ARMA orders and estimating model parameters, respectively.
[…]
Note that, under model misspecification, parameter estimates converge to pseudo‑true values and residuals converge to pseudo‑innovations rather than the true ones (see, e.g., potscher1991). These properties justify using the estimated coefficients and residuals as asymptotically valid approximations to their true counterparts.
3.2 Least Squares estimator applied to the Working Model (1)
To clarify the statistical properties of the coefficients penalized by ARMAr-LASSO, we first analyze the behavior of the corresponding Least Squares estimator applied to the ARMAr working model (ARMAr-LS). Consider the univariate model $y_{t}=\alpha x_{t-1}+\varepsilon_{t}$, where $x_{t}=\phi x_{t-1}+u_{t}$ and $\varepsilon_{t}=\phi_{\varepsilon}\varepsilon_{t-1}+\omega_{t}$, with $u_{t}$ and $\omega_{t}$ being serially uncorrelated innovations. Serial correlation is therefore present in both the predictor and the error term. The ARMAr transformation yields the working model $y_{t}=\alpha u_{t-1}+\phi_{y}y_{t-1}+v_{t}$, where the regressor $u_{t-1}$ is serially uncorrelated by construction, and the composite error $v_{t}$ collects the remaining dynamic components. Although $v_{t}$ is generally correlated with $y_{t-1}$, it is conditionally mean–independent of $u_{t-1}$ under Assumption 1. Consequently, the ARMAr-LS estimator satisfies $\widehat{\alpha}=\alpha+\frac{\sum_{t=1}^{T-1}u_{t}v_{t}}{\sum_{t=1}^{T-1}u_{t}^{2}}$, which immediately implies three key properties.
First, the estimator is unbiased, since $E(v_{t}|\mathbf{u})=0$ and therefore $E(\widehat{\alpha}|\mathbf{u})=\alpha$. Second, it is consistent, because $plim(\widehat{\alpha})=\alpha$ by exogeneity of $u_{t}$. Third, under the common AR(1) restriction $\phi=\phi_{\varepsilon}$, ARMAr-LS attains GLS efficiency, with $v_{t}=\omega_{t}$ and conditional variance $Var(\widehat{\alpha}|\mathbf{u})=\frac{\sigma_{\omega}^{2}}{\sum_{t=1}^{T-1}u_{t}^{2}}$.
These results are crucial for understanding the behavior of ARMAr-LASSO. In contrast to standard LASSO, which penalizes OLS coefficients that may be inefficient, biased, or inconsistent in the presence of serially correlated predictors Keele2005, ARMAr-LASSO penalizes coefficients that are already unbiased and consistent.
3.3 ARMAr-LASSO: Asymptotic Results
This section is devoted to the asymptotic behaviour and feature selection consistency of the LASSO applied to working model (1), within the classical setting with $n$ fixed and $T\rightarrow\infty$. We will extend some known results to our context to demonstrate that the working model (1) retains the usual inferential and selection consistency properties, despite being a misspecification of the true model (1). Our results build upon Theorem 2 of Fu2000 and Theorem 1 of zhaoyu2006. In the classic asymptotic setting, the facts summarized in Section 3.1 allow us to derive properties directly for $u_{t}$ rather than for the estimate $\widehat{u}_{t}$. Let $\boldsymbol{\mu}_{vy}=\mathopen{}\mathclose{{\left(E(v_{t}y_{t-1}),\dots,E(v_{t}y_{t-p_{y}})}}\right)^{\prime}$ be the mean vector and $\boldsymbol{\Gamma}_{vy}$ the $p_{y}\times p_{y}$ covariance matrix of $\mathopen{}\mathclose{{\left(v_{t}y_{t-1},\dots,v_{t}y_{t-p_{y}}}}\right)$. The following theorem provides the asymptotic behaviour of the LASSO solution.
Theorem 1:
Let Assumption 1 holds. If $\lambda\sqrt{T}\rightarrow\lambda_{0}\geq 0$ and $\mathbf{C}_{w}$ is nonsingular, the solution $\widehat{\boldsymbol{\beta}}$ of (2) is such that $\sqrt{T}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*})\hskip 2.84544pt\stackrel{{\scriptstyle d}}{{\rightarrow}}\hskip 2.84544pt\underset{\mathbf{a}\in\mathbb{R}^{n+p_{y}}}{\operatornamewithlimits{argmin\ }}(V(\mathbf{a}))$, where $V(\mathbf{a})=-2\mathbf{a}^{\prime}\mathbf{m}+\mathbf{a}^{\prime}\mathbf{C}_{w}\mathbf{a}+\lambda_{0}\sum_{i=1}^{n+p_{y}}\mathopen{}\mathclose{{\left[a_{i}Sign(\beta_{i}^{*})I(\beta_{i}^{*}\neq 0)+|a_{i}|I(\beta_{i}^{*}=0)}}\right],$ and $\mathbf{m}$ is an $n+p_{y}$ dimensional random vector with a $N\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left(\boldsymbol{0}_{n}^{\prime},\boldsymbol{\mu}_{vy}}}\right)^{\prime},\mathopen{}\mathclose{{\left({\begin{array}[]{cc}\sigma_{v}^{2}\mathbf{C}_{u}&\boldsymbol{0}_{n\times p_{y}}\\ \boldsymbol{0}_{p_{y}\times n}&\boldsymbol{\Gamma}_{vy}\\ \end{array}}}}\right)}}\right)$ distribution.
Next, we consider the feature selection properties of (2). Let $s_{y}\leq p_{y}$ denote the number of relevant lags of $y_{t}$, and separate the coefficients of relevant and irrelevant features into $\boldsymbol{\beta}^{*}(1)=(\alpha_{1}^{*},\dots,\alpha_{s}^{*},\phi_{y1},\dots,\phi_{ys_{y}})^{\prime}$ and $\boldsymbol{\beta}^{*}(2)=(\alpha_{s+1}^{*},\dots,\alpha_{n}^{*},\phi_{ys_{y}+1},\dots,\phi_{yp_{y}})^{\prime}$, respectively. Also, let $\mathbf{W}(1)$ and $\mathbf{W}(2)$ denote the rows of $\mathbf{W}$ corresponding to relevant and irrelevant features. We can rewrite $\widehat{\mathbf{C}}_{w}$ in block-wise form as
$\widehat{\mathbf{C}}_{w}=\mathopen{}\mathclose{{\left({\begin{array}[]{cc}\widehat{\mathbf{C}}_{w}(11)&\widehat{\mathbf{C}}_{w}(12)\\ \widehat{\mathbf{C}}_{w}(21)&\widehat{\mathbf{C}}_{w}(22)\\ \end{array}}}}\right)\ \ ,$
where $\widehat{\mathbf{C}}_{w}(11)=\frac{1}{T}\mathbf{W}(1)\mathbf{W}(1)^{\prime}$, $\widehat{\mathbf{C}}_{w}(22)=\frac{1}{T}\mathbf{W}(2)\mathbf{W}(2)^{\prime}$, $\widehat{\mathbf{C}}_{w}(12)=\frac{1}{T}\mathbf{W}(1)\mathbf{W}(2)^{\prime}$ and $\widehat{\mathbf{C}}_{w}(21)=\frac{1}{T}\mathbf{W}(2)\mathbf{W}(1)^{\prime}$. We then introduce a critical assumption on $\widehat{\mathbf{C}}_{w}$.
Assumption 1:
(strong irrepresentable condition (zhaoyu2006)) Assuming $\widehat{\mathbf{C}}_{w}(11)$ is invertible, $|\widehat{\mathbf{C}}_{w}(21)(\widehat{\mathbf{C}}_{w}(11))^{-1}Sign(\boldsymbol{\beta}^{*}(1))|<\boldsymbol{1}-\varphi,$ where $\varphi\in(0,1)$ and the inequality holds element-wise.
[…]
Theorem 2:
Let Assumptions 1 and 1 hold. If $\lambda\sqrt{T}\rightarrow\lambda_{0}\geq 0$, then the solution $\widehat{\boldsymbol{\beta}}$ of (2) is such that $P\mathopen{}\mathclose{{\left(Sign(\widehat{\boldsymbol{\beta}})=Sign(\boldsymbol{\beta}^{*})}}\right)\rightarrow 1.$
The theoretical results provided in this section show that under Assumptions 1 and 1, and as a consequence of Fact 1, ARMAr-LASSO guarantees consistent estimation, asymptotic normality, as well as consistent feature selection for the vector $\boldsymbol{\alpha}^{*}$.
Remark 3:
The working model (2) underlying ARMAr-LASSO constitutes a misspecification of the true data-generating process, which induces endogeneity due to lags of $y_{t}$. Theorems 1 and 2 are designed to show that, despite this endogeneity, ARMAr-LASSO preserves the classical properties of estimation and inference for the parameters of interest, namely the $\alpha$ coefficients associated with the ARMA residuals. Establishing these results is important to confirm that ARMAr-LASSO is a reliable tool for improving the estimation and forecasting performance of LASSO in settings where both predictors and errors exhibit serial correlation. Building on this foundation, we next turn to asymptotic results in a high-dimensional setting where both $n$ and $T$ diverge. This requires different conditions on the regularization parameter $\lambda$. Indeed, the optimal scaling of $\lambda$ depends on the setting: in high-dimensional asymptotics it typically follows $\sqrt{\log n/T}$, whereas in classical fixed-$n$ asymptotics a different scaling applies. Hence, distinct asymptotic regimes naturally imply different choices of $\lambda$ (see, e.g., Buhlmann2011; zhaoyu2006).
3.4 ARMAr-LASSO: Oracle Inequalities
In this section, we derive the oracle inequalities that provide bounds for the estimation and forecast errors of the ARMAr-LASSO. Here, we allow $n$ to grow as $T$ grows; that is, we pursue results in a framework where $n=n_{T}=O(T^{a})$ for some $a>0$. This condition serves as a broad upper bound and accommodates a variety of growth rates. For example, $n_{T}$ could grow as slowly as $log(T)$ or as fast as $T^{a}$. In this context, we replace the predictor vector $\mathbf{w}_{t}$ with $\widehat{\mathbf{w}}_{t}=(\widehat{u}_{1,t},\dots,\widehat{u}_{n_{T},t},y_{t-1},\dots,y_{t-p_{y}})^{\prime}$, where the $\widehat{u}$’s are obtained by employing BIC and ML as described in Section 3.1. We need the following Assumption, which bounds the unconditional moments of the predictors in the true model (1), and of $\widehat{\mathbf{w}}_{t}$ and $v_{t}$.
Assumption 2:
Consider $\mathbf{q}_{t}=(\widehat{\mathbf{w}}_{t}^{\prime},v_{t})^{\prime}$. There exist constants $c_{2}>c_{1}>2$ such that $\underset{i\leq n_{T}+p_{y}+1,\ t\leq T}{\text{max}}E(\mathopen{}\mathclose{{\left|q_{i,t}}}\right|^{2c_{2}})\leq C$ and $\underset{i\leq n_{T},t\leq T}{\text{max}}E(\mathopen{}\mathclose{{\left|x_{i,t}}}\right|^{2c_{2}})\leq C$.
Remark 4:
The error term $v_{t}$ is modeled as a stationary Near-Epoch-Dependent (NED) process which, under appropriate decay conditions, can be approximated by a strongly mixing sequence. In particular, this includes the case of stationary, and finite-order ARMA processes, which are well known to be strongly mixing with geometric decay of the mixing coefficients (see, e.g., davidson1994). For valid asymptotic results, we assume that the error term has finite $q$-th moments for some $q>4$, ensuring tails sufficiently light for our limit theory, although stronger assumptions such as exponential moment bounds could also be accommodated.
To derive the error bound of the ARMAr-LASSO estimator from (2), we follow the typical procedure presented in technical textbooks (see, e.g., Buhlmann2011, ch. 6). We need $\lambda$ to be sufficiently large as to exceed the empirical process $\underset{i\leq n_{T}+p_{y},t\leq T}{\text{max}}\mathopen{}\mathclose{{\left|\sum_{t=1}^{T}w_{i,t}v_{t}}}\right|$ with high probability.
Theorem 3:
Let Assumption 1 and 2 hold and define $\mathcal{A}_{T}\coloneqq\mathopen{}\mathclose{{\left\{\underset{i\leq n_{T}+p_{y},l\leq T}{\text{max}}\mathopen{}\mathclose{{\left|\sum_{t=1}^{l}\widehat{w}_{i,t}v_{t}}}\right|\leq\frac{T\lambda}{4}}}\right\}$. Furthermore, assume that $T$ and $n_{T}$ are sufficiently large as to have $\lambda\geq C\mathopen{}\mathclose{{\left(\sqrt{log(T)}}}\right)^{1/c_{1}}\frac{(2n_{T}+p_{y})^{1/c_{1}}}{\sqrt{T}}$. Then $\Pr(\mathcal{A}_{T})\geq 1-C\mathopen{}\mathclose{{\left(\sqrt{\text{log(T)}}}}\right)^{-1}$.
Theorem 3 establishes that the inequalities we need for the error bound of the proposed ARMAr-LASSO estimator hold with high probability. The bounds used in the proof of Theorem 3 put implicit limits on the divergence rate of $n_{T}$ relative to $T$. The term $\sqrt{\text{log($T$)}}$ is chosen arbitrarily as a sequence that grows slowly as $T\rightarrow\infty$. However, we can use any sequence that tends to infinity sufficiently slowly. For example, ADAMEK2023 use $\text{log}(\text{log}(T))$ to derive properties of the LASSO in a high-dimensional time series model under weak sparsity. We introduce the assumption on the “restricted” positive definiteness of the covariance matrix of the predictors, which allows us to generalize subsequent results to the high-dimensional framework.
Remark 5:
The concentration arguments in Theorem 3 are based on the NED-mixingale framework of ADAMEK2023, which allows for very general forms of temporal dependence, random regressors, and only finite moment assumptions. This choice is particularly suited to our projection-based and potentially misspecified regression setting with lagged dependent variables and estimated components. Sharper deviation bounds, and hence more aggressive tuning rates, could be obtained under stronger structural assumptions using, for example, the functional dependence approach of WuWu2016 or Fuk-Nagaev type inequalities for $\tau$- or $s$-mixing processes as in Babii2024. These alternatives, however, require additional restrictions on mixing rates, tail behavior, and the design structure. We therefore favor the more general framework of ADAMEK2023, which yields slightly more conservative but broadly valid theoretical guarantees.
Assumption 3:
[…]
Remark 6:
Let $\gamma_{x}^{2}$ be the compatibility constant of the restricted eigenvalue of $\frac{1}{T}\mathbf{XX}^{\prime}$. Since this captures how strongly predictors are correlated in the sample, as a consequence of the theoretical treatment of Sections 2.1 and 2.2, we have $\gamma_{\widehat{w}}^{2}>\gamma_{x}^{2}$ with high probability as the degree of serial correlation increases (when both $\frac{1}{T}\widehat{\mathbf{W}}\widehat{\mathbf{W}}^{\prime}$ and $\frac{1}{T}\mathbf{XX}^{\prime}$ are nonsingular, we have $\widehat{\psi}_{min}^{\widehat{w}}>\widehat{\psi}_{min}^{x}$ with high probability). Of course, $\gamma_{\widehat{w}}^{2}$ and $\gamma_{x}^{2}$ also depend on the cardinalities $\tilde{s}$ and $s$. However, here we emphasize the role of serial correlation.
The following theorem, which expresses the oracle inequalities for the ARMAr-LASSO, is a direct consequence of Theorem 3.
Theorem 4:
Let Assumptions 1, 2 and 3 hold. Furthermore, let the conditions of Theorem 3 hold. When assume that $T$ and $n_{T}$ are sufficiently large, the following oracle inequalities hold simultaneously with probability at least $1-C\mathopen{}\mathclose{{\left(\sqrt{\text{log(T)}}}}\right)^{-1}$: ($a$) $\frac{1}{T}\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left|\widehat{\mathbf{W}}^{\prime}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*})}}\right|}}\right|_{2}^{2}\leq\frac{4\tilde{s}\lambda^{2}}{\gamma_{\widehat{w}}^{2}}$; ($b$) $\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left|\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}}}\right|}}\right|_{1}\leq\frac{4\tilde{s}\lambda}{\gamma_{\widehat{w}}^{2}}$. In addition, if $\underset{j\in\tilde{S}}{\text{min}}|\beta_{j}^{*}|>\frac{4\tilde{s}\lambda}{\gamma_{\widehat{w}}^{2}},$ ARMAr-LAS enjoys the variable screening property; that is, it correctly identifies all true non-zero coefficients.
Corollary 1:
Under the additional assumption that $\tilde{s}\lambda\rightarrow 0$ one can also establish, as an immediate corollary to Theorem 4, the following convergence rates: ($a$) $\frac{1}{T}\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left|\widehat{\mathbf{W}}^{\prime}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*})}}\right|}}\right|_{2}^{2}=O_{P}\mathopen{}\mathclose{{\left(\frac{\tilde{s}}{T}\mathopen{}\mathclose{{\left((n_{T}+p_{y})\mathopen{}\mathclose{{\left(\sqrt{log(T)}}}\right)}}\right)^{2/c_{1}}}}\right)$; ($b$) $\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left|\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}}}\right|}}\right|_{1}=O_{P}\mathopen{}\mathclose{{\left(\frac{\tilde{s}}{\sqrt{T}}\mathopen{}\mathclose{{\left((n_{T}+p_{y})\mathopen{}\mathclose{{\left(\sqrt{log(T)}}}\right)}}\right)^{1/c_{1}}}}\right)$.
4 Simulations and Empirical Application
In this section, we analyse the performance of the ARMAr-LASSO by means of both simulations and a real data application.
4.1 Simulation Experiments
The response variable is generated using the model $y_{t}=\sum_{i=1}^{n}\alpha_{i}^{*}x_{i,t-1}+\varepsilon_{t}$, and we consider the following data generating processes (DGPs) for predictors and error terms:
  • (A)
    Common AR(2) restriction: $x_{i,t}=0.45x_{i,t-1}+0.45x_{i,t-2}+u_{i,t},\ \varepsilon_{t}=0.45\varepsilon_{t-1}+0.45\varepsilon_{t-2}+\omega_{t}$.
  • (B)
    General AR/ARMA: $x_{i,t}=qf_{t}+z_{i,t}$, where $f_{t}=0.9f_{t-1}+\delta_{t}$, ${z}_{j,t}=0.8{z}_{j,t-1}+\eta_{j,t}$; ${z}_{h,t}=0.6{z}_{h,t-1}+0.3{z}_{h,t-2}+\eta_{h,t}$; $z_{w,t}=0.5z_{w,t-1}+0.4z_{w,t-2}+\eta_{w,t}+0.3\eta_{w,t-1}$; $z_{k,t}=0.7z_{k,t-1}+\eta_{k,t}+0.4\eta_{k,t-1}$, for $t=1,\dots,T$, and $j=1,\dots,4;\ h=5,\dots,7;\ w=7,\dots,10;\ k=11,\dots,n$. The error terms are generated as $\varepsilon_{t}=0.7\varepsilon_{t-1}+0.2\varepsilon_{t-2}+\omega_{t}$.
The shocks are generated as follows: $u_{i,t}\sim i.i.d.\ N(0,1)$ with $\mathopen{}\mathclose{{\left(\mathbf{C}_{u}}}\right){ij}=c_{ij}^{u}=\rho^{|i-j|}$, $\delta_{t},\ \eta_{i,t}\sim i.i.d.\ N(0,1)$ with $\mathopen{}\mathclose{{\left(\mathbf{C}_{\eta}}}\right){ij}=c_{ij}^{\eta}=\rho^{|i-j|}$, and $\omega_{t}\sim i.i.d.\ N(0,\sigma_{\omega}^{2})$. For the DGP (A) and for the DGP (B) with $q=0$ we set $\rho=0.8$, while for the DGP (B) with $q=1$ we set $\rho=0.4$ to generate predictors primarily influenced by the common factor, with weakly correlated AR or ARMA idiosyncratic components. Finally, we vary the value of $\sigma_{\omega}^{2}$ to explore different signal-to-noise ratios (SNRs). For each DGPs, the performance of ARMAr-LASSO and benchmarks is evaluated based on average results from 1000 independent simulations, focusing on the coefficient estimation error (CoEr) obtained as $\widehat{\boldsymbol{\alpha}}-\boldsymbol{\alpha}||_{2}$, the Root Mean Square Forecast Error (RMSFE), and the percentages of true positives (%TP) and false positives (%FP) in selecting relevant predictors. Regardless of the choice of $n$, $\boldsymbol{\alpha}^{*}$ is always taken to have the first $10$ entries equal to $1$ and all others equal to 0.
4.1.1 DGP(A): Filters Evaluation Under Common AR restriction.
We test several ARMAr-LASSO settings in terms of the number of lags of $y_{t}$ included in the model ($p_{y}$), the order of the AR filter used to obtain the estimated $\widehat{u}$’s ($p_{i}$), and the model selection method for the filter. We consider as reference setting the case where $p_{y}=3$ and the $\widehat{u}$’s are obtained by filtering each variable with an AR($p_{i}$), where $p_{i}$ (max 3) is selected via BIC and the AR parameters are estimated via ML (see Section 3.1). The dimensionality is kept fixed at $n=150$, while the sample size varies; we consider $T=75,150,300$. Both CoEr and RMSFE are relative to the working model 1, which includes the true $u$’s. Results are summarized in Table 1. To assess the impact of including lags of $y_{t}$, we exclude them from the penalization process and compare the cases with one lag ($y_{t-1}$) and three lags ($y_{t-1}^{t-3}$) treated as fixed predictors in the model. Note that under DGP (A), the number of lags of $y_{t}$ that yields white-noise error terms is $p_{y}=2$. Using too few lags ($y_{t-1}$) reduces both estimation and forecast accuracy, whereas using an excessive number of lags ($y_{t-1}^{t-3}$) does not provide additional benefits since, especially for large $T$, the performance coincides with that of ARMAr-LS. To evaluate the effect of filter misspecification, instead of selecting $p_{i}$ using BIC, we filter the variables with AR(1), AR(2), and AR(3) processes, thus under-specifying, correctly-specifying, and over-specifying the true order, respectively. Underestimating the true order worsens coefficient estimation and forecast accuracy, particularly at high SNRs, while over-specifying the true order provides no gains. Finally, we evaluate alternative model selection criteria other than BIC. In particular, we consider the Akaike Information Criterion (AIC), its small-sample corrected version (AICc), and block cross-validation (blCV), which is a time-series variant of cross-validation (racine1997). For blCV, we use a rolling window of size T3T-3, and for each variable, we select the order $p_{i}$ (max 3) that best predicts one step over the three out-of-sample horizons. All methods yield similar results, except for blCV, which performs worse than the others for high SNRs and large $T$.
4.1.2 DGP(B): General AR/ARMA.
We compare our ARMAr-LASSO (ARMAr-LAS) with the standard LASSO applied to the observed time series (LAS), LASSO applied to the observed time series plus lags of $y_{t}$ (LASy), GLS-LASSO as proposed by chronopoulos2023 (GLS-LAS), autoregressive distributed lag LASSO (ARDL-LAS), and FarmSelect as proposed by Fan2020, which employs LASSO on factor model residuals (FaSel). For all methods, the tuning parameter $\lambda$ is selected using the Bayesian Information Criterion (BIC). For GLS-LAS, we filter both response and predictors using the coefficients of an AR($p_{\varepsilon}$) model applied to $\widehat{\varepsilon}_{t}$, with the order $p_{\varepsilon}$ (max 2) selected with BIC. For ARDL-LAS, we consider two lags of the response and two lags of each predictor as additional regressors – bringing the number of term undergoing selection to $n\times 3+2$. For the working model underlying ARMAr-LAS, the $\widehat{u}$’s are obtained by filtering each series with an ARMA($p_{i},q_{i}$) process, with the orders $p_{i}$ and $q_{i}$ (max 2) selected via BIC. We consider $p_{y}=3$; that is, three lags of $y_{t}$ as additional predictors. Simulations have varying numbers of predictors (dimensionality), $n=75,150,300$, and a fixed sample size, $T=150$. In this way, we cover low ($n=75$), intermediate ($n=150$), and high ($n=300$) dimensional scenarios, and also cover different levels of sparsity, consequently to $\boldsymbol{\alpha}^{*}||_{0}=10$. Results are presented in Table 2. ARMAr-LAS outperforms all other LASSO-based methods in terms of estimation accuracy, forecasting, and feature selection, regardless of the SNR and the presence of a common factor (i.e., when $q=1$). In particular, ARMAr-LASSO removes serial correlation in the predictors and the resulting spurious correlations, yielding more accurate estimation and forecasts than GLS-LASSO. Unlike GLS-LASSO, which fully eliminates predictor serial correlation only under the restrictive common AR($p$) condition, ARMAr-LASSO achieves this goal without requiring the same dynamic structure for predictors and errors. Moreover, ARMAr-LASSO requires only a few lags of $y_{t}$ as additional predictors, making it considerably more parsimonious than ARDL-LASSO, which quickly becomes over-parameterized when multiple lags are included. The effectiveness of our proposal in this realistic setting highlights its suitability also when tackling differing AR and ARMA processes and common factors, where the common AR($p$) restriction does not hold.
In Supplement D.3, we compare the minimum eigenvalues of the predictors correlation matrix of ARMAr-LASSO with those of LASSO and GLS-LASSO. Results show that ARMAr-LASSO relies on a correlation matrix that exhibits a larger minimum eigenvalue than the classical LASSO and GLS-LASSO. Notably, this corroborates the statement of Remark 6.
Table 1: DGP (A). CoEr, RMSFE (relative to WM), %TP and %FP for various ARMAr-LASSO settings. For each $T$ setting the best CoEr and RMSFE are in bold.
SNR = 1 $T$ 75 150 300 CoEr RMSFE % TP % FP CoEr RMSFE % TP % FP CoEr RMSFE % TP % FP ARMAr-LAS 1.06 1.02 0.49 0.13 1.02 1.00 0.61 0.01 1.00 1.00 0.77 0.01 $y_{t-1}$ 1.30 1.20 0.49 0.15 1.08 1.21 0.61 0.01 1.08 1.24 0.77 0.01 $y_{t-1}^{(3)}$ 1.06 1.02 0.44 0.18 1.02 1.00 0.52 0.01 1.00 1.00 0.69 0.01 AR(1) 1.02 1.00 0.49 0.16 1.03 1.05 0.61 0.01 1.02 1.06 0.75 0.01 AR(2) 1.10 1.02 0.48 0.15 1.02 1.00 0.57 0.02 1.00 1.00 0.72 0.01 AR(3) 1.09 1.00 0.49 0.15 1.02 1.00 0.61 0.01 1.00 1.00 0.77 0.01 AIC 1.06 1.01 0.49 0.15 1.02 1.00 0.61 0.01 1.00 1.00 0.77 0.01 AICc 1.05 0.99 0.49 0.14 1.02 1.00 0.61 0.01 1.00 1.00 0.77 0.01 blCV 1.07 1.00 0.49 0.14 1.02 1.01 0.61 0.01 1.02 1.01 0.77 0.01 SNR = 10 $T$ 75 150 300 CoEr RMSFE % TP % FP CoEr RMSFE % TP % FP CoEr RMSFE % TP % FP ARMAr-LAS 1.05 1.03 0.87 0.05 1.03 1.01 0.97 0.02 1.00 1.00 1.00 0.01 $y_{t-1}$ 1.69 1.91 0.86 0.05 1.63 1.91 0.97 0.02 1.81 1.86 1.00 0.01 $y_{t-1}^{(3)}$ 1.05 1.03 0.65 0.09 1.03 1.01 0.80 0.01 1.00 1.00 0.93 0.01 AR(1) 1.13 1.04 0.85 0.06 1.17 1.12 0.96 0.02 1.19 1.10 1.00 0.01 AR(2) 1.05 1.01 0.84 0.07 1.03 1.01 0.95 0.03 1.00 1.00 0.99 0.02 AR(3) 1.07 1.01 0.86 0.05 1.03 1.01 0.97 0.02 1.00 1.00 1.00 0.01 AIC 1.06 1.02 0.86 0.05 1.03 1.02 0.97 0.02 1.00 1.00 1.00 0.01 AICc 1.05 1.02 0.85 0.05 1.03 1.02 0.97 0.02 1.00 1.00 1.00 0.01 blCV 1.07 1.04 0.85 0.05 1.10 1.09 0.97 0.02 1.15 1.07 1.00 0.01
Table 2: DGP (B). CoEr, RMSFE (relative to LAS), %TP and %FP for LASSO-based benchmarks and ARMAr-LASSO. For each $n$ setting the best CoEr and RMSFE are in bold.
$q=0$ $q=1$
SNR 1 10 1 10
n 75 150 300 75 150 300 75 150 300 75 150 300
CoEr
LASSOy 0.49 0.59 0.66 0.92 0.95 0.97 0.50 0.56 0.63 0.93 0.95 0.97
GLS-LAS 0.76 0.87 0.92 0.90 0.93 0.97 0.72 0.85 0.91 0.90 0.93 0.97
ARDL-LAS 0.45 0.54 0.61 0.81 0.89 0.91 0.46 0.52 0.58 0.80 0.87 0.91
FaSel 1.04 1.03 1.06 1.17 1.07 1.01 1.02 1.04 1.04 1.06 1.06 1.03
ARMAr-LAS 0.41 0.50 0.55 0.53 0.60 0.64 0.43 0.48 0.53 0.52 0.59 0.62
RMSFE
LASSOy 0.80 0.88 0.90 0.97 0.97 0.98 0.81 0.83 0.87 0.94 0.97 0.98
GLS-LAS 0.82 0.88 0.94 0.88 0.91 0.96 0.82 0.89 0.92 0.88 0.91 0.95
ARDL-LAS 0.80 0.89 0.90 1.00 0.99 0.97 0.78 0.80 0.85 0.92 0.93 0.98
FaSel 0.96 0.95 0.96 0.96 0.94 0.94 0.99 0.97 0.98 0.98 0.94 0.96
ARMAr-LAS 0.67 0.74 0.79 0.71 0.75 0.80 0.70 0.74 0.79 0.70 0.74 0.79
% TP
LASSO 57.10 52.00 52.60 87.60 87.00 87.50 48.10 39.40 33.40 82.10 79.30 76.60
LASSOy 46.50 43.80 44.10 87.20 86.50 87.40 31.90 27.60 22.90 81.70 78.40 76.00
GLS-LAS 55.30 51.10 51.30 89.10 86.30 86.20 43.30 37.90 32.40 83.30 80.70 77.30
ARDL-LAS 46.70 42.20 42.20 85.90 86.50 87.20 35.00 29.70 24.00 81.10 78.60 75.20
FaSel 47.10 45.60 49.70 73.60 81.80 87.10 45.60 37.90 32.10 78.60 79.00 78.20
ARMAr-LAS 65.60 62.40 60.80 97.20 96.60 96.40 54.80 50.00 43.40 96.40 95.30 94.30
% FP
LASSO 32.30 18.40 11.30 29.50 14.60 9.10 31.80 19.70 11.30 31.50 16.50 9.70
LASSOy 10.90 6.50 4.40 27.60 13.50 8.70 12.20 7.80 5.00 29.50 15.30 9.20
GLS-LAS 19.70 13.80 9.30 21.60 10.50 7.70 19.60 15.40 9.80 25.30 14.20 9.10
ARDL-LAS 4.10 2.30 1.70 8.60 5.70 3.60 5.20 3.10 1.80 8.70 5.40 3.20
FaSel 29.80 17.70 11.60 32.20 16.50 9.60 32.30 21.10 12.70 35.90 22.40 13.50
ARMAr-LAS 4.10 2.00 0.90 5.10 2.40 1.20 10.50 6.20 3.60 13.30 8.30 5.10
In addition to the results presented in this section, Supplement D reports further simulations based on a common AR(1) structure with varying autoregressive coefficients $\phi$, larger sample sizes $T$, and cases where the ARMAr-LASSO misspecifies the AR model of the predictors.
4.2 Empirical Application
We consider Euro Area (EA) data composed of 309 monthly macroeconomic time series spanning the period between January 1997 and December 2018. We note that the data used here are not real-time vintages. Rather, all macroeconomic series were downloaded at a single point in time and correspond to the most recent available releases as of the download date. The series are listed in Supplement E, grouped according to their measurement domain: Industry & Construction Survey (ICS), Consumer Confidence Indicators (CCI), Money & Interest Rates (M&IR), Industrial Production (IP), Harmonized Consumer Price Index (HCPI), Producer Price Index (PPI), Turnover & Retail Sale (TO), Harmonized Unemployment Rate (HUR), and Service Surveys (SI). Supplement E also reports transformations applied to the series to achieve stationarity (we did not attempt to identify or remove outliers), as well as an analysis of the autocorrelation functions that justifies the use of our ARMAr-LASSO in this context. The target variable is the Overall EA Consumer Price Index (CPI), which is transformed as I(2) (i.e. integration of order 2) following SeW2002b: $y_{t+h}=(1200/h)\text{log}(CPI_{t+h}/CPI_{t})-1200\ \text{log}(CPI_{t}/CPI_{t-1})$, where $y_{t}=1200\ \text{log}(CPI_{t}/CPI_{t-1})-1200\ \text{log}(CPI_{t-1}/CPI_{t-2})$, and $h$ is the forecasting horizon. We compute forecasts of $y_{t+h}$ at horizons $h=12$ and $24$ using a rolling $\omega$-year window $t-\omega,t+1]$; the models are re-estimated at each $t$, adding one observation on the right of the window and removing one observation on the left. The last forecast is December 2018. The methods employed for our empirical exercise are:
  • (a)
    Univariate AR($p$): the autoregressive forecasting model based on $p$ lagged values of the target variable, i.e. $\widehat{y}_{t+h}=\widehat{\alpha}_{0}+\sum_{i=1}^{p}\widehat{\phi}_{i}y_{t-i+1}$, which serves as a benchmark.
  • (b)
    LAS (tibshirani96): forecasts are obtained from the equation $\widehat{y}_{t+h}=\widehat{\alpha}_{0}+\sum_{l=0}^{11}\widehat{\phi}_{l}y_{t-l}+\sum_{i=1}^{308}\widehat{\alpha}_{i}x_{it},$ where $(\widehat{\phi}_{0},\dots,\widehat{\phi}_{11},\widehat{\alpha}_{1},\dots,\widehat{\alpha}_{308})$ is the sparse vector of penalized regression coefficients estimated by the LASSO.
  • (c)
    GLS-LAS (chronopoulos2023): forecasts are obtained from the equation $\widehat{y}_{t+h}=\widehat{\alpha}_{0}+\sum_{l=1}^{p_{\varepsilon}}\widehat{\phi}_{l}y_{t-l+1}+\sum_{i=1}^{308}\widehat{\alpha}_{i}\widetilde{x}_{it},$ where $(\widehat{\alpha}_{1},\dots,\widehat{\alpha}_{308})$ is the sparse vector of penalized regression coefficients estimated by the LASSO using pre-filtered response and predictors (the $\widetilde{x}$’s) as detailed in Supplement C.
  • (d)
    ARDL-LAS: forecasts are obtained from the equation $\widehat{y}_{t+h}=\widehat{\alpha}_{0}+\sum_{l=0}^{11}\widehat{\phi}_{l}y_{t-l}+\sum_{i=1}^{308}\sum_{j=0}^{2}\widehat{\alpha}_{i,t-j}x_{i,t-j},$ where $(\widehat{\phi}_{0},\dots,\widehat{\phi}_{11},\widehat{\alpha}_{1,t},\dots,\widehat{\alpha}_{308,t-2})$ is the sparse vector of penalized regression coefficients estimated by the LASSO, which in this case contains two lagged values for each predictor.
  • (e)
    FaSel (Fan2020): FarmSelector applies feature selection on factors residuals. Forecasts are obtained from the equation: $\widehat{y}_{t+h}=\widehat{\alpha}_{0}+\widehat{\boldsymbol{\Lambda}}\widehat{\boldsymbol{f}}_{t}+\widehat{\boldsymbol{\alpha}}^{\prime}\widehat{\mathbf{z}}_{t}+\sum_{i=1}^{p}\widehat{\phi}_{i}y_{t-i+1},$ where $\widehat{\boldsymbol{f}}_{t}$ is a $r$-dimensional vector of factors estimated with PCA (as in SeW2002a; SeW2002b), $\widehat{\mathbf{z}}_{t}=\widehat{\boldsymbol{\Lambda}}\widehat{\boldsymbol{f}}_{t}-\mathbf{x}_{t}$, $\widehat{\boldsymbol{\Lambda}}^{\prime}$ is the $n\times r$ matrix of loadings, and $\widehat{\boldsymbol{\alpha}}$ is the sparse vector obtained applying the LASSO. The number of factors $r$ is chosen with the approach described in AhnHorenstein2013.
  • (f)
    ARMAr-LAS: our proposal, where LASSO is applied to the estimated ARMA residuals. Forecasts are obtained from the equation $\widehat{y}_{t+h}=\widehat{\alpha}_{0}+\sum_{l=0}^{11}\widehat{\phi}_{l}y_{t-l}+\sum_{i=1}^{308}\widehat{\alpha}_{i}\widehat{u}_{it},$ where $(\widehat{\phi}_{0},\dots,\widehat{\phi}_{11},\widehat{\alpha}_{1},\dots,\widehat{\alpha}_{308})$ is the sparse vector of penalized regression coefficients estimated by the LASSO.
For the AR($p$) benchmark and the GLS-LAS, the lag orders $p$ and $p_{\varepsilon}$ are selected by BIC within $0\leq p,p_{\varepsilon}\leq 12$. For the ARMAr-LAS, estimated residuals (the $\widehat{u}$’s) are obtained filtering each time series with an ARMA($p_{i},q_{i}$), where $p_{i}$ and $q_{i}$ are selected by BIC within $0\leq p_{i},q_{i}\leq 12$, $i=1,\dots,n$. For all the LASSO-based methods (including our ARMAr-LAS), the shrinkage parameter $\lambda$ is also selected by BIC. Forecasting accuracy is evaluated using the root mean square forecast error (RMSFE), defined as $RMSFE=\sqrt{\frac{1}{T_{1}-T_{0}}\sum_{\tau=T_{0}}^{T_{1}}\Bigl(\widehat{y}_{\tau}-y_{\tau}\Bigr)^{2}},$ where $T_{0}$ and $T_{1}$ are the first and last time points used for the out-of-sample evaluation. We also consider the number of selected variables.
Table 3 reports ratios of RMSFEs between pairs of methods (RMSFE (ratio)), as well as significance of the corresponding Diebold-Mariano test (DiebMar1995). Furthermore, the column “Selected Variables (Av.)” reports the average number of selected variables with ARMAr-LAS (on the left side of the vertical bar), and other LASSO-based methods (on the right side of the vertical bar). Notably, ARMAr-LAS produces significantly better forecasts than AR($p$) and LASSO-based methods, and provides a more parsimonious model than the LAS, ARDL-LAS and FaSel. This is, in principle, consistent with the theoretical analysis we provided earlier. The sparser ARMAr-LAS output may be due to fewer false positives, as compared to other methods which suffer from the effects of spurious correlations induced by serial correlation. Notably, GLS-LAS selects fewer predictors than ARMAr-LAS but provides significantly worse predictions. However, since in this real data application we do not know the true DGP, any comment regarding accuracy in feature selection is necessarily speculative.
Table 3: RMSFE (ratio): ratios of RMSFE contrasting pairs of employed methods; for each ratio, we perform a Diebold-Mariano test (alternative: the second method is less accurate in forecasting) and report p-values as 0 ’***’ 0.001 ’**’ 0.01 ’*’ $^{\mathbin{\vbox{\hbox{\scalebox{0.5}{$\bullet$}}}}}’ 0$.1”. Selected Variables (Av.): average of the number of variables selected by ARMAr-LAS (left side of the vertical bar) and LASSO-based benchmarks (right side of the vertical bar).
Method 1 Method 2 RMSFE (ratio) Selected Variables (Av.)
hh=12 $h=24$ hh=12 $h=24$
ARMAr-LAS LAS 0.69*** 0.82* 6.0||67.9 6.2||60.9
ARMAr-LAS GLS-LAS 0.66*** 0.61*** 6.0||3.5 6.2||3.8
[…]
ARMAr-LAS FarSel 0.71*** 0.73*** 6.0||77.2 6.2||72.5
ARMAr-LAS AR(pp) 0.94 0.89*
LAS AR(pp) 1.36 1.08
GLS-LAS AR(pp) 1.43 1.44
ARDL-LAS AR(pp) 1.53 1.07
FarSel AR(pp) 1.32 1.21
Figure 4 summarizes patterns of selected predictors over time for LAS and ARMAr-LAS. The heatmaps represent the number of selected variables categorized according to the nine main domains (see above). LAS selects predictors largely, though not exclusively, from the domains ICS, M&IR and HUR. ARMAr-LAS is more targeted, selecting predictors almost exclusively in the HCPI domain (in Supplement F, we report the top 5 predictors in terms of selection frequency across forecasting samples). Note, however, that in a few instances (3 for $h=12$ and 2 for $h=24$) ARMAr-LAS does select many more predictors across multiple groups. Interestingly, these correspond to the period of the financial crisis (between the end of 2008 and the beginning of 2010), when negative shocks in some of the variables under consideration may indeed create a more complex picture in terms of feature selection. In summary, ARMAr-LAS exploits cross-sectional information mainly focusing on prices, and accrues a forecasting advantage – as LAS uses many more variables to produce significantly worse forecasts.
Figure 4: Heatmaps representing the number of variables selected by LAS (left) and ARMAr-LAS (right) in the nine main domains. The tuning procedure is BIC.
5 Concluding Remarks
In this paper, we demonstrated that the probability of spurious correlations between stationary orthogonal or weakly correlated processes depends not only on the sample size, but also on the degree of predictors serial correlation. Through this result, we pointed out that serial correlation negatively affects the estimation and forecasting error bounds of LASSO. In order to improve the performance of LASSO in a time series context, we proposed an approach based on applying LASSO to pre-whitened (i.e., ARMA filtered) time series. This proposal relies on a working model that mitigates large spurious correlation and improves both estimation and forecast accuracy. We characterized limiting distribution and feature selection consistency, as well as forecast and estimation error bounds, for our proposal. Furthermore, we assessed its performance through Monte Carlo simulations and an empirical application to Euro Area macroeconomic time series. Through simulations, we observed that ARMAr-LASSO, i.e., LASSO applied to ARMA residuals, reduces the probability of large spurious correlations and outperforms other LASSO-based methods from the literature in terms of both coefficient estimation and forecasting. The empirical application confirms that ARMAr-LASSO improves forecasting performance and produces more parsimonious models.
Based on the results obtained so far, we envision several avenues for future work. For instance, it would be of interest to derive the rate at which the distribution of the sample correlation coefficient approaches $\mathcal{D}(r)$, thus formalizing what we observed numerically in Figure 2. Another promising avenue for future work is the development of a desparsified ARMAr-LASSO to enable valid inference in high-dimensional time series with serially correlated predictors and error terms. Such an extension could build on the Bartlett-kernel Newey-West long run covariance estimator and could be compared with the recent inferential frameworks proposed by Chernozhukov2021 and Babii2024. We also note that the density in Proposition 1 provides a theoretical foundation for further advancement in testing correlations that link autoregressive processes.
Finally, we intend to explore additional econometric applications; for instance, the analysis of EA macroeconomic data presented here could be replicated on other data sets, such as the FRED-MD dataset for the U.S.
Appendix A Proofs
In this section, we provide the proof of our theoretical results.
A.1 Proof of Proposition 1
Let $\mathbf{x}_{t}=\boldsymbol{\phi}\mathbf{x}_{t-1}+\mathbf{u}_{t}$, $t=1,\dots,T$, be a first order $n$-variate autoregressive process as in Section 2.1. We focus on the probability density of $\widehat{c}_{ij}^{x}$. Following Anderson03 ch. 4, we shall consider $r=\frac{a_{ij}}{\sqrt{a_{ii}}\sqrt{a_{jj}}}$, where $a_{ij}=\sum_{t=1}^{T}(x_{i,t}-\overline{x}_{i})(x_{j,t}-\overline{x}_{j})$. In particular, when $c_{ij}^{u}=0$, $b=a_{ji}/a_{ii}$ and $v=a_{jj}-a_{ji}^{2}/a_{ii}$,
$\cfrac{\sqrt{a_{ii}}\ b}{\sqrt{v}}=\cfrac{a_{ij}/\sqrt{a_{ii}a_{jj}}}{\sqrt{1-a_{ij}^{2}/(a_{ii}a_{jj})}}=\cfrac{r}{\sqrt{1-r^{2}}}\ \ .$ (1)
Note that $b$ is the least squares regression coefficient of $x_{jt}$ on $x_{it}$, and $v$ is the sum of the square of residuals of such regression. Thus, according to (1), to obtain the probability density of $\widehat{c}_{ij}^{x}$, we need the distributions of $b$ and $v$.
Remark 1:
In contrast to asymptotic statements, our theoretical analysis is intended to derive distributions and densities of estimators that hold for finite $T$ as in Proposition 1. Hence, we will not employ the usual concepts of convergence in probability and in distribution; rather, we will use a notion of approximation, whose “precision” has been numerically evaluated in Section 2.1 ans Supplement B.
Throughout, the symbol $\overset{\mathrm{BE}}{\approx}$ denotes a finite–sample Gaussian approximation justified by Berry–Esseen type bounds and assessed via Monte Carlo simulations; it is not an asymptotic statement. The symbol $\approx$ is used exclusively to indicate a classical numerical approximation and does not refer to any asymptotic notion of convergence.
Distribution of $b$. We start by deriving the sample distribution of $b$, the OLS regression coefficient for $x_{j}$ on $x_{i}$. The same holds if we regress $x_{i}$ on $x_{j}$.
Lemma 1:
The sample probability distribution of $b$ is approximately $N\mathopen{}\mathclose{{\left(0,\frac{(1-\phi_{i}^{2}\phi_{j}^{2})(1-\phi_{i}^{2})}{(T-1)(1-\phi_{j}^{2})(1-\phi_{i}\phi_{j})^{2}}}}\right)$.
Proof of Lemma 1 We first focus on the distribution of the sample covariance between $x_{i,t}$ and $x_{j,t}$. Let $\widehat{Cov}(x_{i[-l]},x_{j})=\sum_{t=l+1}^{T}(x_{i,t-l}-\overline{x}_{i[-l]})(x_{j,t}-\overline{x}_{j})/(T-l-1)$, for $i\neq j$, where $\overline{x}_{i[-l]}=\frac{1}{T-l}\sum_{t=l+1}^{T}x_{i,t-l}$ and $\overline{x}_{j}=\frac{1}{T-l}\sum_{t=l+1}^{T}x_{j,t}$. By considering the AR(1) decomposition of $x_{i,t}$ and $x_{j,t}$, we have
$\frac{a_{ij}}{(T-1)}=\widehat{Cov}(x_{i},x_{j})\\ =\phi_{i}\phi_{j}\widehat{Cov}(x_{i[-1]},x_{j[-1]})+\phi_{i}\widehat{Cov}(x_{i[-1]},u_{j})+\phi_{j}\widehat{Cov}(x_{j[-1]},u_{i})+\widehat{Cov}(u_{i},u_{j})\ \ .$ (2)
Note that by moving $\phi_{i}\phi_{j}\widehat{Cov}(x_{i[-1]},x_{j[-1]})$ on the left side of the equality and adding and removing the quantity $\phi_{i}\phi_{j}\widehat{Cov}(x_{i},x_{j})$, after few algebra we obtain
$(1-\phi_{i}\phi_{j})\widehat{Cov}(x_{i},x_{j})=\phi_{i}\widehat{Cov}(x_{i[-1]},u_{j})+\phi_{j}\widehat{Cov}(x_{j,[-1]},u_{i})+\widehat{Cov}(u_{i},u_{j})\\ -\phi_{i}\phi_{j}\mathopen{}\mathclose{{\left(\widehat{Cov}(x_{i},x_{j})-\widehat{Cov}(x_{i[-1]},x_{j[-1]})}}\right)\ \ .$ (3)
Note that,
$\phi_{i}\widehat{Cov}(x_{i[-1]},u_{j})=\sum_{l=1}^{T-3}\phi_{i}^{l}\widehat{Cov}(u_{i[-l]},u_{j})+\phi_{i}^{T-2}\widehat{Cov}(x_{i[-(T-2)]},u_{j}).$
The remainder terms $\phi_{i}^{T-2}\widehat{Cov}(x_{i[-(T-2)]},u_{j})$, $\phi_{j}^{T-2}\widehat{Cov}(x_{j[-(T-2)]},u_{i})$, and $\phi_{i}\phi_{j}\mathopen{}\mathclose{{\left(\widehat{Cov}(x_{i},x_{j})-\widehat{Cov}(x_{i[-1]},x_{j[-1]})}}\right)$ are negligible, and we may write
$\widehat{Cov}(x_{i},x_{j})\approx\\ \mathopen{}\mathclose{{\left[\sum_{l=1}^{T-3}\phi_{i}^{l}\widehat{Cov}(u_{i[-l]},u_{j})+\sum_{l=1}^{T-3}\phi_{j}^{l}\widehat{Cov}(u_{j[-l]},u_{i})+\widehat{Cov}(u_{i},u_{j})}}\right](1-\phi_{i}\phi_{j})^{-1}\ \ .$ (4)
We can write $\widehat{Cov}(u_{i},u_{j})=\frac{1}{T-1}\sum_{t=1}^{T}u_{i,t}u_{j,t}-\frac{T}{T-1}\overline{u}_{i}\overline{u}_{j}$, where $\frac{T}{T-1}\overline{u}_{i}\overline{u}_{j}$ is negligible. Let $W_{t}\coloneq u_{i,t}u_{j,t}$, and $S_{T}\coloneq\frac{1}{\sqrt{T-1}}\sum_{t=1}^{T}W_{t}$. By the Berry–Esseen theorem berry1941; esseen1942, we have
$\sup_{a}\mathopen{}\mathclose{{\left|P\!\mathopen{}\mathclose{{\left(S_{T}\leq a}}\right)-\Phi(a)}}\right|\;\leq\;\frac{C\,E(|W_{1}|^{3})}{\sqrt{T-1}}\ \ ,$
for some universal constant $C<0.5$, and where $\Phi$ is the cumulative distribution function of the standard normal distribution. Moreover, $E(|W_{1}|^{3})=E(|u_{i,1}u_{j,1}|^{3})=E(|u_{i,1}|^{3})^{2}=E(|u_{j,1}|^{3})^{2}=(2\sqrt{2/\pi})^{2}=8/\pi\approx 2.546$. Thus, $\frac{C\,E(|W_{1}|^{3})}{\sqrt{T-1}}\;<\frac{0.5\times 2.55}{\sqrt{T-1}}\;=\;\frac{1.275}{\sqrt{T-1}}.$ Then, the Berry-Esseen theorem guarantees that
$\widehat{Cov}(u_{i,t},u_{j,t})\overset{\mathrm{BE}}{\approx}N\mathopen{}\mathclose{{\left(0\>,\>\cfrac{1}{(T-1)}}}\right)\ \ .$
For a sample size $T>20$, the Berry-Esseen bound guarantees that $\frac{1.275}{\sqrt{T-1}}\;<\;0.292$. See Supplement G for numerical results.
Moreover, define
$\eta_{ij}=\sum_{l=1}^{T-3}\phi_{i}^{l}\widehat{Cov}(u_{i,t-l},u_{j,t})+\sum_{l=1}^{T-3}\phi_{j}^{l}\widehat{Cov}(u_{j,t-l},u_{i,t}).$ (5)
The quantity $\eta_{ij}$ is a linear combination of sample cross–covariances between the innovations of one series and lagged innovations of the other series. For each fixed lag $l$, by the Berry–Esseen theorem,
$\widehat{Cov}(u_{i,t-l},u_{j,t})\overset{\mathrm{BE}}{\approx}N\!\mathopen{}\mathclose{{\left(0,\frac{1}{T-l-1}}}\right),$
and analogously for $\widehat{Cov}(u_{j,t-l},u_{i,t})$. Although the sample covariances at different lags are not independent, their mutual covariances are of smaller order with respect to the leading variance terms and are therefore negligible in the present approximation. Hence, exploiting $|\phi_{i}|,|\phi_{j}|<1$ and the convergence of the associated geometric series, we obtain
$\mathrm{Var}(\eta_{ij})\approx\sum_{l=1}^{T-3}\frac{\phi_{i}^{2l}}{T-l-1}+\sum_{l=1}^{T-3}\frac{\phi_{j}^{2l}}{T-l-1}\approx\frac{1}{T-1}\mathopen{}\mathclose{{\left(\frac{\phi_{i}^{2}}{1-\phi_{i}^{2}}+\frac{\phi_{j}^{2}}{1-\phi_{j}^{2}}}}\right).$
The approximation above treats the lag–specific sample cross–covariances as effectively uncorrelated when computing $\mathrm{Var}(\eta_{ij})$. Under temporally independent Gaussian innovations, the remaining cross–lag covariance terms arise only from finite–sample index overlap and contribute at a smaller order (typically $O(T^{-1})$) relative to the leading variance terms; we therefore neglect them as a second–order effect. Therefore,
$\eta_{ij}\overset{\mathrm{BE}}{\approx}N\mathopen{}\mathclose{{\left(0,\frac{\phi_{i}^{2}+\phi_{j}^{2}-2\phi_{i}^{2}\phi_{j}^{2}}{(T-1)(1-\phi_{i}^{2})(1-\phi_{j}^{2})}}}\right)\ \ ,$
and
$\widehat{Cov}(x_{i,t},x_{j,t})\overset{\mathrm{BE}}{\approx}N\mathopen{}\mathclose{{\left(0\>,\>\frac{1-\phi_{i}^{2}\phi_{j}^{2}}{(T-1)(1-\phi_{i}^{2})(1-\phi_{j}^{2})(1-\phi_{i}\phi_{j})^{2}}}}\right)\ \ .$
To obtain a closed–form representation of the distribution of $b=a_{ji}/a_{ii}$, we adopt a plug–in approximation by replacing $a_{ii}$ with its expectation $E[a_{ii}]=(T-1)/(1-\phi_{i}^{2})$. Following standard approximations for ratios of random variables (see Stuart1998), we have
$b\overset{\mathrm{BE}}{\approx}N\mathopen{}\mathclose{{\left(0,\frac{(1-\phi_{i}^{2}\phi_{j}^{2})(1-\phi_{i}^{2})}{(T-1)(1-\phi_{j}^{2})(1-\phi_{i}\phi_{j})^{2}}}}\right)\ \ .$
$\square$
Distribution of $v$. Here, we derive the sample distribution of the sum of the square of residuals obtained by regressing $x_{j}$ on $x_{i}$. Since $v=a_{jj}-a_{ji}^{2}/a_{ii}$, we start by deriving the distribution of $a_{jj}$ and $a_{ji}^{2}/a_{ii}$ in the following two Lemmas.
Lemma 2:
The sample probability distribution of $a_{jj}$ is approximately $\Gamma\mathopen{}\mathclose{{\left(\frac{(T-1)^{2}}{\xi_{a}},\frac{\xi_{a}}{(T-1)(1-\phi_{j}^{2})}}}\right)$, where the quantity $\xi_{a}=\mathopen{}\mathclose{{\left[3(T-1)-(T-1)^{2}+2\sum_{t=1}^{T-2}(T-1-t)(1+2\phi_{j}^{2t})}}\right]$.
Proof of Lemma 2 Let $z_{j,t}$ be the standardized version of $x_{j,t}$, so that $x_{j,t}=z_{j,t}/\sqrt{1-\phi_{j}^{2}}$. Then $a_{jj}=\sum_{t=1}^{T}(x_{j,t}-\bar{x}_{j})^{2}\;\approx\;\sum_{t=1}^{T}x_{j,t}^{2}=\frac{1}{1-\phi_{j}^{2}}\sum_{t=1}^{T}z_{j,t}^{2}$. Using the T1T-1 normalization for sample variances and covariances, we approximate the above quadratic form by a sum of T1T-1 correlated $\chi^{2}_{1}$ terms, that is, we work with an effective number of degrees of freedom equal to T1T-1. Thus, $a_{jj}$ is approximated by a Gamma distribution with shape parameter $k_{a}$ and scale parameter $\theta_{a}$, obtained by matching the first two moments. Thus, $a_{jj}$ is the sum of T1T-1 correlated $\chi_{1}^{2}$ multiplied by $\frac{1}{1-\phi_{j}^{2}}$, approximate a Gamma distribution with shape parameter $k_{a}$ and a scale parameter $\theta_{a}$. Thus, we have to define such parameters via moments matching. We have $E(a_{jj})=\frac{T-1}{1-\phi_{j}^{2}}$ and, consequently to the dependency between the elements of $a_{jj}$, $Var(a_{jj})=\xi_{a}(1-\phi_{j}^{2})^{-2}$, where $\xi_{a}=\mathopen{}\mathclose{{\left[3(T-1)-(T-1)^{2}+2\sum_{t=1}^{T-2}(T-1-t)(1+2\phi_{j}^{2t})}}\right]$. We can use these moments to obtain $k_{a}=\frac{E(a_{jj})^{2}}{Var(a_{jj})}=\frac{(T-1)^{2}}{\xi_{a}}$ and $\theta_{a}=\frac{Var(a_{jj})}{E(a_{jj})}=\frac{\xi_{a}}{(T-1)(1-\phi_{j}^{2})}$. Therefore $a_{jj}\approx\Gamma\mathopen{}\mathclose{{\left(\frac{(T-1)^{2}}{\xi_{a}},\frac{\xi_{a}}{(T-1)(1-\phi_{j}^{2})}}}\right)$. $\square$
Lemma 3:
The sample probability distribution of $a_{ij}^{2}/a_{ii}$ is approximately $\Gamma\mathopen{}\mathclose{{\left(\frac{1}{2},\frac{2(1-\phi_{i}^{2}\phi_{j}^{2})}{(1-\phi_{j}^{2})(1-\phi_{i}\phi_{j})^{2}}}}\right)$.
Proof of Lemma 3 Note that $a_{ij}/\sqrt{a_{ii}}=\sqrt{a_{ii}}b$. Thus, by Lemma 1 we have that $a_{ij}/\sqrt{a_{ii}}=\sqrt{a_{ii}}b\approx N\mathopen{}\mathclose{{\left(0,\frac{(1-\phi_{i}^{2}\phi_{j}^{2})}{(1-\phi_{j}^{2})(1-\phi_{i}\phi_{j})^{2}}}}\right)$. Let $z$ be the variable obtained by standardizing $a_{ij}/\sqrt{a_{ii}}$, we have $a_{ij}^{2}/a_{ii}=\frac{z^{2}(1-\phi_{i}^{2}\phi_{j}^{2})}{(1-\phi_{j}^{2})(1-\phi_{i}\phi_{j})^{2}}$ where $E(a_{ij}^{2}/a_{ii})=\frac{(1-\phi_{i}^{2}\phi_{j}^{2})}{(1-\phi_{j}^{2})(1-\phi_{i}\phi_{j})^{2}}$ and $Var(a_{ij}^{2}/a_{ii})=2\mathopen{}\mathclose{{\left(\frac{(1-\phi_{i}^{2}\phi_{j}^{2})}{(1-\phi_{j}^{2})(1-\phi_{i}\phi_{j})^{2}}}}\right)^{2}$. Using the same argument as in Lemma 2, we obtain $a_{ij}^{2}/a_{ii}\approx\Gamma\mathopen{}\mathclose{{\left(\frac{1}{2},\frac{2(1-\phi_{i}^{2}\phi_{j}^{2})}{(1-\phi_{j}^{2})(1-\phi_{i}\phi_{j})^{2}}}}\right)$. $\square$
Lemmas 2 and 3 allow us to derive the sample distribution of $v$.
Lemma 4:
The sample probability distribution of $v=a_{jj}-a_{ji}^{2}/a_{ii}$ is approximately $\Gamma\mathopen{}\mathclose{{\left(\frac{T_{v}^{2}}{\xi_{v}},\frac{\xi_{v}}{T_{v}(1-\phi_{j}^{2})}}}\right)$, where the quantities $T_{v}=\mathopen{}\mathclose{{\left\lfloor\frac{(T-1)(1-\phi_{i}\phi_{j})^{2}-(1-\phi_{i}^{2}\phi_{j}^{2})}{(1-\phi_{i}\phi_{j})^{2}}}}\right\rceil$, and $\xi_{v}=\mathopen{}\mathclose{{\left[3T_{v}-T_{v}^{2}+2\sum_{t=1}^{T_{v}-1}(T_{v}-t)(1+2\phi_{j}^{2t})}}\right]$.
Proof of Lemma 4 While Lemmas 23 provide marginal approximations for $a_{jj}$ and $a_{ji}^{2}/a_{ii}$, the exact finite–sample law of $v=a_{jj}-a_{ji}^{2}/a_{ii}$ does not admit a tractable closed-form expression, since it depends on the joint distribution of these two terms and, in particular, on ${Cov}\!\big(a_{jj},a_{ji}^{2}/a_{ii}\big)$. We therefore introduce an additional moment–matching approximation: we model $v$ as a scaled quadratic form with the same dependence structure as $x_{j}$, but with an effective number of degrees of freedom $T_{v}$ chosen to match $E(v)$. The corresponding variance is approximated by that of a sum of $T_{v}$ correlated $\chi^{2}_{1}$ components (with AR(1) dependence parameter $\phi_{j}$), yielding the Gamma law in Lemma 4.
We combine the results in Lemmas 2 and 3. Considering $E(v)=E\!\mathopen{}\mathclose{{\left(a_{jj}-\frac{a_{ij}^{2}}{a_{ii}}}}\right)=E(a_{jj})-E\!\mathopen{}\mathclose{{\left(\frac{a_{ij}^{2}}{a_{ii}}}}\right)=\frac{(T-1)(1-\phi_{i}\phi_{j})^{2}-(1-\phi_{i}^{2}\phi_{j}^{2})}{(1-\phi_{j}^{2})(1-\phi_{i}\phi_{j})^{2}},$ we define $T_{v}=\mathopen{}\mathclose{{\left\lfloor\frac{(T-1)(1-\phi_{i}\phi_{j})^{2}-(1-\phi_{i}^{2}\phi_{j}^{2})}{(1-\phi_{i}\phi_{j})^{2}}}}\right\rceil$. Therefore, using the same moment–matching argument as in Lemma 2, we approximate $v$ by a sum of $T_{v}$ correlated $\chi^{2}_{1}$ terms with the same AR(1) dependence structure as $x_{j}$. Under this effective degrees–of–freedom approximation, we approximate $\mathrm{Var}(v)$ by the variance of a sum of $T_{v}$ correlated $\chi^{2}_{1}$ terms with AR(1) dependence parameter $\phi_{j}$, yielding $\mathrm{Var}(v)\approx\frac{\xi_{v}}{(1-\phi_{j}^{2})^{2}},$ where $\xi_{v}=\mathopen{}\mathclose{{\left[3T_{v}-T_{v}^{2}+2\sum_{t=1}^{T_{v}-1}(T_{v}-t)(1+2\phi_{j}^{2t})}}\right]$ incorporates the serial dependence among the $\{x_{j,t}\}$. Matching the first two moments then leads to the Gamma approximation $v=a_{jj}-\frac{a_{ji}^{2}}{a_{ii}}\approx\Gamma\!\mathopen{}\mathclose{{\left(\frac{T_{v}^{2}}{\xi_{v}},\frac{\xi_{v}}{T_{v}(1-\phi_{j}^{2})}}}\right).$ $\square$
Note that the distribution of $b$ and $v$ in the case of independent observations (i.e., without serial correlation) are known (see Anderson03, ch, 4). Here, Lemmas 1 and 4 derive the sample distributions of $b$ and $v$ in the case of serial correlation, namely, by taking into account the dependence of the vectors $\boldsymbol{x}_{i}$ and $\boldsymbol{x}_{j}$.
Proof of Proposition 1 Although $\mathbf{x}_{j}$ is Gaussian, its temporal dependence implies that, under OLS, $\sqrt{a_{ii}}\,b$ and $v$ are not exactly independent in finite samples. However, in line with our finite–sample approximation framework, we treat their dependence as a second–order effect and approximate them as independent. Using Lemmas 1 and 4 and equation (1) we can now derive the probability density of $\widehat{c}_{ij}^{x}$. Because of Lemma 1, $\sqrt{a_{ii}}$ is approximately $N\mathopen{}\mathclose{{\left(0,\frac{1-\phi_{i}^{2}\phi_{j}^{2}}{(1-\phi_{j}^{2})(1-\phi_{i}\phi_{j})^{2}}}}\right)$. Let $\delta^{2}=\frac{1-\phi_{i}^{2}\phi_{j}^{2}}{(1-\phi_{j}^{2})(1-\phi_{i}\phi_{j})^{2}}$, $k_{v}=\frac{T_{v}^{2}}{\xi_{v}}$, $\theta_{v}=\frac{\xi_{v}}{T_{v}(1-\phi_{j}^{2})}$ and $t=\frac{\sqrt{a_{ii}}b}{\sqrt{v}}$. In the remainder of the proof, we consider the distributions of $\sqrt{a_{ii}}$ and $v$ in Lemmas 1 and 4 as exact. Thus, we have the densities
$g\mathopen{}\mathclose{{\left(\sqrt{a_{ii}}b}}\right)$ $=$ $\frac{1}{\delta\sqrt{2\pi}}\text{exp}\mathopen{}\mathclose{{\left(-\frac{a_{ii}b^{2}}{2\delta^{2}}}}\right)\ ,\ \,\ \ h(v)=\frac{1}{(\theta_{v})^{k_{v}}\Gamma\mathopen{}\mathclose{{\left(k_{v}}}\right)}v^{k_{v}-1}\text{exp}\mathopen{}\mathclose{{\left(-\frac{v}{\theta_{v}}}}\right)\ .$
We focus on
$f(t)$ $=$ $\int\sqrt{v}g\mathopen{}\mathclose{{\left(\sqrt{v}t}}\right)h(v)dv=\int_{0}^{\infty}\sqrt{v}\frac{1}{\delta\sqrt{2\pi}}\text{exp}\mathopen{}\mathclose{{\left(-\frac{vt^{2}}{2\delta^{2}}}}\right)\frac{v^{k_{v}-1}\text{exp}\mathopen{}\mathclose{{\left(-\frac{v}{\theta_{v}}}}\right)}{(\theta_{v})^{k_{v}}\Gamma\mathopen{}\mathclose{{\left(k_{v}}}\right)}dv$
$=$ $\frac{1}{\sqrt{2\pi}\delta(\theta_{v})^{k_{v}}\Gamma\mathopen{}\mathclose{{\left(k_{v}}}\right)}\int_{0}^{\infty}v^{k_{v}-\frac{1}{2}}\text{exp}\mathopen{}\mathclose{{\left(-\mathopen{}\mathclose{{\left(\frac{1}{\theta_{v}}+\frac{t^{2}}{2\delta^{2}}}}\right)v}}\right)dv\ .$
Now define $\Upsilon=\frac{1}{\sqrt{2\pi}\delta(\theta_{v})^{k_{v}}\Gamma\mathopen{}\mathclose{{\left(k_{v}}}\right)}$ and $x=\mathopen{}\mathclose{{\left(\frac{1}{\theta_{v}}+\frac{t^{2}}{2\delta^{2}}}}\right)v$. Then
$f(t)$ $=$ $\Upsilon\int_{0}^{\infty}\mathopen{}\mathclose{{\left[x\mathopen{}\mathclose{{\left(\frac{1}{\theta_{v}}+\frac{t^{2}}{2\delta^{2}}}}\right)^{-1}}}\right]^{k_{v}-\frac{1}{2}}\text{exp}\mathopen{}\mathclose{{\left(-x}}\right)dx$
$=$ $\Upsilon\ \mathopen{}\mathclose{{\left(\frac{1}{\theta_{v}}+\frac{t^{2}}{2\delta^{2}}}}\right)^{-(k_{v}+\frac{1}{2})}\int_{0}^{\infty}x^{k_{v}+\frac{1}{2}-1}\text{exp}\mathopen{}\mathclose{{\left(-x}}\right)dx\ .$
The integral on the right-hand side can be represented by using the gamma function $\Gamma(\alpha)=\int_{0}^{\infty}x^{\alpha-1}\text{exp}\mathopen{}\mathclose{{\left(-x}}\right)dx$. Thus we obtain
$f(t)$ $=$ $\Upsilon\ \mathopen{}\mathclose{{\left(\frac{1}{\theta_{v}}+\frac{t^{2}}{2\delta^{2}}}}\right)^{-(k_{v}+\frac{1}{2})}\Gamma\mathopen{}\mathclose{{\left(k_{v}+\frac{1}{2}}}\right)=\Upsilon\ \mathopen{}\mathclose{{\left(\frac{2\delta^{2}+\theta_{v}t^{2}}{\theta_{v}2\delta^{2}}}}\right)^{-(k_{v}+\frac{1}{2})}\Gamma\mathopen{}\mathclose{{\left(k_{v}+\frac{1}{2}}}\right)$
$=$ $\frac{\Gamma\mathopen{}\mathclose{{\left(k_{v}+\frac{1}{2}}}\right)}{\sqrt{2\pi}\delta(\theta_{v})^{k_{v}}\Gamma\mathopen{}\mathclose{{\left(k_{v}}}\right)}\mathopen{}\mathclose{{\left(\frac{2\delta^{2}+\theta_{v}t^{2}}{\theta_{v}2\delta^{2}}}}\right)^{-(k_{v}+\frac{1}{2})}=\frac{\Gamma\mathopen{}\mathclose{{\left(k_{v}+\frac{1}{2}}}\right)\sqrt{\theta_{v}}}{\delta\sqrt{2\pi}\Gamma\mathopen{}\mathclose{{\left(k_{v}}}\right)}\mathopen{}\mathclose{{\left[\frac{2\delta^{2}+\theta_{v}t^{2}}{2\delta^{2}}}}\right]^{-(k_{v}+\frac{1}{2})}.$
Substituting $\delta^{2}$ with $\frac{1-\phi_{i}^{2}\phi_{j}^{2}}{(1-\phi_{j}^{2})(1-\phi_{i}\phi_{j})^{2}}$ and $\theta_{v}$ with $\frac{\xi_{v}}{T_{v}(1-\phi_{j}^{2})}$, we obtain the density
$f(t)$ $=$ $\frac{\Gamma\mathopen{}\mathclose{{\left(k_{v}+\frac{1}{2}}}\right)\sqrt{\xi_{v}(1-\phi_{j}^{2})(1-\phi_{i}\phi_{j})^{2}}}{\Gamma\mathopen{}\mathclose{{\left(k_{v}}}\right)\sqrt{2\pi T_{v}(1-\phi_{i}^{2}\phi_{j}^{2})(1-\phi_{j}^{2})}}\mathopen{}\mathclose{{\left(1+\frac{t^{2}\xi_{v}(1-\phi_{i}\phi_{j})^{2}(1-\phi_{j}^{2})}{2T_{v}(1-\phi_{i}^{2}\phi_{j}^{2})(1-\phi_{j}^{2})}}}\right)^{-(k_{v}+\frac{1}{2})}$
$=$ $\frac{\Gamma\mathopen{}\mathclose{{\left(k_{v}+\frac{1}{2}}}\right)(1-\phi_{i}\phi_{j})\sqrt{\xi_{v}}}{\Gamma\mathopen{}\mathclose{{\left(k_{v}}}\right)\sqrt{2\pi T_{v}(1-\phi_{i}^{2}\phi_{j}^{2})}}\mathopen{}\mathclose{{\left(1+\frac{t^{2}\xi_{v}(1-\phi_{i}\phi_{j})^{2}}{2T_{v}(1-\phi_{i}^{2}\phi_{j}^{2})}}}\right)^{-(k_{v}+\frac{1}{2})}\ .$
The density of $w=r\mathopen{}\mathclose{{\left[1-r^{2}}}\right]^{-\frac{1}{2}}$, where $r\in[-1,1]$, is thus
$f(w)$ $=$ $\frac{\Gamma\mathopen{}\mathclose{{\left(k_{v}+\frac{1}{2}}}\right)(1-\phi_{i}\phi_{j})\sqrt{\xi_{v}}}{\Gamma\mathopen{}\mathclose{{\left(k_{v}}}\right)\sqrt{2\pi T_{v}(1-\phi_{i}^{2}\phi_{j}^{2})}}\mathopen{}\mathclose{{\left[1+\frac{w^{2}\xi_{v}(1-\phi_{i}\phi_{j})^{2}}{2T_{v}(1-\phi_{i}^{2}\phi_{j}^{2})}}}\right]^{-(k_{v}+\frac{1}{2})}\ .$
Next, define $\kappa(r)=w=r\mathopen{}\mathclose{{\left[1-r^{2}}}\right]^{-\frac{1}{2}}$, from which $\kappa^{\prime}(r)=\mathopen{}\mathclose{{\left[1-r^{2}}}\right]^{-\frac{3}{2}}$, $\ddot{\phi}=\phi_{i}\phi_{j}$ and $\Theta=\frac{\Gamma\mathopen{}\mathclose{{\left(k_{v}+\frac{1}{2}}}\right)(1-\ddot{\phi})\sqrt{\xi_{v}}}{\Gamma\mathopen{}\mathclose{{\left(k_{v}}}\right)\sqrt{2\pi T_{v}(1-\ddot{\phi}^{2})}}$. We can use these quantities to write
$\mathcal{D}(r)$ $=$ $f_{w}(\kappa(r))\kappa^{\prime}(r)=\Theta\mathopen{}\mathclose{{\left[1+\mathopen{}\mathclose{{\left(r(1-r^{2})^{-\frac{1}{2}}}}\right)^{2}\frac{\xi_{v}(1-\ddot{\phi})^{2}}{2T_{v}(1-\ddot{\phi}^{2})}}}\right]^{-(k_{v}+\frac{1}{2})}\mathopen{}\mathclose{{\left[1-r^{2}}}\right]^{-\frac{3}{2}}$
$=$ $\Theta\mathopen{}\mathclose{{\left[1-r^{2}}}\right]^{k_{v}-1}\mathopen{}\mathclose{{\left[\frac{2T_{v}(1-\ddot{\phi}^{2})}{(1-r^{2})2T_{v}(1-\ddot{\phi}^{2})+r^{2}\xi_{v}(1-\ddot{\phi})^{2}}}}\right]^{k_{v}+\frac{1}{2}}\ .$
Thus, the (finite) sample probability density of $\widehat{c}_{ij}^{x}$, tacking the densities in Lemmas 1 and 4 as exact, is
$\mathcal{D}(r)=\frac{\Gamma\mathopen{}\mathclose{{\left(k_{v}+\frac{1}{2}}}\right)(1-\ddot{\phi})\sqrt{\xi_{v}}}{\Gamma\mathopen{}\mathclose{{\left(k_{v}}}\right)\sqrt{\pi}}\frac{\mathopen{}\mathclose{{\left[1-r^{2}}}\right]^{k_{v}-1}\mathopen{}\mathclose{{\left[2T_{v}(1-\ddot{\phi}^{2})}}\right]^{k_{v}}}{\mathopen{}\mathclose{{\left[(1-r^{2})2T_{v}(1-\ddot{\phi}^{2})+r^{2}\xi_{v}(1-\ddot{\phi})^{2}}}\right]^{k_{v}+\frac{1}{2}}}\ \ ,\ \ r\in[-1,1]\ \ .$
$\square$
A.2 Proof of Proposition 2
By lemma 1 and Hastie2015, ch. 11 , for any $i\in\{1,\dots,n\}$ we have that for $T$ sufficiently large $\mathbf{x}_{i}^{\prime}\boldsymbol{\varepsilon}/T$ is stochastically dominated by a $N\mathopen{}\mathclose{{\left(0,\sigma_{x\varepsilon}^{2}/T}}\right)$, where $\sigma_{x\varepsilon}^{2}=\frac{1-\phi^{2}\phi_{\varepsilon}^{2}}{(1-\phi_{\varepsilon}^{2})(1-\phi\phi_{\varepsilon})^{2}}$. Thus, we have that
$P\mathopen{}\mathclose{{\left(\frac{|\mathbf{x}^{\prime}\boldsymbol{\varepsilon}|}{T}\geq t}}\right)\leq 2e^{-\frac{Tt^{2}}{2\sigma_{x\varepsilon}^{2}}},$
and the union bound yields
[…]
where the second equality follows by setting $t=\sigma_{x\varepsilon}\sqrt{\frac{c_{0}log(n)}{T}}$ for some $c_{0}>2$. Consequently, the inequalities hold with probability at least $1-2e^{-\frac{1}{2}(c_{0}-2)log(n)}$ by setting $\ddot{\lambda}=2\sigma_{x\varepsilon}\sqrt{\frac{c_{0}log(n)}{T}}$, for some $c_{0}>2$. $\square$
A.3 Proof of Theorem 1
Remember that
$\widehat{\boldsymbol{\beta}}$ $=$ $\underset{\boldsymbol{\beta}\in\mathbb{R}^{n+p_{y}}}{\operatornamewithlimits{argmin\ }}\>\mathopen{}\mathclose{{\left\{\>\frac{1}{2T}\>\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left|\mathbf{y}-\mathbf{W}^{\prime}\boldsymbol{\beta}}}\right|}}\right|_{2}^{2}+\lambda||\boldsymbol{\beta}||_{1}\>}}\right\}$
$=$ $\underset{\boldsymbol{\beta}\in\mathbb{R}^{n+p_{y}}}{\operatornamewithlimits{argmin\ }}\>\mathopen{}\mathclose{{\left\{\>\>\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left|\mathbf{y}-\mathbf{W}^{\prime}\boldsymbol{\beta}}}\right|}}\right|_{2}^{2}+2T\lambda||\boldsymbol{\beta}||_{1}\>}}\right\}\ \ \ ,$
Define $V_{T}(\mathbf{a})=\sum_{t=1}^{T}\mathopen{}\mathclose{{\left[(v_{t}-\mathbf{a}^{\prime}\mathbf{w}_{t}/\sqrt{T})^{2}-v_{t}^{2}}}\right]+T\lambda\sum_{i=1}^{n+p_{y}}\mathopen{}\mathclose{{\left[|\beta_{i}^{*}+a_{i}/\sqrt{T}|-|\beta_{i}^{*}|}}\right],$ where, $\mathbf{a}=(a_{1}\dots,a_{n+p_{y}})^{\prime}$. We claim that $V_{T}(\mathbf{a})$ is minimized at $\sqrt{T}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*})$ and
$V_{T}(\mathbf{a})$ $=$ $\sum_{t=1}^{T}(v_{t}-\mathbf{a}^{\prime}\mathbf{w}_{t}/\sqrt{T})^{2}+T\lambda\sum_{i=1}^{n+p_{y}}|\beta_{i}^{*}+a_{i}/\sqrt{T}|-\mathopen{}\mathclose{{\left(\sum_{t=1}^{T}v_{t}^{2}+T\lambda\sum_{i=1}^{n+p_{y}}|\beta_{i}^{*}|}}\right)=$
$=$ $\sum_{t=1}^{T}(y_{t}-\mathbf{w}_{t}^{\prime}\boldsymbol{\beta}^{*}-\mathbf{a}^{\prime}\mathbf{w}_{t}/\sqrt{T})^{2}+T\lambda\sum_{i=1}^{n+p_{y}}|\beta_{i}^{*}+a_{i}/\sqrt{T}|$
$-\mathopen{}\mathclose{{\left(\sum_{t=1}^{T}(y_{t}-\mathbf{w}_{t}^{\prime}\boldsymbol{\beta}^{*})^{2}+T\lambda\sum_{i=1}^{n+p_{y}}|\beta_{i}^{*}|}}\right)$
$=$ $A_{T}(\mathbf{a})-A\ \ ,$
where
$A_{T}(\mathbf{a})=\sum_{t=1}^{T}(y_{t}-\mathbf{w}_{t}^{\prime}\boldsymbol{\beta}^{*}-\mathbf{a}^{\prime}\mathbf{w}_{t}/\sqrt{T})^{2}+T\lambda\sum_{i=1}^{n+p_{y}}|\beta_{i}^{*}+a_{i}/\sqrt{T}|\ \ ,$
and
$A=\sum_{t=1}^{T}(y_{t}-\mathbf{w}_{t}^{\prime}\boldsymbol{\beta}^{*})^{2}+T\lambda\sum_{i=1}^{n+p_{y}}|\beta_{i}^{*}|\ \ .$
Since $A$ does not depend on $\mathbf{a}$, minimizing $V_{T}(\mathbf{a})$ with respect to $\mathbf{a}$ is equivalent to minimizing $A_{T}(\mathbf{a})$ with respect to $\mathbf{a}$. Thus, in order to show that $\sqrt{T}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*})$ is the minimizer of $V_{T}(\mathbf{a})$ it is sufficient to show that it is the minimizer of $A_{T}(\mathbf{a})$.
$A_{T}\mathopen{}\mathclose{{\left(\sqrt{T}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*})}}\right)$ $=$ $\sum_{t=1}^{T}\mathopen{}\mathclose{{\left(y_{t}-(\boldsymbol{\beta}^{*}+\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*})^{\prime}\mathbf{w}_{t}}}\right)^{2}+T\lambda\sum_{i=1}^{n+p_{y}}|\beta_{i}^{*}+\widehat{\beta}_{i}-\beta_{i}^{*}|=$
$=$ $\sum_{t=1}^{T}\mathopen{}\mathclose{{\left(y_{t}-\widehat{\boldsymbol{\beta}}^{\prime}\mathbf{w}_{t}}}\right)^{2}+T\lambda\sum_{i=1}^{n+p_{y}}|\widehat{\beta}_{i}|$
$\leq$ $\sum_{t=1}^{T}\mathopen{}\mathclose{{\left(y_{t}-(\boldsymbol{\beta}^{*}+\mathbf{a}/\sqrt{T})^{\prime}\mathbf{w}_{t}}}\right)^{2}+T\lambda\sum_{i=1}^{n+p_{y}}|\beta_{i}^{*}+\mathbf{a}/\sqrt{T}|(\mathbf{a})$
$=$ $A_{T}\ \ \ ,$
for all $\mathbf{a}$. Note that the inequality (A.3) follows from the definition of $\widehat{\boldsymbol{\beta}}$. Thus, we see that
$\underset{\mathbf{a}\in\mathbb{R}^{n+p_{y}}}{\operatornamewithlimits{argmin\ }}V_{T}(\mathbf{a})=\sqrt{T}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*})\ \ .$
By the Argmin Theorem (geyer1996), we can claim that $\underset{\mathbf{a}\in\mathbb{R}^{n+p_{y}}}{\operatornamewithlimits{argmin\ }}V_{T}(\mathbf{a})\hskip 2.84544pt\stackrel{{\scriptstyle d}}{{\rightarrow}}\hskip 2.84544pt\underset{\mathbf{a}\in\mathbb{R}^{n+p_{y}}}{\operatornamewithlimits{argmin\ }}V(\mathbf{a})$, which implies that $\sqrt{T}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*})\hskip 2.84544pt\stackrel{{\scriptstyle d}}{{\rightarrow}}\hskip 2.84544pt\underset{\mathbf{a}\in\mathbb{R}^{n+p_{y}}}{\operatornamewithlimits{argmin\ }}V(\mathbf{a})$, which would prove the Theorem. In what follows we show that $V_{T}(\mathbf{a})\hskip 2.84544pt\stackrel{{\scriptstyle d}}{{\rightarrow}}\hskip 2.84544ptV(\mathbf{a})$ for all $\mathbf{a}$. Note that
$V_{T}(\mathbf{a})=\sum_{t=1}^{T}\mathopen{}\mathclose{{\left[(v_{t}-\mathbf{a}^{\prime}\mathbf{w}_{t}/\sqrt{T})^{2}-v_{t}^{2}}}\right]+T\lambda\sum_{i=1}^{n+p_{y}}\mathopen{}\mathclose{{\left(|\beta_{i}^{*}+a_{i}/\sqrt{T}|-|\beta_{i}^{*}|}}\right)=I(\mathbf{a})+II(\mathbf{a})\ \ .$
Recall that
$I(\mathbf{a})=\sum_{t=1}^{T}\mathopen{}\mathclose{{\left[(v_{t}-\mathbf{a}^{\prime}\mathbf{w}_{t}/\sqrt{T})^{2}-v_{t}^{2}}}\right]=\mathbf{a}^{\prime}\mathopen{}\mathclose{{\left(\frac{1}{T}\sum_{t=1}^{T}\mathbf{w}_{t}\mathbf{w}_{t}^{\prime}}}\right)\mathbf{a}-\frac{2}{\sqrt{T}}\sum_{t-1}^{T}v_{t}\mathbf{a}^{\prime}\mathbf{w}_{t}\ \ .$
As $T\to\infty$ we have $\mathbf{a}^{\prime}\mathopen{}\mathclose{{\left(\frac{1}{T}\sum_{t=1}^{T}\mathbf{w}_{t}\mathbf{w}_{t}^{\prime}}}\right)\mathbf{a}\to\mathbf{a}^{\prime}\boldsymbol{C}_{w}\mathbf{a}$. Note that $\{v_{t}y_{t-l}\},l\geq 1$, has mean 0, autocovariance function $\gamma(\cdot)$ such that $\sum_{h=-\infty}^{\infty}|\gamma(h)|<\infty$, and autocorrelation coefficient $\phi_{l}$ such that $\sum_{j=0}^{\infty}\phi_{j}\neq 0$. Thus, we can apply the CLT under weak dependence (see billingsley1995, Thm. 27.4) to obtain
$\frac{1}{\sqrt{T}}\sum_{t=1}^{T}v_{t}\mathbf{a}^{\prime}\mathbf{w}_{t}\hskip 2.84544pt\stackrel{{\scriptstyle d}}{{\rightarrow}}\hskip 2.84544ptN\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left(\boldsymbol{0}_{n}^{\prime},\boldsymbol{\mu}_{vy}}}\right)^{\prime},\mathbf{a}^{\prime}\mathopen{}\mathclose{{\left({\begin{array}[]{cc}\sigma_{v}^{2}\mathbf{C}_{u}&\boldsymbol{0}_{n\times p_{y}}\\ \boldsymbol{0}_{p_{y}\times n}&\boldsymbol{\Gamma}_{vy}\\ \end{array}}}}\right)\mathbf{a}}}\right)\ \ .$
Therefore,
$\frac{1}{\sqrt{T}}\sum_{t=1}^{T}v_{t}\mathbf{a}^{\prime}\mathbf{w}_{t}\hskip 2.84544pt\stackrel{{\scriptstyle d}}{{\rightarrow}}\hskip 2.84544pt\mathbf{a}^{\prime}\mathbf{m}\ \ ,$
where, $\mathbf{m}\sim N\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left(\boldsymbol{0}_{n}^{\prime},\boldsymbol{\mu}_{vy}}}\right)^{\prime},^{\prime}\mathopen{}\mathclose{{\left({\begin{array}[]{cc}\sigma_{v}^{2}\mathbf{C}_{u}&\boldsymbol{0}_{n\times p_{y}}\\ \boldsymbol{0}_{p_{y}\times n}&\boldsymbol{\Gamma}_{vy}\\ \end{array}}}}\right)}}\right)\ \ .$
Applying Slutsky’s theorem, we have $I(\mathbf{a})\hskip 2.84544pt\stackrel{{\scriptstyle d}}{{\rightarrow}}\hskip 2.84544pt\mathbf{a}^{\prime}\mathbf{C}_{w}\mathbf{a}-2\mathbf{a}^{\prime}\mathbf{m}.$
Recall $II(\mathbf{a})=T\lambda\sum_{i=1}^{n+p_{y}}\mathopen{}\mathclose{{\left(|\beta_{i}^{*}+a_{i}/\sqrt{T}|-|\beta_{i}^{*}|}}\right)$. When $\beta_{i}^{*}=0$,
$II(\mathbf{a})=\lambda\sqrt{T}\sum_{i=1}^{n+p_{y}}|a_{i}|\overset{T\to\infty}{\to}\lambda_{0}\sum_{i=1}^{n+p_{y}}|a_{i}|\ \ ,$
that is a consequence of the assumption $\lambda\sqrt{T}\to\lambda_{0}\geq 0$. Thus, when $\beta_{i}^{*}\neq 0$, we have to show that $\lambda\sum_{i=1}^{n+p_{y}}a_{i}Sign(\beta_{i}^{*})I(\beta_{i}^{*}\neq 0)$. Observe that
$|\beta_{i}^{*}+a_{i}/\sqrt{T}|-|\beta_{i}^{*}|=\frac{1}{\sqrt{T}}\mathopen{}\mathclose{{\left(|\sqrt{T}\beta_{i}^{*}+a_{i}|-|\sqrt{T}\beta_{i}^{*}|}}\right)=$
$=\frac{1}{\sqrt{T}}\mathopen{}\mathclose{{\left(\sqrt{T}Sign(\beta_{i}^{*})\beta_{i}^{*}+Sign(\beta_{i}^{*})a_{i}-|\sqrt{T}\beta_{i}^{*}|}}\right)=\frac{1}{\sqrt{T}}Sign(\beta_{i}^{*})a_{i}\ \ ,$
where the last equality is due to $Sign(\beta_{i})\beta_{i}=|\beta_{i}|$. Therefore ,
$T\lambda\mathopen{}\mathclose{{\left(|\beta_{i}^{*}+a_{i}/\sqrt{T}|-|\beta_{i}^{*}|}}\right)=\lambda\sqrt{T}Sign(\beta_{i}^{*})a_{i}\overset{T\to\infty}{\to}\lambda_{0}Sign(\beta_{i}^{*})a_{i}\ \ .$
We can now say that $T\lambda\sum_{i=1}^{n+p_{y}}\mathopen{}\mathclose{{\left(|\beta_{i}^{*}+a_{i}/\sqrt{T}|-|\beta_{i}^{*}|}}\right)\to\lambda_{0}\sum_{i=1}^{n+p_{y}}a_{i}Sign(\beta_{i}^{*})I(\beta_{i}^{*}\neq 0)$. Hence ,
$II(\mathbf{a})\to\lambda_{0}\sum_{i=1}^{n+p_{y}}\mathopen{}\mathclose{{\left[a_{i}Sign(\beta_{i}^{*})I(\beta_{i}^{*}\neq 0)+|a_{i}|I(\beta_{i}^{*}=0)}}\right]\ \ .$
Therefore, using Slutsky’s theorem, and by combining the two results, we have
$I(\mathbf{a})+II(\mathbf{a})\hskip 2.84544pt\stackrel{{\scriptstyle d}}{{\rightarrow}}\hskip 2.84544pt\mathbf{a}^{\prime}\mathbf{C}_{w}\mathbf{a}-2\mathbf{a}^{\prime}\mathbf{m}+\lambda_{0}\sum_{i=1}^{n+p_{y}}\mathopen{}\mathclose{{\left[a_{i}Sign(\beta_{i}^{*})I(\beta_{i}^{*}\neq 0)+|a_{i}|I(\beta_{i}^{*}=0)}}\right]\ \ ,$
which shows that $V_{T}(\mathbf{a})\hskip 2.84544pt\stackrel{{\scriptstyle d}}{{\rightarrow}}\hskip 2.84544ptV(\mathbf{a}).$ $\square$
Remark 2:
Under the common AR($p$) restriction (see Remark 2 of the main text), $v_{t}=\omega_{t}$ and $E(v_{t}|y_{t-l-1},y_{t-l-2},\dots)=0$, $\forall l\geq 1$. Thus, if $\lambda\rightarrow 0$ and $T^{\frac{1-c}{2}}\lambda\rightarrow\infty$, $c\in[0,1)$, then Theorem 1 holds with $\underset{\mathbf{a}\in\mathbb{R}^{n+p_{y}}}{\operatornamewithlimits{argmin\ }}(V(\mathbf{a}))=\mathbf{C}_{w}^{-1}\mathbf{m}\sim N(\bf 0_{n+p_{y}},\sigma_{v}^{2}\mathbf{C}_{w})$ (Fu2000, Thm. 2) and Theorem 2 ensures $P\mathopen{}\mathclose{{\left(Sign(\widehat{\boldsymbol{\beta}})=Sign(\boldsymbol{\beta}^{*})}}\right)=1-o(e^{-T^{c}})$ for $c\in[0,1)$ (zhaoyu2006, Thm. 1).
A.4 Proof of Theorem 2
Define two distinct events:
$\mathcal{E}.1_{T}$ $=$ $\mathopen{}\mathclose{{\left\{\mathopen{}\mathclose{{\left|\widehat{\mathbf{c}}_{i}(11)^{-1}_{i}b_{i}(1)}}\right|<\sqrt{T}\mathopen{}\mathclose{{\left(|\beta_{i}^{*}|-\frac{\lambda}{2T}\mathopen{}\mathclose{{\left|\widehat{\mathbf{c}}_{i}(11)^{-1}Sign(\beta_{i}^{*})}}\right|}}\right)}}\right\},$
$i=1,\dots,s,n+1,\dots,n+s_{y}\ \ ,$
$\mathcal{E}.2_{T}$ $=$ $\mathopen{}\mathclose{{\left\{|b_{i}-b_{i}(2)|\leq\frac{\lambda\varphi}{2\sqrt{T}}}}\right\},\ i=s+1,\dots,n,n+s_{y}+1,\dots,n+p_{y}\ \ ,$
where $\widehat{\mathbf{c}}_{i}(11)_{i}$, $b_{i}$, $b_{i}(1)$ and $b_{i}(2)$ are elements of $\widehat{\mathbf{C}}_{w}(11),\ \mathbf{b}=\mathopen{}\mathclose{{\left(\widehat{\mathbf{C}}_{21}(\widehat{\mathbf{C}}_{11})^{-1}\mathbf{W}(1)\mathbf{v}}}\right),\ \mathbf{b}(1)=\frac{1}{\sqrt{T}}\mathbf{W}(1)\mathbf{v}$ and $\mathbf{b}(2)=\sqrt{T}\mathbf{W}(2)\mathbf{v}$, respectively. $\mathcal{E}.1_{T}$ implies that the signs of the relevant predictors are correctly estimated, while $\mathcal{E}.1_{T}$ and $\mathcal{E}.2_{T}$ together imply that the signs of the irrelevant predictors are shrunk to zero. To show $P\mathopen{}\mathclose{{\left(\exists\lambda\geq 0:Sign(\widehat{\boldsymbol{\beta}})=Sign(\boldsymbol{\beta}^{*})}}\right)\rightarrow 1$, it is sufficient to show that $P\mathopen{}\mathclose{{\left(\exists\lambda\geq 0:Sign(\widehat{\boldsymbol{\beta}})=Sign(\boldsymbol{\beta}^{*})}}\right)\geq P\mathopen{}\mathclose{{\left(\mathcal{E}.1_{T}\cap\mathcal{E}.2_{T}}}\right)$ (see Proposition 1 in zhaoyu2006). Using the identity of $1-P\mathopen{}\mathclose{{\left(\mathcal{E}.1_{T}\cap\mathcal{E}.2_{T}}}\right)\leq P\mathopen{}\mathclose{{\left(\mathcal{E}.1_{T}^{c}}}\right)+P\mathopen{}\mathclose{{\left(\mathcal{E}.2_{T}^{c}}}\right)$ we have that
$P\mathopen{}\mathclose{{\left(\mathcal{E}.1_{T}^{c}}}\right)+P\mathopen{}\mathclose{{\left(\mathcal{E}.2_{T}^{c}}}\right)\leq\\ \sum_{i=1}^{s,n+1,\dots,n+s_{y}}P\mathopen{}\mathclose{{\left(\frac{1}{\sqrt{T}}|\widehat{\mathbf{c}}_{i}(11)^{-1}\mathbf{w}_{i}^{\prime}\mathbf{v}|\geq\sqrt{T}\mathopen{}\mathclose{{\left(|\beta_{i}^{*}|-\frac{\lambda}{2T}|\widehat{\mathbf{c}}_{i}(11)^{-1}Sign(\beta_{i}^{*})|}}\right)}}\right)\\ +\sum_{i=1}^{s+1,\dots,n,n+s_{y}+1,\dots,n+p_{y}}P\mathopen{}\mathclose{{\left(\frac{1}{\sqrt{T}}|b_{i}-\mathbf{w}_{i}^{\prime}\mathbf{v}|\geq\frac{\lambda\varphi}{2\sqrt{T}}}}\right)=I_{T}+II_{T}\ \ .$ (8)
Note that by Assumption 3 of the main text, $\widehat{\psi}_{max}^{w}\geq\widehat{\psi}_{min}^{w}\geq 0$, hence
$\frac{\lambda}{2T}\mathopen{}\mathclose{{\left|\widehat{\mathbf{c}}_{i}(11)^{-1}Sign(\beta_{i}^{*})}}\right|\leq\frac{\lambda}{2c_{0}T}||Sign(\beta^{*})||_{2}\leq\sqrt{s+s_{y}}\frac{\lambda}{2c_{0}T},$
for some $c_{0}>0$ (see zhaoyu2006, Thm. 3 and 4). Therefore, by the union bound, Markov’s inequality and the mixingale concentration inequality (see Hansen1991, Lemma 2), we have that
$I_{T}$ $\leq$ $(s+s_{y})P\mathopen{}\mathclose{{\left(\underset{i,j}{\operatornamewithlimits{max\ }}\mathopen{}\mathclose{{\left|\sum_{t=1}^{T}\hat{c}_{ij}(11)^{-1}w_{i,t}v_{t}}}\right|\geq T\mathopen{}\mathclose{{\left(|\beta_{i}^{*}|-\frac{\lambda\sqrt{s+s_{y}}}{2c_{0}T}}}\right)}}\right)$
$\leq$ $\mathopen{}\mathclose{{\left[T\mathopen{}\mathclose{{\left(|\beta_{i}^{*}|-\frac{\lambda\sqrt{s+s_{y}}}{2c_{0}T}}}\right)}}\right]^{-c_{1}}(s+s_{y})E\mathopen{}\mathclose{{\left[\max_{l\leq T}\mathopen{}\mathclose{{\left|\sum_{t=1}^{l}\hat{c}_{ij}(11)^{-1}w_{i,t}v_{t}}}\right|^{c_{1}}}}\right]$
$\leq$ $\mathopen{}\mathclose{{\left[T\mathopen{}\mathclose{{\left(|\beta_{i}^{*}|-\frac{\lambda\sqrt{s+s_{y}}}{2c_{0}T}}}\right)}}\right]^{-c_{1}}(s+s_{y})C_{A}^{c_{1}}\mathopen{}\mathclose{{\left(\sum_{t=1}^{T}d_{t}^{2}}}\right)^{c_{1}/2}$
$\leq$ $C(s+s_{y})T^{c_{1}/2}\mathopen{}\mathclose{{\left[T\mathopen{}\mathclose{{\left(|\beta_{i}^{*}|-\frac{\lambda\sqrt{s+s_{y}}}{2c_{0}T}}}\right)}}\right]^{-c_{1}}$
$=$ $C(s+s_{y})\mathopen{}\mathclose{{\left(\frac{1}{T\mathopen{}\mathclose{{\left(|\beta_{i}^{*}|-\frac{\lambda\sqrt{s+s_{y}}}{2c_{0}T}}}\right)}}}\right)^{c_{1}}\overset{T\to\infty}{\to}0\ \ ,$
where $c_{1}>2$ (see Assumption 2 ($b$) in the main text). Conducting a similar analysis for $II_{T}$, and considering that by assumption $\sqrt{T}\lambda\rightarrow\lambda_{0}\geq 0$, we obtain $P\mathopen{}\mathclose{{\left(Sign(\widehat{\boldsymbol{\beta}})=Sign(\boldsymbol{\beta}^{*})}}\right)\rightarrow 1.$ $\square$
A.5 Proof of Theorem 3
We start by introducing some important definitions.
Definition 1:
Let $\mathopen{}\mathclose{{\left(\Omega,\mathcal{F},P}}\right)$ be a probability space and let $\mathcal{G}$ and $\mathcal{H}$ be sub-$\mathbf{\sigma}$-fields of $\mathcal{F}$. Then
$\alpha(\mathcal{G},\mathcal{H})=\sup_{G\in\mathcal{G},H\in\mathcal{H}}|\Pr(G\cap H)-\Pr(G)\Pr(H)|$
[…]
$\alpha_{m}=\sup_{t}\alpha(\mathcal{F}_{-\infty}^{t},\mathcal{F}_{t+m}^{\infty}).$
Definition 2:
(Mixingale, davidson1994, ch. 16). The sequence of pairs $\mathopen{}\mathclose{{\left\{X_{t},\mathcal{F}}}\right\}_{-\infty}^{+\infty}$ in a filtered probability space $\mathopen{}\mathclose{{\left(\Omega,\mathcal{F},P}}\right)$ where the $X_{t}$ are integrable r.v.s is called $L_{p}$-mixingale if, for $p\geq 1$, there exist sequences of non-negative constants $\mathopen{}\mathclose{{\left\{d_{t}}}\right\}_{-\infty}^{+\infty}$ and $\mathopen{}\mathclose{{\left\{\nu_{m}}}\right\}_{0}^{\infty}$ such that $\nu_{m}\rightarrow 0$ as $m\rightarrow\infty$ and
$||E(X_{t}|\mathcal{F}_{t-m})||_{p}\leq d_{t}\nu_{m}$
$||X_{t}-E(X_{t}|\mathcal{F}_{t+m})||_{p}\leq d_{t}\nu_{m+1},$
hold for all $t$ and $m\geq 0$. Furthermore, we say that $\mathopen{}\mathclose{{\left\{X_{t}}}\right\}$ is $L_{p}$-mixingale of size -$a$ with respect to $\mathcal{F}_{t}$ if $\nu_{m}=O(m^{-a-\epsilon})$ for some $\epsilon>0$.
Definition 3:
(Near-Epoch Dependence, davidson1994, ch. 17). For a possibly vector-valued stochastic sequence $\mathopen{}\mathclose{{\left\{\mathbf{V}_{t}}}\right\}_{-\infty}^{+\infty}$, in a probability space $\mathopen{}\mathclose{{\left(\Omega,\mathcal{F},P}}\right)$ let $\mathcal{F}_{t-m}^{t+m}=\mathbf{\sigma\mathopen{}\mathclose{{\left(\mathbf{V}_{t-m},\dots,\mathbf{V}_{t+m}}}\right)}$, such that $\mathopen{}\mathclose{{\left\{\mathcal{F}_{t-m}^{t+m}}}\right\}_{m=0}^{\infty}$ is a non-decreasing sequence of $\mathbf{\sigma}$-fields. If for $p>0$ a sequence of integrable r.v.s $\mathopen{}\mathclose{{\left\{X_{t}}}\right\}_{-\infty}^{+\infty}$ satisfies
$||X_{t}-E(X_{t}|{F}_{t-m}^{t+m})||_{p}\leq d_{t}\nu_{m},$
where $\nu_{m}\rightarrow 0$ and $\mathopen{}\mathclose{{\left\{d_{t}}}\right\}_{-\infty}^{+\infty}$ is a sequence of positive constants, $X_{t}$ will be said to be near-epoch dependent in $L_{p}$-norm ($L_{p}$-NED) on $\mathopen{}\mathclose{{\left\{\mathbf{V}_{t}}}\right\}_{-\infty}^{+\infty}$. Furthermore, we say that $\mathopen{}\mathclose{{\left\{X_{t}}}\right\}$ is $L_{p}$-NED of size -$a$ on $\mathbf{V}_{t}$ if $\nu_{m}=O(m^{-a-\epsilon})$ for some $\epsilon>0$.
Note that we use the same notation for the constants $d_{t}$ and sequence $\nu_{m}$ as for the near-epoch dependence, since they play the same role in both types of dependence.
To simplify the analysis, we frequently make use of arbitrary positive finite constants $C$, as well as of its sub-indexed version $C_{i}$, whose values may change from line to line throughout the paper, but they are always independent of the time and cross-sectional dimension. Generic sequences converging to zero as $T\rightarrow\infty$ are denoted by $\zeta_{T}$. We say a sequence $\zeta_{T}$ is of size $\phi_{0}$ if $\zeta_{T}=O(T-\phi_{0}-\varepsilon)$ for some $\varepsilon>0$.
Remark 3:
Under Assumption 1 of the main text the process $\{x_{i,t}\}$ is $L_{2b_{1}}$-NED of size a-a, with $a\geq 1$, while the process $\{q_{i,t}\}$ is $L_{2c_{1}}$-NED of size d-d, with $d\geq 1$. By Theorems 17.5 in ch.17 of davidson1994, they are also $L_{b_{1}}$ and $L_{c_{1}}$-Mixingale, respectively. In later theorems, the NED order and sequence size are important for asymptotic rates. Assumption 2 ($b$) of the main text requires $\mathbf{q}_{t}$ to have slightly more moments than $c_{1}$. More moments mean tighter error bounds and weaker tuning parameter conditions, but a high $c_{2}$ imposes stronger model restrictions. Under strong dependence, fewer moments are needed, and the reduction from $c_{2}$ to $c_{1}$ reflects the cost of allowing greater dependence through a smaller mixing rate.
Proof of Theorem 3 Let $\widehat{x}_{i,t}^{(\boldsymbol{\phi},\boldsymbol{\theta})}=\sum_{l=1}^{\widehat{p}_{i}}\widehat{\phi}_{i,l}x_{i,t-l}+\sum_{k=1}^{\widehat{q}_{i}}\widehat{\theta}_{i,k}u_{i,t-k}$, $\widehat{\mathbf{x}}_{t}^{(\boldsymbol{\phi},\boldsymbol{\theta})}=(\widehat{x}_{1,t}^{(\boldsymbol{\phi},\boldsymbol{\theta})},\dots,\widehat{x}_{n_{T},t}^{(\boldsymbol{\phi},\boldsymbol{\theta})})^{\prime}$, and $\mathbf{y}_{t-1}^{(p_{y})}=(y_{t-1},\dots,y_{t-p_{y}})^{\prime}$. Note that $\sum_{t=1}^{T}v_{t}\mathbf{w}_{t}^{\prime}=\sum_{t=1}^{T}v_{t}\mathopen{}\mathclose{{\left(\widehat{\mathbf{u}}_{t}^{\prime},\mathbf{y}_{t-1}^{(p_{y})^{\prime}}}}\right)$. Therefore,
$\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left|\sum_{t=1}^{T}v_{t}\mathbf{w}_{t}^{\prime}}}\right|}}\right|_{\infty}=\max\mathopen{}\mathclose{{\left\{\max_{i\leq n_{T}}\mathopen{}\mathclose{{\left|\sum_{t=1}^{T}v_{t}\widehat{u}_{i,t}}}\right|,\max_{j\leq p_{y}}\mathopen{}\mathclose{{\left|\sum_{t=1}^{T}v_{t}y_{t-j}}}\right|}}\right\}.$ (9)
Consequently to (9) we have that $Pr(\mathcal{A}_{T})=1-Pr(\mathcal{A}_{T}^{c})\geq 1-Pr\mathopen{}\mathclose{{\left(\max_{i\leq n_{T},l\leq T}\mathopen{}\mathclose{{\left|\sum_{t=1}^{l}v_{t}\widehat{u}_{i,t}}}\right|>\frac{T\lambda}{4}}}\right)-Pr\mathopen{}\mathclose{{\left(\max_{j\leq p_{y},l\leq T}\mathopen{}\mathclose{{\left|\sum_{t=1}^{l}v_{t}y_{t-j}}}\right|>\frac{T\lambda}{4}}}\right)=1-Pr(I)-Pr(II)$.
We first bound $Pr(I)$. Note that $\widehat{u}_{i,t}=u_{i,t}+x_{i,t}^{(\boldsymbol{\phi},\boldsymbol{\theta})}-\widehat{x}_{i,t}^{(\boldsymbol{\phi},\boldsymbol{\theta})}$. Thus,
$Pr(I)$ $\leq\sum_{i=1}^{n_{T}}Pr\mathopen{}\mathclose{{\left(\max_{l\leq T}\mathopen{}\mathclose{{\left|\sum_{t=1}^{l}v_{t}\widehat{u}_{i,t}}}\right|>\frac{T\lambda}{4}}}\right)$
$\leq\sum_{i=1}^{n_{T}}Pr\mathopen{}\mathclose{{\left(\max_{l\leq T}\mathopen{}\mathclose{{\left|\sum_{t=1}^{l}v_{t}u_{i,t}}}\right|>\frac{T\lambda}{8}}}\right)+\sum_{i=1}^{n_{T}}Pr\mathopen{}\mathclose{{\left(\max_{l\leq T}\mathopen{}\mathclose{{\left|\sum_{t=1}^{l}v_{t}\mathopen{}\mathclose{{\left(x_{i,t}^{(\boldsymbol{\phi},\boldsymbol{\theta})}-\widehat{x}_{i,t}^{(\boldsymbol{\phi},\boldsymbol{\theta})}}}\right)}}\right|>\frac{T\lambda}{8}}}\right)$ (10)
$=\sum_{i=1}^{n_{T}}Pr(I_{1})+\sum_{i=1}^{n_{T}}Pr(I_{2}).$
We proceed to analyze $I_{1}$. By Assumptions 1, 2 ($b$) and Theorems 17.5, 17.9 and 17.10 in davidson1994, we have that $\{v_{t}u_{i,t}\}$ is an $L_{m}$-mixingale of appropriate size. By the union bound, the Markov’s inequality and the Hansen’s mixingale concentration inequality, it follows that
$Pr\mathopen{}\mathclose{{\left(\max_{i\leq n_{T},l\leq T}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left|\sum_{t=1}^{l}v_{t}u_{i,t}}}\right|}}\right]>\frac{T\lambda}{8}}}\right)\leq\sum_{i=1}^{n_{T}}Pr\mathopen{}\mathclose{{\left(\max_{l\leq T}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left|\sum_{t=1}^{l}v_{t}u_{i,t}}}\right|}}\right]>\frac{T\lambda}{8}}}\right)\leq$
$\mathopen{}\mathclose{{\left(\frac{T\lambda}{8}}}\right)^{-c_{1}}\sum_{i=1}^{n_{T}}E\mathopen{}\mathclose{{\left[\max_{l\leq T}\mathopen{}\mathclose{{\left|\sum_{t=1}^{l}v_{t}u_{i,t}}}\right|^{c_{1}}}}\right]\leq\mathopen{}\mathclose{{\left(\frac{T\lambda}{8}}}\right)^{-c_{1}}\sum_{i=1}^{n_{T}}C_{1}^{c_{1}}\mathopen{}\mathclose{{\left(\sum_{t=1}^{T}d_{t}^{2}}}\right)^{c_{1}/2}\leq$
$Cn_{T}T^{c_{1}/2}\mathopen{}\mathclose{{\left(\frac{T\lambda}{8}}}\right)^{-c_{1}}\ \ .$
For $I_{2}$, note that $v_{t}(x_{i,t}^{(\boldsymbol{\phi},\boldsymbol{\theta})}-\widehat{x}_{i,t}^{(\boldsymbol{\phi},\boldsymbol{\theta})})=v_{t}\mathopen{}\mathclose{{\left(\sum_{l=1}^{{p}_{i}}(\phi_{i,l}-\widehat{\phi}_{i,l})x_{i,t-l}+\sum_{k=1}^{{q}_{i}}(\theta_{i,k}-\widehat{\theta}_{i,k})u_{i,t-k}}}\right)$. We assume uniform ARMA estimation, namely, there exists $c_{0}>0$ such that $Pr(\mathcal{B}_{T}^{c})\rightarrow 0$, where $\mathcal{B}_{T}\coloneq\mathopen{}\mathclose{{\left\{\max_{1\leq i\leq n_{T}}||\boldsymbol{\widehat{\vartheta}}_{i}-\boldsymbol{\vartheta}_{i}||_{1}\leq c_{0}}}\right\}$, where $\boldsymbol{\vartheta}_{i}=(\phi_{1}\dots,\phi_{p_{i}},\theta_{1},\dots,\theta_{q_{i}})^{\prime}$. Therefore,
$Pr(I_{2})$ $\leq\sum_{i=1}^{n_{T}}Pr\mathopen{}\mathclose{{\left(\max_{l\leq T}\mathopen{}\mathclose{{\left|\sum_{t=1}^{l}v_{t}\sum_{l=1}^{{p}_{i}}(\phi_{i,l}-\widehat{\phi}_{i,l})x_{i,t-l}}}\right|>\frac{T\lambda}{16}}}\right)$
$+\sum_{i=1}^{n_{T}}Pr\mathopen{}\mathclose{{\left(\max_{l\leq T}\mathopen{}\mathclose{{\left|\sum_{t=1}^{l}v_{t}\sum_{k=1}^{{q}_{i}}(\theta_{i,k}-\widehat{\theta}_{i,k})u_{i,t-l}}}\right|>\frac{T\lambda}{16}}}\right)$
$\leq\sum_{i=1}^{n_{T}}Pr\mathopen{}\mathclose{{\left(\max_{l\leq T}\mathopen{}\mathclose{{\left|\sum_{t=1}^{l}v_{t}\sum_{l=1}^{{p}_{i}}x_{i,t-l}}}\right|>\frac{T\lambda}{c_{0}16}}}\right)+Pr\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left|\sum_{l=1}^{{p}_{i}}\mathopen{}\mathclose{{\left(\phi_{i,l}-\widehat{\phi}_{i,l}}}\right)}}\right|>c_{0}}}\right)+$
$+\sum_{i=1}^{n_{T}}Pr\mathopen{}\mathclose{{\left(\max_{l\leq T}\mathopen{}\mathclose{{\left|\sum_{t=1}^{l}v_{t}\sum_{k=1}^{{q}_{i}}u_{i,t-l}}}\right|>\frac{T\lambda}{c_{0}16}}}\right)+Pr\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left|\sum_{k=1}^{{q}_{i}}\mathopen{}\mathclose{{\left(\theta_{i,k}-\widehat{\theta}_{i,k}}}\right)}}\right|>c_{0}}}\right)$
$=\sum_{i=1}^{n_{T}}Pr(I_{2}^{\prime})+\sum_{i=1}^{n_{T}}Pr(I_{2}^{\prime\prime})+2Pr(\mathcal{B}_{T}^{c}),\ \ \ \ \ \ \text{for }c_{0}>0.$
Since $\Pr(\mathcal{B}_{T}^{c})=o(1)$, it is sufficient to bound $I_{2}^{\prime}$. Following the same procedure for $I_{1}$, we have
$Pr(I_{2}^{\prime})\leq Cn_{T}T^{c_{1}/2}\mathopen{}\mathclose{{\left(\frac{T\lambda}{c_{0}16}}}\right)^{-c_{1}}.$
Therefore, we have that
$Pr(I_{1})\leq Cn_{T}T^{c_{1}/2}\mathopen{}\mathclose{{\left(\frac{T\lambda}{8}}}\right)^{-c_{1}},\ \ \ \ \ \ \ Pr(I_{2})\leq Cn_{T}T^{c_{1}/2}\mathopen{}\mathclose{{\left(\frac{T\lambda}{c_{0}16}}}\right)^{-c_{1}},$
which implies that
$Pr(I)\leq C_{1}n_{T}T^{c_{1}/2}\mathopen{}\mathclose{{\left(\frac{T\lambda}{8}}}\right)^{-c_{1}},$ (11)
for some large enough constant $c_{0}$.
For $Pr(II)$, we follow the same procedure and obtain
$Pr\mathopen{}\mathclose{{\left(\max_{j\leq p_{y},l\leq T}\mathopen{}\mathclose{{\left|\sum_{t=1}^{l}v_{t}y_{t-j}}}\right|>\frac{T\lambda}{4}}}\right)\leq C2p_{y}T^{c_{1}/2}\mathopen{}\mathclose{{\left(\frac{T\lambda}{4}}}\right)^{-c_{1}}.$ (12)
Combining the results from (11) and (12), we obtain that, for $T$ and $n_{T}$ large enough,
$Pr(\mathcal{A}^{c})\leq C(2n_{T}+p_{y})T^{c_{1}/2}\mathopen{}\mathclose{{\left(\frac{T\lambda}{4}}}\right)^{-c_{1}}.$
This means that $\Pr(\mathcal{A}_{T})\geq 1-C\mathopen{}\mathclose{{\left(2n_{T}+p_{y}}}\right)\mathopen{}\mathclose{{\left(\frac{1}{\sqrt{T}\lambda}}}\right)^{c_{1}}$. We impose that the probability of the complement event is bounded by a sequence $\zeta_{T}\rightarrow 0$. Thus, $(2n_{T}+p_{y})(\lambda\sqrt{T})^{-c_{1}}\leq\zeta_{T}$, from which $\lambda\geq\frac{C(2n_{T}+p_{y})^{1/c_{1}}\zeta_{T}^{-1/c_{1}}}{\sqrt{T}}$. The Theorem follows from choosing $\zeta_{t}=C\mathopen{}\mathclose{{\left(\sqrt{\text{log}(T)}}}\right)^{-1}$, for a large enough constant $C>0$. $\square$
A.6 Proof of Theorem 4
The proof of Theorem 4 follows that of Theorem 1 in ADAMEK2023.
Proof of Theorem 4.
By Lemma 6.1 in Buhlmann2011 we obtain
$\frac{1}{T}\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left|\mathbf{W}^{\prime}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*})}}\right|}}\right|_{2}^{2}\leq\frac{2}{T}\mathbf{W}\mathbf{v}\bigl(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\bigl)+\lambda\Bigl(||\boldsymbol{\beta}^{*}||_{1}-||\widehat{\boldsymbol{\beta}}||_{1}\Bigl).$
Note that the empirical process $\frac{2}{T}\mathbf{W}\mathbf{v}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*})$, i.e., the random part can be easily bounded in terms of the $\ell_{1}$ norm of the parameters, such that,
$\frac{1}{T}\mathopen{}\mathclose{{\left|\mathbf{W}^{\prime}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*})}}\right|\leq\frac{2}{T}\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left|\mathbf{W}\mathbf{v}}}\right|}}\right|_{\infty}||\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}||_{1}.$
The penalty $\lambda$ is chosen such that $T^{-1}\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left|\mathbf{W}\mathbf{v}}}\right|}}\right|_{\infty}\leq\lambda$. Theorem 3, the event $\mathcal{A}_{T}\coloneqq\mathopen{}\mathclose{{\left\{T^{-1}\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left|\mathbf{W}\mathbf{v}}}\right|}}\right|_{\infty}\leq\frac{\lambda_{0}}{2}}}\right\}$ holds with high probability, where $\lambda_{0}\leq\frac{\lambda}{2}$. Since $\lambda\geq 2\lambda_{0}$ under $\mathcal{A}_{T}$ and by Assumption 3 of the main text, we can use the following dual norm inequality (Theorem 6.1 Buhlmann2011)
$\frac{1}{T}\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left|\mathbf{W}^{\prime}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*})}}\right|}}\right|_{2}^{2}+\lambda\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left|\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}}}\right|}}\right|_{1}\leq\frac{4\tilde{s}\lambda}{\gamma_{w}^{2}},$
which leads to
$\frac{1}{T}\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left|\mathbf{W}^{\prime}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*})}}\right|}}\right|_{2}^{2}\leq\frac{4\tilde{s}\lambda^{2}}{\gamma_{w}^{2}},$
$\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left|\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}}}\right|}}\right|_{1}\leq\frac{4\tilde{s}\lambda}{\gamma_{w}^{2}},$
with probability at least $1-\zeta_{t}$. The result of the Theorem follows from choosing $\zeta_{t}=C\mathopen{}\mathclose{{\left(\sqrt{\text{log}(T)}}}\right)^{-1}$, for a large enough constant $C>0$. $\square$
Appendix B Upper Bound for $\psi_{min}$
Here, we would like to point out the role of $\widehat{c}_{ij}^{x}$ for $\widehat{\psi}_{min}^{x}$. To this end, we start by recalling an inequality that links off-diagonal elements and eigenvalues of $\widehat{\mathbf{C}}_{x}$; namely, $\widehat{\psi}_{min}^{x}\leq 1-\underset{i\neq j}{\text{max}}|\widehat{c}_{ij}^{x}|$. Because of this, for any given $\tau\in[0,1)$ we have
$\Pr\mathopen{}\mathclose{{\left(\widehat{\psi}_{min}^{x}\leq 1-\tau}}\right)\geq\Pr\mathopen{}\mathclose{{\left(1-\max_{i\neq j}|\widehat{c}_{ij}^{x}|\leq 1-\tau}}\right)\geq\Pr\mathopen{}\mathclose{{\left(1-|\widehat{c}_{i\neq j}^{x}|\leq 1-\tau}}\right)=\Pr\mathopen{}\mathclose{{\left(|\widehat{c}_{i\neq j}^{x}|\geq\tau}}\right)$
which emphasizes how the probability of a generic sample correlation being large in absolute value affects the probability of the minimum eigenvalue being small – and thus the estimation error bounds of the LASSO, as established by Bickel2009. As the next example shows, point the inequality $\widehat{\psi}_{min}^{x}\leq 1-\underset{i\neq j}{\text{max}}|\widehat{c}_{ij}^{x}|$ can be easily fixed.
Example 1:
Let $e_{i}$ and $e_{j}$ be vectors from the standard basis of $\mathbb{R}^{n}$, $i,j\in{1,\dots,N}$. Moreover, let $x_{\pm}=2^{-1/2}(e_{i}\pm e_{j})$, satisfying $x_{\pm}||_{2}=1$, and let $A$ be a correlation matrix with $a_{k}$ be the $k$-th column. Then we have
$x_{\pm}^{\prime}Ax_{\pm}=\frac{1}{2}(e_{i}\pm e_{j})^{\prime}(a_{i}\pm a_{j})=\frac{1}{2}(a_{ii}\pm 2a_{ij}+a_{jj})=1\pm a_{ij}\ \ .$
Thus, $\psi_{min}\leq 1-|a_{ij}|$ for all $i\neq j$ and so
$\psi_{min}\leq 1-\max_{i\neq j}|a_{ij}|\ \ .$
Acknowledgements
The authors wish to thank Marco Lippi for valuable suggestions on the theoretical developments underlying this work. We are also grateful to Sebastiano Michele Zema, Luca Insolia and Mario Martinoli for helpful comments and stimulating conversations. This work was partially supported by the Huck Institutes of the Life Sciences at Penn State (F.C.), the L’EMbeDS Department of Excellence of the Sant’Anna School of Advanced Studies (F.C. and S.T.), the SMaRT COnSTRUCT project (CUP J53C24001460006, as part of FAIR, PE0000013, CUP B53C22003630006, Italian National Recovery and Resilience Plan funded by NextGenerationEU; F.C. and S.T.) and the Italian Ministry of Education, University and Research, Progetti di Ricerca di Interesse Nazionale, research project 2020-2023, project 2020N9YFFE (A.G.).
Supplement - ARMAr-LASSO: Mitigating the Impact of Predictor Serial Correlation on the LASSO
Appendix A ARMAr-LS: Simulation Experiments
Consider the univariate regression model
$y_{t}=\alpha x_{t-1}+\varepsilon_{t},\ \ \ \ \ \ t=\,\dots,T,$ (S.13)
where $x_{t}=\phi x_{t-1}+u_{t}$, $\varepsilon_{t}=\phi_{\varepsilon}\varepsilon_{t-1}+\omega_{t}$, $u_{t}\sim i.i.d.N(0.\sigma^{2})$, and $\omega_{t}\sim i.i.d(0,\sigma^{2}_{\omega})$. In this section, we provide the estimation and inferential properties of the ARMAr Least Squares (ARMAr-LS) estimator relative to DGP (S.13). In this case, the ARMAr-LS model is
$y_{t}=\alpha u_{t-1}+\phi_{y}y_{t-1}+v_{t},\ \ \ \ \ \ t=\,\dots,T,$ (S.14)
where $v_{t}=(\phi-\phi_{y})x_{t-1}+(\phi_{\varepsilon}-\phi_{y})\varepsilon_{t-1}+\omega_{t}$, and $\phi_{y}=\mathopen{}\mathclose{{\left(\sum_{i=1}^{n}\frac{\phi_{i}\alpha_{i}^{*2}}{1-\phi_{i}^{2}}+\frac{\phi_{\varepsilon}}{1-\phi_{\varepsilon}^{2}}}}\right)/\mathopen{}\mathclose{{\left(\sum_{i=1}^{n}\frac{\alpha_{i}^{*2}}{1-\phi_{i}^{2}}+\frac{1}{1-\phi_{\varepsilon}^{2}}}}\right)$. By Assumption 1, we have $E(u_{t}y_{y-1})=0$, $E(v_{t}|u_{t})=0$, and $E(v_{t}|y_{t-1})=(\phi-\phi_{y})x_{t-1}+(\phi_{\varepsilon}-\phi_{y})\varepsilon_{t-1}\neq 0$ (see Example 2). Consequently, $\widehat{\alpha}=\alpha+\frac{\sum_{t=1}^{T-1}u_{t}v_{t}}{\sum_{t=1}^{T-1}u_{t}^{2}}$ and we have:
  • Unbiasedness. $E(\widehat{\alpha}|\mathbf{u})=\alpha+\frac{\sum_{t=1}^{T-1}u_{t}E(v_{t}|\mathbf{u})}{\sum_{t=1}^{T-1}u_{t}^{2}}=\alpha$;
  • Consistency. $plim(\widehat{\alpha})=\alpha+\frac{E(u_{y}v_{t})}{E(u_{t}^{2})}=\alpha$, consequently to exogeneity of $u_{t}$.
  • Efficiency under common AR(1) restriction. Let $\phi=\phi_{\varepsilon}$, then $v_{t}=\omega_{t}$ and $Var(\widehat{\alpha}|\mathbf{u})=\frac{\sigma_{\omega}^{2}}{\sum_{t=1}^{T-1}u_{t}^{2}}$
We examine, via simulation, the sampling properties of ARMAr-LAS. We explore performances under DGP (S.13) with $T=100,1000$, $\phi=\phi_{\varepsilon}=0.3,0.6,0.9$, and $\sigma^{2}=\sigma_{\omega}^{2}=1$. We compare OLS, OLS augmented with $y_{t-1}$ as an additional regressor (OLSy), and our proposed ARMAr-LS in estimating $\alpha$. For completeness, we also report the estimates obtained from the working model (S.14) (WM). Results are obtained on 1000 Monte Carlo simulations. Figure S.5 illustrates that, as $\phi$ increases, the variance of OLS estimates rises, reflecting their inefficiency under stronger serial correlation, while OLSy becomes increasingly biased. As the sample size $T$ grows, however, all methods exhibit reduced variance. The ARMAr-LASSO estimates outperform those of OLS and OLSy and do not differ significantly from those of the WM, confirming that the estimation of $u$ does not pose any issue.
Figure S.5: Estimates of $\alpha$ on 1000 Monte Carlo simulations for various values of $T$ and $\phi$.
Appendix B Distribution of Sample Correlation Between Serially Correlated Processes: Simulation Experiments
In this Section, we conduct Monte Carlo experiments to assess numerically the approximation of the density of $\widehat{c}_{ij}^{x}$ to $\mathcal{D}(r)$, as described in Section 2.1 of the main text. In particular, we compare the density of $\widehat{c}_{ij}^{x}$ obtained by simulations (indicated as $d(r)$) the distribution provided in Proposition 1 of the main text (indicated as $\mathcal{D}(r)$). After, we expand the theoretical results in more generic contexts, relaxing the assumption that the covariates are orthogonal Gaussian AR(1) processes.
B.1 Numerical Approximation of $d(r)$ to $\mathcal{D}(r)$
We generate data from the bivariate process $\mathbf{x}_{t}=\mathbf{D}_{\phi}\mathbf{x}_{t-1}+\mathbf{u}_{t}$ for $t=1,\dots,T$, where $\mathbf{D}_{\phi}$ is a $2\times 2$ diagonal matrix with the same autocorrelation coefficient $\phi$ in both positions along the diagonal, and $u_{t}\sim N(\boldsymbol{0}_{2},\boldsymbol{I}_{2})$. We consider $T=50,100,250$ and $\phi=0.3,0.6,0.9,0.95$ – thus, the parameter $\ddot{\phi}$ in $\mathcal{D}(r)$, here equal to $\phi^{2}$, takes values 0.09, 0.36, 0.81, 0.90. The first row of Figure S.6 (Plots (a), (b), (c)) shows, for various values of $T$ and $\ddot{\phi}$, the density $d(r)$ generated through 5000 Monte Carlo replications. The second row of Figure S.6 (Plots (d), (e), (f)) shows the corresponding $\mathcal{D}(r)$. These were plotted using 5000 values of the argument starting at -1 and increasing by steps of size 0.0004 until 1. As expected, we observe that the degree of approximation of $d(r)$ to $\mathcal{D}(r)$ improves as $T$ increases and/or $\ddot{\phi}$ decreases. In particular, Plots (a), (d) and (g) in Figure S.6, where $T=50$, show that $\mathcal{D}(r)$ approximates $d(r)$ well for a low-to-intermediate degree of serial correlation ($\ddot{\phi}\leq 0.36$, i.e. $\phi\leq 0.6$). In contrast, in cases with high degree of serial correlation ($\ddot{\phi}\geq 0.81$, i.e. $\phi\geq 0.9$), $\mathcal{D}(r)$ has larger tails compared to $d(r)$; that is, the latter over-estimates the probability of large spurious correlations. However, it is noteworthy that the difference between the two densities is negligible for $T\geq 100$ (Figure S.6, Plots (b), (e) and (h) for $T=100$, and Plots (c), (f) and (i) for $T=250$), also with high degree of serial correlation ($\ddot{\phi}\approx 0.90$, i.e. $\phi=0.95$). These numerical experiments corroborate that the sample cross-correlation between orthogonal Gaussian AR(1) processes is affected by the degree of serial correlation in a way that is well approximated by $\mathcal{D}(r)$. In fact, for a sufficiently large finite $T$, we observe that $\Pr\mathopen{}\mathclose{{\left\{|\widehat{c}_{12}^{x}|\geq\tau}}\right\}$, $\tau>0$, increases with $\ddot{\phi}$ in a similar way for $d(r)$ and $\mathcal{D}(r)$.
Figure S.6: Monte Carlo densities for $\widehat{c}_{12}^{x}$ (top) and asymptotic $\mathcal{D}(r)$ (bottom) for various $T$ and $\phi$.
The Impact of $Sign(\ddot{\phi})$
In Section 2.1 of the main text, we pointed out that the impact of $\ddot{\phi}$ on $\mathcal{D}(r)$ depends on $Sign(\ddot{\phi})$. In particular, when $1<\ddot{\phi}<0$, an increment on $|\ddot{\phi}|$ makes the density of $\widehat{c}_{12}^{x}$ more concentrated around 0. In order to validate this result, we run simulations with $T=100$ and different values for the second element of the diagonal of $\mathbf{D}_{\phi}$; namely, 0.3,0.6,0.9,0.95-0.3,-0.6,-0.9,-0.95. Results are shown in Plots (a) and (b) of Figure S.7. In this case, we see that when $Sign(\phi_{1})\neq Sign(\phi_{2})$ and $|\ddot{\phi}|$ increases, $d(r)$ increases its concentration around 0 in a way that is, again, well approximated by $\mathcal{D}(r)$.
General Case
To generalize our findings to the case of non-Gaussian weakly correlated AR and ARMA processes, we generate covariates according to the following DGPs: ${x}_{1t}=(\phi+0.1){x}_{1t-1}+(\phi+0.1){x}_{1t-2}-0.2{x}_{1t-3}+u_{1t}$, and ${x}_{2t}=\phi x_{2t-1}+\phi x_{2t-2}+{u}_{2t}+0.8{u}_{2t-1}$, where $t=1,\ldots,100$ and $\phi=0.15,0.3,0.45,0.475$. Moreover, we generate $u_{1t}$ and $u_{2t}$ from a bivariate Laplace distribution with means 0, variances $1$, and $c_{12}^{u}=0.2$. In these more general cases, we do not know an approximate theoretical density for $\widehat{c}_{12}^{u}$. Therefore, we rely entirely on simulations to show the effect of serial correlation on $\Pr\mathopen{}\mathclose{{\left\{|\widehat{c}_{12}^{x}|\geq\tau}}\right\}$. Figure S.8 shows $d(r)$ obtained from 5000 Monte Carlo replications for the different values of $\phi$. In short, also in the more general cases where covariates are non-Gaussian, weakly correlated AR(3) and ARMA(2,1) processes, the probability of getting large sample cross-correlations depends on the degree of serial correlation. More simulation results are provided below.
Figure S.7: Monte Carlo densities for $\widehat{c}_{12}^{x}$ (a) and corresponding $\mathcal{D}(r)$ (b), for TT=100 and various (negative) $\ddot{\phi}$’s.
Figure S.8: Densities for $\widehat{c}_{12}^{x}$ in the case of Laplace weakly correlated AR(3) and ARMA(2,1) processes, for $T=100$ and various $\phi$’s.
B.2 More General Cases
We study the density of $\widehat{c}_{12}^{x}$ in three different cases: non-Gaussian processes; weakly and high cross-correlated processes; and ARMA processes with different order. Note that for the first two cases the variables are AR(1) processes with $T=100$ and autocorrelation coefficient $\phi=0.3,0.6,0.9,0.95$. Since we do not have $\mathcal{D}(r)$ for these cases, we rely on the densities obtained on 5000 Monte Carlo replications, i.e. $d(r)$, to show the effect of serial correlation on $\Pr\mathopen{}\mathclose{{\left\{|\widehat{c}_{12}^{x}|\geq\tau}}\right\}$.
The Impact of non-Gaussianity
The theoretical contribution reported in Section 2.1 of the main text requires the Gaussianity of $u_{1}$ and $u_{2}$. With the following simulation experiments we show that the impact of $\ddot{\phi}$ on the density of $\widehat{c}_{12}^{x}$ is relevant also when $u_{1t}$ and $u_{2t}$ are non-Gaussian random variables. To this end, we generate $u_{1t}$ and $u_{2t}$ from the following distributions: Laplace with mean 0 and variance 1 (case (a)); Cauchy with location parameter 0 and scale parameter 1 (case (b)); and from a $t$-student with 1 degree of freedom (case (c)). Figure S.9 reports the results of the simulation experiment. We can state that regardless the distribution of the processes, whenever $Sign(\phi_{1})=Sign(\phi_{2})$, the probability of large values of $\widehat{c}_{12}^{x}$ increases with $\ddot{\phi}$. As a curiosity, this result is more evident for the case of Laplace variables, whereas for Cauchy and $t$-student the effect of $\ddot{\phi}$ declines.
Figure S.9: Simulated density of $\widehat{c}_{12}^{x}$ in the case of non-Gaussian processes, for $T=100$ and various values of $\ddot{\phi}$.
The Impact of Population Cross-Correlation
Since orthogonality is an unrealistic assumption for most economic applications, here we admit population cross-correlation. In Figure S.10 we report $d(r)$ when the processes are weakly cross-correlated with $c_{12}^{u}=0.2$, and when the processes are multicollinear with $c_{12}^{u}=0.8$ (usually we refer to multicollinearity when $c_{12}^{u}\geq 0.7$). We observe that the impact of $\ddot{\phi}$ on $d(r)$ depends on the degree of (population) cross-correlation as follows. In the case of weakly correlated processes, an increase in $\ddot{\phi}$ yields a high probability of observing large sample correlations in absolute value. In the case of multicollinear processes, on the other hand, an increase in $\ddot{\phi}$ leads to a high probability of underestimating the true population cross-correlation.
Figure S.10: $d(r)$ obtained through simulations in the case of $c_{12}^{x}=0.2$ (a) and $c_{12}^{x}=0.8$ (b), for $T=100$ and various values of $\ddot{\phi}$.
[…]
To show the effect of serial correlation on a more general case, we generate $x_{1}$ and $x_{2}$ through the following ARMA processes
${x}_{1t}={\phi}{x}_{1t-1}+\phi{x}_{1t-2}-\phi{x}_{1t-3}{u}_{1t}+0.5u_{1t-1},$
${x}_{2t}={\phi}{x}_{2t-1}+\phi{x}_{2t-2}+{u}_{2t}+0.7{u}_{2t-1}-0.4u_{3t-2},$
where $t=1,\dots,100$ and $u_{i}\sim N(0,1)$. In Figure S.11 we report the density of $\widehat{c}_{12}^{x}$ in the case of $T=100$ and $\phi=0.1,0.2,0.3,0.33$. With no loss of generality we can observe that $d(r)$ gets larger as $\phi$ increases, that is $\Pr\mathopen{}\mathclose{{\left\{|\widehat{c}_{12}^{x}|\geq\tau}}\right\}$ increases with $|\phi|$.
Figure S.11: Densities of $d(r)$ between two uncorrelated ARMA Gaussian processes, for $T=100$ and various values of $\ddot{\phi}$.
Appendix C Comparison with ARDL and GLS Estimators
Two natural points of comparison for our proposal are the AutoRegressive Distributed Lag (ARDL) and the Generalized Least Square (GLS) estimators, which are widely used in the literature to tackle serial correlation.
The ARDL consists of regressing the response on its past realizations – the autoregressive component – as well as on current and past values of the predictors – the distributed lag component (see, e.g., Panopoulou2004). Although this method does mitigate serial correlation, it has the drawback of requiring a very large number of coefficients to be estimated. This issue becomes particularly relevant when the sample size is small. In contrast, our proposal only requires the addition of a few response lags.
The popular Cochrane-Orcutt GLS estimator approximates the serial correlation structure of the error term while retaining consistent coefficient estimation (see, e.g., CocOrc1949). Although this improves statistical efficiency and inference compared to conventional least squares, it does not tackle directly the risk of spurious correlations due to predictors’ serial correlations, as described in Section 2.1 of the main text. In particular, while the GLS filter may reduce predictors’ serial correlations, it does not remove them completely if the AR structure of the error term differs from the AR or ARMA structures of the predictors. The GLS-LASSO (chronopoulos2023) can be summarized in the following steps:
Step 1: Estimation of $\varepsilon_{t}$. The estimates of the error term are obtained as $\widehat{\varepsilon}_{t}=y_{t}-\mathbf{x}^{\prime}_{t}\widetilde{\boldsymbol{\alpha}}$, where $\widetilde{\boldsymbol{\alpha}}$ is the solution to the classical Lasso problem using $\mathbf{X}$ as a design matrix.
Step 2: Estimation of $\phi_{\varepsilon}$. The estimates of the parameters of model (3) of the main text, i.e $\phi_{\varepsilon 1},\dots,\phi_{\varepsilon p_{\varepsilon}}$, is obtained as a solution of the following AR($p_{\varepsilon}$) model $\widehat{\varepsilon}_{t}=\phi_{1}\widehat{\varepsilon}_{t-1}+\dots+\phi_{\varepsilon p_{\varepsilon}}\widehat{\varepsilon}_{t-p_{\varepsilon}},$ where $\widehat{\varepsilon}_{t},\dots,\widehat{\varepsilon}_{t-p_{\varepsilon}}$ are obtained at step 1.
Step 3: GLM-LASSO. The LASSO based on the Cochrane-Orcutt GLS filter is
$\widehat{\boldsymbol{\alpha}}=\underset{\boldsymbol{\alpha}\in\mathbb{R}^{n}}{\operatornamewithlimits{argmin\ }}\>\mathopen{}\mathclose{{\left\{\>\frac{1}{2(T)}\>\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left|\widetilde{\mathbf{y}}-\widetilde{\mathbf{X}}\boldsymbol{\alpha}}}\right|}}\right|_{2}^{2}+\lambda||\boldsymbol{\alpha}||_{1}\>}}\right\},$ (S.15)
where, in scalar representation,
$\widetilde{y}_{t}=y_{t}-\sum_{j=1}^{p_{\varepsilon}}\widehat{\phi}_{\varepsilon j}y_{t-j},\ \ \ \ \ \widetilde{x}_{it}=x_{it}-\sum_{j=1}^{p_{\varepsilon}}\widehat{\phi}_{\varepsilon j}x_{it-j},\ \ \ \ \ t=1,\dots,T,\ \ \ \ \ i=1,\dots,n.$
The loss function in (S.15) corresponds to the $\ell_{1}$-penalized regression considering the estimates of $\phi_{\varepsilon l}$, $l=1,\dots,p_{\varepsilon}$. chronopoulos2023 provide the theoretical properties of this procedure and support them through simulation results. Thus, the working model of GLS-LASSO is¡
$y_{t}-\sum_{j=1}^{p_{\varepsilon}}\phi_{\varepsilon j}y_{t-j}=\sum_{i=1}^{n}\alpha_{i}^{*}\mathopen{}\mathclose{{\left(x_{it}-\sum_{j=1}^{p_{\varepsilon}}\phi_{\varepsilon j}x_{it-j}}}\right)+\omega_{t}.$ (S.16)
Here we compare ARMAr-LASSO and GLM-LASSO in two different cases, namely when the common factor restriction holds and when it does not hold.
Common Factor Restriction. The common factor restriction holds when predictors and error term are generated by the same AR($p$) process (Mizon1995), as in the Example 1 and Remark 2 of the main text. In this case, we can easily observe that the working model of ARMAr-LASSO (see (1) in the main text) and (S.16) estimate the true coefficients $\alpha_{i}^{*}$ by means of the AR($p$) residuals $u_{it}$. To this end, we consider the simplest case where both predictors and error term are AR(1) processes with autoregression coefficient $\phi$. In this case the GLM-filter leads to $\widetilde{x}_{it}=x_{it}-\phi x_{it-1}=u_{it}$.
However, also in this case two main differences emerge between the procedures. First, GLS-LASSO requires one more estimation step compared to ARMAr-LASSO. In step 1 GLS-LASSO estimates $\varepsilon_{t}$ by means of classical LASSO applied directly on time series, which we know to be a non-optimal procedure for the LASSO for the problems listed so far. In particular, without removing residuals serial correlation the variance of $T^{-1}\mathbf{x}_{i}^{\prime}\boldsymbol{\varepsilon}$ depends on both $\phi$ and $\phi_{\varepsilon}$ also after the standardization of $\mathbf{x}_{i}$. In fact, after the standardization of $\mathbf{x}_{i}$, $\widehat{Cov}(x_{it},\varepsilon_{t})\approx N\mathopen{}\mathclose{{\left(0\>,\>\frac{1-\phi^{2}\phi_{\varepsilon}^{2}}{(T-1)(1-\phi_{\varepsilon}^{2})(1-\phi_{1}\phi_{\varepsilon})^{2}}}}\right).$ Therefore estimates of $\varepsilon_{t}$ can be problematic in finite samples. Second, GLS-LASSO has poor forecasting performance compared to ARMAr-LASSO. GLS-LASSO reduces the explained variance of $y_{t}$ compared to ARMAr-LASSO since it does not consider the past of $y_{t}$. This can be mitigated by considering the term $\widehat{\phi}y_{t}$ in the forecasting equation, but $\widehat{\phi}$ obtained at step 2 of GLS-LASSO is affected by estimation issues due to the estimate of $\varepsilon_{t}$ at step 1.
Out of the Common Factor Restriction. Here we consider the case where $\phi\neq\phi_{\varepsilon}$, namely, all predictors have the same autoregressive coefficient, which differs from that of the error term. Without loss of generality, we note that in this case $\widetilde{x}_{it}=x_{it}-\phi_{\varepsilon}x_{it-1}=(\phi-\phi_{\varepsilon})x_{it-1}+u_{it}$ exhibits the following variance
$\frac{(1-2\phi\phi_{\varepsilon}+\phi_{\varepsilon}^{2})\sigma_{u_{i}}^{2}}{1-\phi^{2}},$
which corresponds to the variance of an ARMA(1,1) with AR coefficient $\phi$ and MA coefficient $\phi_{\varepsilon}$. This implies that $\widetilde{x}_{it}\neq u_{it}$ and the probability of spurious correlation between $\widetilde{x}_{it}$ and $\widetilde{x}_{jt}$ increases as $|\phi-\phi_{\varepsilon}|$ increases. Therefore, when the common factor restriction does not hold, under Assumptions 1 and 3 of the main text the non-asymptotic error bounds of GLS-LASSO are greater than those of ARMAr-LASSO since, with high probability, the minimum eigenvalue relative to the covariance matrix $\widetilde{\mathbf{X}}\widetilde{\mathbf{X}}^{\prime}/T$ will be smaller than $\widehat{\psi}_{min}^{\omega}$. This will be numerically validated in Supplement D.3.
Appendix D Simulations
D.1 $\widehat{u}$’s Estimation Error
We generated n variables from an AR(3) process and applied three different filters. In the first case, we fitted an AR(1) (underspecified order) model to each variable and used the corresponding parameter for filtering. In the second and third cases, we repeated the same exercise fitting AR(3) (correct order) and AR(5) (overspecified order) processes, respectively. For each of the three scenarios, we reported $\underset{i\leq i\leq n,1\leq t\leq T}{\operatornamewithlimits{max\ }}|\widehat{u}_{i,t}-u_{i,t}|$ under three regimes: $n<T$ (classical setting), $n>T$ (high-dimensional setting), and $n\ll T$ (asymptotic setting). Results, reported in Figure S.12, are obtained on 1000 Monte Carlo replications. When the applied filter is of order at least as large as that of the true AR process, the estimated residuals converge to the true residuals.
Figure S.12: Maximum AR residuals estimation error for different $n$, $T$ and filter settings.
D.2 Simulation Experiments with More DGPs
The response variable is generated using the model $y_{t}=\sum_{i=1}^{n}\alpha_{i}^{*}x_{i,t-1}+\varepsilon_{t}$, and we consider the following data generating processes (DGPs) for predictors and error terms:
  • (A)
    Common AR(1) Restriction: $x_{i,t}=\phi x_{i,t-1}+u_{i,t},\ \varepsilon_{t}=\phi\varepsilon_{t-1}+\omega_{t}$.
  • (B)
    Common AR(1) Restriction with Common Factor: $x_{i,t}=f_{t}+z_{i,t}$, where $f_{t}=\phi f_{t-1}+\delta_{t},$ $z_{i,t}=\phi z_{i,t-1}+\eta_{i,t},$ $\varepsilon_{t}=\phi\varepsilon_{t-1}+\omega_{t}$.
The shocks are generated as follows: $u_{i,t}\sim i.i.d.\ N(0,1)$ with $\mathopen{}\mathclose{{\left(\mathbf{C}_{u}}}\right){ij}=c_{ij}^{u}=\rho^{|i-j|}$, $\delta_{t},\ \eta_{i,t}\sim i.i.d.\ N(0,1)$ with $\mathopen{}\mathclose{{\left(\mathbf{C}_{\eta}}}\right){ij}=c_{ij}^{\eta}=\rho^{|i-j|}$, and $\omega_{t}\sim i.i.d.\ N(0,\sigma_{\omega}^{2})$. For the DGP in (A) we set $\rho=0.8$, while for the DGP in (B) we set $\rho=0.4$ to generate predictors primarily influenced by the common factor, with weakly correlated AR and/or ARMA idiosyncratic components. Finally, we vary the value of $\sigma_{\omega}^{2}$ to explore different signal-to-noise ratios (SNRs).
We compare our ARMAr-LASSO (ARMAr-LAS) with the LASSO-based benchmarks employed in Section 4.1 of the main text.
For the DGP in (A) we set $\rho=0.8$, while for the DGP in (B) we set $\rho=0.4$ to generate predictors primarily influenced by the common factor, with weakly correlated AR or ARMA idiosyncratic components. Finally, we vary the value of $\sigma_{\omega}^{2}$ to explore different signal-to-noise ratios (SNRs).
We compare our ARMAr-LASSO (ARMAr-LAS) with the standard LASSO applied to the observed time series (LAS), LASSO applied to the observed time series plus lags of $y_{t}$ (LASy), GLS-LASSO as proposed by chronopoulos2023 (GLS-LAS), autoregressive distributed lag LASSO (ARDL-LAS), and FarmSelect as proposed by Fan2020, which employs LASSO on factor model residuals (FaSel). The performance of each method is evaluated based on average results from 1000 independent simulations, focusing on the coefficient estimation error (CoEr) obtained as $\widehat{\boldsymbol{\alpha}}-\boldsymbol{\alpha}||_{2}$, the Root Mean Square Forecast Error (RMSFE), and the percentages of true positives (%TP) and false positives (%FP) in selecting relevant predictors. Simulations have varying numbers of predictors (dimensionality), $n=50,150,300$, and a fixed sample size, $T=150$. In this way we cover low ($n=50$), intermediate ($n=150$), and high ($n=300$) dimensional scenarios. For all methods, the tuning parameter $\lambda$ is selected using the Bayesian Information Criterion (BIC). Finally, regardless of the choice of $n$, $\boldsymbol{\alpha}^{*}$ is always taken to have the first $10$ entries equal to $1$ and all others equal to 0. In this way, as $n$ varies, we also cover different levels of sparsity. In addition to the results presented below, Supplement D provides simulations under other DGPs, simulations with a much larger sample size $T$, and simulations where our ARMAr-LASSO misspecifies the AR model of the predictors.
For DGP (A), we investigate settings with different $\phi$ ($0.3,0.6,0.9,0.95$) and different SNR ($0.5,1,5,10$). For GLS-LAS, we estimate an AR(1) model on $\widehat{\varepsilon}_{t}$ (see Supplement C) and use the resulting autoregressive coefficient to filter both response and predictors. For ARDL-LAS, we consider one lag of the response and one lag of each predictor as additional regressors, bringing the number of terms undergoing selection to $n\times 2+1$. For the working model underlying ARMAr-LAS, the $\widehat{u}$’s are obtained by filtering each series with an AR(1) process, and we consider $p_{y}=1$; that is, we take one lag of $y_{t}$ as additional predictor. Results are presented in Table S.4 for SNR values of 1 and 10 (complete results are provided in Supplement D.2). For each SNR, CoEr and RMSFE (both expressed in relative terms to the values obtained by LAS), as well as %TP and %FP are given for every $n$ and $\phi$ considered (the best CoEr and RMSFE are in bold). Results have ARMAr-LAS as the best performer in terms of CoEr and RMSFE across values of $\phi$, $n$, and SNR, demonstrating superior accuracy in both estimation and forecasting compared to the other LASSO-based methods considered. ARMAr-LAS also shows superior performance in feature selection, with a higher %TP and a lower %FP. These gains are more evident when serial correlations are stronger ($\phi=0.6$ or higher).
Notably, under the common AR(1) restriction, the ARMAr-LAS and GLS-LAS estimators should be equivalent (this is the one case where the GLS-LAS estimator removes the serial correlations of the predictors). Nevertheless, GLS-LAS performs on par with ARMAr-LAS only when serial correlations are low; ARMAr-LAS outperforms GLS-LAS when serial correlations are medium/high, likely because the latter requires the estimation of $\widehat{\varepsilon}_{t}$ (see Supplement C). Also, in some instances, ARDL-LAS exhibits a slightly lower %FP than ARMAr-LAS. However, this metric is calculated on $n\times 2+1$ predictors for the former; in terms of the absolute number of false positives, ARDL-LAS has more than ARMAr-LAS (see Supplement C).
Finally, we note that the superior performance of ARMAr-LAS in DGP (B) indicates its effectiveness in handling factor structures, where multicollinearities are more complex than for simple AR processes (DGP (A)).
Table S.4: DGPs (A) and (B). CoEr, RMSFE (relative to LAS), %TP and %FP for LASSO-based benchmarks and ARMAr-LASSO. For each $n$ and $\phi$ setting the best CoEr and RMSFE are in bold.
(A) (B) n 75 150 300 50 150 300 $\phi$ 0.3 0.6 0.9 0.95 0.3 0.6 0.9 0.95 0.3 0.6 0.9 0.95 0.3 0.6 0.9 0.95 0.3 0.6 0.9 0.95 0.3 0.6 0.9 0.95 SNR 0.5 CoEr LASSOy 0.99 0.81 0.37 0.36 0.99 0.82 0.46 0.46 0.99 0.83 0.53 0.53 0.99 0.83 0.41 0.39 0.99 0.85 0.47 0.46 1.00 0.86 0.53 0.51 GLS-LASSO 0.96 0.82 0.64 0.70 0.97 0.81 0.75 0.79 0.97 0.83 0.82 0.85 0.96 0.79 0.64 0.73 0.96 0.80 0.76 0.79 0.96 0.82 0.82 0.84 ARDL-LAS 0.96 0.79 0.34 0.32 0.97 0.80 0.42 0.42 1.32 0.83 0.50 0.48 0.97 0.81 0.38 0.35 0.99 0.83 0.44 0.42 1.46 0.88 0.50 0.48 FaSel 1.01 0.99 1.04 1.00 1.04 1.01 1.00 1.00 2.38 1.60 1.01 0.99 0.91 0.86 1.01 1.02 0.89 0.85 1.03 1.04 5.04 2.81 1.04 1.04 ARMAr-LAS 0.98 0.82 0.34 0.31 0.99 0.82 0.41 0.41 0.99 0.82 0.47 0.45 0.99 0.82 0.37 0.34 0.99 0.83 0.43 0.41 1.01 0.83 0.48 0.46 RMSFE LASSOy 0.98 0.89 0.77 0.76 0.99 0.88 0.82 0.78 0.98 0.90 0.84 0.81 0.99 0.88 0.76 0.72 1.00 0.89 0.79 0.79 0.99 0.89 0.85 0.82 GLS-LASSO 0.97 0.83 0.84 0.86 0.96 0.83 0.86 0.88 0.95 0.85 0.91 0.91 0.96 0.83 0.82 0.85 0.98 0.84 0.90 0.89 0.97 0.86 0.93 0.91 ARDL-LAS 1.00 0.89 0.77 0.75 1.00 0.89 0.82 0.78 1.05 0.91 0.84 0.82 0.99 0.88 0.72 0.70 1.00 0.89 0.77 0.77 1.05 0.88 0.82 0.80 FaSel 0.96 0.97 0.96 0.96 0.96 0.99 0.97 0.94 0.90 0.92 0.97 0.97 0.97 1.00 0.99 0.99 0.98 1.00 0.97 0.98 1.29 1.07 0.99 0.94 ARMAr-LAS 0.98 0.83 0.69 0.68 0.98 0.84 0.73 0.72 0.98 0.86 0.75 0.74 0.97 0.83 0.67 0.64 0.99 0.83 0.72 0.72 1.01 0.86 0.77 0.77 % TP LASSO 44.80 41.40 47.50 48.60 42.60 38.60 43.70 43.80 40.60 37.10 42.30 40.40 34.50 30.40 42.00 43.50 28.50 23.30 33.00 30.00 20.90 18.70 24.10 23.40 LASSOy 45.40 42.10 32.40 31.90 43.70 39.00 31.50 31.00 41.40 37.80 30.40 29.40 35.30 32.00 21.70 22.50 28.90 25.00 17.70 16.40 21.50 20.00 13.20 13.20 GLS-LASSO 45.70 46.20 43.10 43.80 44.20 43.50 41.40 41.50 41.90 42.10 40.50 39.00 36.20 35.10 32.00 34.80 28.60 27.70 28.00 26.50 21.50 21.20 21.70 20.80 ARDL-LAS 43.20 40.20 33.20 33.30 41.70 37.30 31.80 30.90 41.90 36.20 28.80 27.70 33.90 31.60 25.30 24.60 28.00 24.80 20.30 18.60 28.30 20.80 14.10 14.60 FaSel 11.80 18.00 39.80 42.00 19.30 24.70 39.30 42.10 39.90 35.00 38.90 40.30 2.60 8.10 38.60 40.50 2.70 7.10 30.30 29.70 56.50 38.40 22.90 22.30 ARMAr-LAS 48.10 49.90 49.10 48.90 46.20 47.50 47.80 48.10 42.80 45.10 45.20 44.20 38.30 40.80 40.50 40.80 30.30 33.90 33.80 33.60 22.60 26.20 27.00 27.80 % FP LASSO 1.50 4.90 31.40 34.20 0.70 3.00 17.90 17.60 0.30 1.90 9.90 9.80 6.30 9.30 32.50 34.20 3.80 5.50 18.90 18.30 2.20 3.50 10.40 9.90 LASSOy 1.60 2.80 7.80 10.60 0.80 1.50 5.10 5.90 0.40 0.80 3.00 3.60 6.70 8.00 10.20 12.20 3.90 4.80 6.30 6.70 2.30 2.90 3.80 3.90 GLS-LASSO 1.00 1.30 16.90 22.00 0.50 0.60 11.10 12.10 0.20 0.50 6.90 7.30 5.90 6.20 17.50 22.80 3.40 3.50 12.20 13.00 2.00 2.30 7.40 7.50 ARDL-LAS 0.80 1.40 5.00 6.00 0.50 0.70 2.90 3.50 4.60 0.80 1.60 1.90 3.00 4.40 6.90 7.50 2.30 2.60 3.90 4.10 8.10 2.40 2.20 2.30 FaSel 0.90 3.80 29.20 31.50 0.60 2.70 16.60 17.10 10.70 7.50 9.50 9.60 0.30 3.10 31.80 33.80 0.20 1.80 19.10 19.80 48.30 29.80 11.30 10.80 ARMAr-LAS 1.80 1.80 2.00 2.00 0.80 0.90 0.90 1.00 0.40 0.50 0.50 0.50 7.20 8.20 8.00 8.20 4.10 5.00 5.00 4.90 2.60 3.00 2.90 2.90 1 CoEr LASSOy 0.99 0.85 0.51 0.50 0.99 0.86 0.61 0.59 0.99 0.87 0.66 0.66 0.99 0.87 0.56 0.54 0.99 0.89 0.63 0.62 0.99 0.90 0.69 0.66 GLS-LASSO 0.97 0.81 0.70 0.74 0.97 0.83 0.80 0.83 0.97 0.83 0.86 0.88 0.96 0.81 0.71 0.75 0.96 0.82 0.80 0.82 0.96 0.84 0.85 0.85 ARDL-LAS 0.97 0.83 0.44 0.41 0.98 0.84 0.55 0.54 1.01 0.87 0.61 0.60 0.98 0.85 0.49 0.45 0.98 0.86 0.56 0.55 1.07 0.89 0.63 0.60 FaSel 1.10 1.05 1.04 1.00 1.08 1.04 1.02 1.02 1.93 1.34 1.00 0.98 1.01 0.92 1.03 1.02 0.98 0.90 1.04 1.06 3.86 2.29 1.04 1.05 ARMAr-LAS 0.98 0.82 0.41 0.38 0.98 0.83 0.50 0.49 0.98 0.83 0.55 0.53 0.98 0.81 0.44 0.41 0.99 0.83 0.50 0.50 1.00 0.84 0.57 0.54 RMSFE LASSOy 0.98 0.90 0.81 0.76 0.99 0.92 0.87 0.79 1.00 0.90 0.89 0.85 0.99 0.90 0.81 0.83 0.99 0.93 0.83 0.83 1.00 0.92 0.88 0.85 GLS-LASSO 0.95 0.82 0.84 0.86 0.95 0.85 0.88 0.87 0.96 0.82 0.94 0.92 0.96 0.85 0.82 0.91 0.96 0.87 0.89 0.89 0.95 0.86 0.92 0.90 ARDL-LAS 0.99 0.91 0.79 0.74 1.00 0.92 0.86 0.79 1.01 0.91 0.88 0.83 1.00 0.89 0.76 0.78 0.99 0.91 0.77 0.80 1.01 0.92 0.84 0.81 FaSel 0.96 0.98 0.97 0.95 0.97 0.98 0.96 0.95 0.90 0.95 0.96 0.96 1.00 1.03 0.99 0.99 0.99 1.02 0.98 0.94 1.16 1.05 0.99 0.96 ARMAr-LAS 0.96 0.82 0.67 0.65 0.97 0.85 0.72 0.71 0.98 0.84 0.76 0.73 0.97 0.86 0.68 0.71 0.98 0.85 0.72 0.75 0.99 0.85 0.77 0.76 % TP LASSO 60.10 54.00 55.80 56.30 58.40 52.10 53.50 51.70 56.70 51.10 52.30 50.20 53.90 45.30 49.00 50.80 47.60 39.10 41.10 38.60 40.30 33.70 33.70 32.50 LASSOy 60.50 54.90 46.10 46.10 58.70 52.80 46.00 43.90 56.80 51.90 44.80 42.50 54.20 46.60 35.40 36.40 47.80 40.30 30.50 28.00 40.60 34.50 25.10 24.20 GLS-LASSO 62.10 61.90 54.10 53.40 60.20 58.80 52.10 50.60 58.80 58.60 51.50 49.40 56.30 55.20 43.90 44.50 49.40 48.50 38.80 35.20 41.90 42.30 31.90 30.90 ARDL-LAS 59.10 53.60 49.20 48.00 57.30 51.50 45.70 44.20 55.70 50.30 44.20 41.70 53.20 47.30 39.30 38.50 47.00 41.10 33.40 30.00 40.90 34.70 26.50 26.20 FaSel 19.10 26.20 45.60 49.00 30.70 36.30 48.00 50.60 51.90 47.80 50.20 50.80 7.40 15.40 45.30 47.20 6.50 13.00 38.50 37.30 60.60 44.90 32.30 32.30 ARMAr-LAS 63.40 64.00 63.40 64.00 61.30 60.80 61.90 61.70 59.90 60.00 60.20 59.70 57.80 57.60 58.50 57.40 50.50 51.20 51.30 50.20 43.20 44.80 44.50 45.00 % FP LASSO 1.80 5.80 31.70 33.60 0.90 3.40 17.40 16.80 0.40 2.10 9.80 9.40 8.30 11.10 32.50 34.40 5.20 7.10 18.30 17.60 3.30 4.40 10.20 9.70 LASSOy 1.80 3.60 12.80 15.90 0.90 1.80 7.70 8.10 0.50 1.10 4.40 4.90 8.40 9.80 15.50 17.60 5.30 6.20 9.10 9.30 3.30 3.80 5.40 5.30 GLS-LASSO 1.30 1.70 17.90 22.10 0.60 0.70 11.00 11.70 0.30 0.50 6.80 7.20 8.00 8.50 19.20 22.90 5.00 5.30 12.50 12.50 3.10 3.40 7.50 7.20 ARDL-LAS 0.80 2.00 7.80 8.50 0.40 0.90 4.60 4.80 0.70 0.80 2.50 2.80 3.90 5.70 9.20 9.30 2.60 3.40 5.20 5.30 3.50 2.30 3.00 3.00 FaSel 1.40 5.20 29.40 30.70 0.90 3.30 16.60 16.90 9.30 6.20 9.30 9.20 0.40 4.10 33.10 34.30 0.20 2.30 19.80 19.80 46.40 28.60 11.40 11.00 ARMAr-LAS 2.10 2.30 2.20 2.40 1.00 1.10 1.20 1.20 0.50 0.60 0.60 0.60 9.40 9.60 9.80 9.50 5.80 6.20 6.00 6.10 3.70 3.70 3.80 3.60 5 CoEr LASSOy 1.00 0.95 0.83 0.83 1.00 0.96 0.88 0.87 1.00 0.96 0.91 0.89 1.00 0.96 0.87 0.85 1.00 0.97 0.91 0.89 1.00 0.98 0.93 0.92 GLS-LASSO 0.95 0.81 0.79 0.82 0.95 0.82 0.87 0.88 0.96 0.83 0.92 0.92 0.95 0.81 0.82 0.84 0.95 0.83 0.87 0.86 0.96 0.85 0.91 0.90 ARDL-LAS 1.00 0.93 0.68 0.68 1.00 0.95 0.81 0.81 1.00 0.95 0.85 0.84 0.99 0.93 0.70 0.70 0.99 0.94 0.81 0.81 0.99 0.95 0.85 0.85 FaSel 1.45 1.27 1.06 1.07 1.27 1.14 1.04 1.01 1.42 1.15 0.99 0.98 1.18 1.09 1.05 1.06 1.14 1.06 1.07 1.06 2.08 1.46 1.06 1.05 ARMAr-LAS 0.95 0.80 0.50 0.49 0.96 0.81 0.60 0.58 0.97 0.81 0.64 0.61 0.95 0.78 0.52 0.50 0.96 0.79 0.60 0.58 0.97 0.81 0.65 0.63 RMSFE LASSOy 1.00 0.97 0.93 0.92 1.00 0.97 0.95 0.91 1.00 0.98 0.96 0.95 1.00 0.98 0.94 0.90 1.00 0.97 0.96 0.91 1.00 0.99 0.94 0.95 GLS-LASSO 0.95 0.83 0.86 0.86 0.96 0.84 0.88 0.83 0.94 0.85 0.94 0.90 0.95 0.84 0.88 0.87 0.97 0.87 0.88 0.83 0.96 0.91 0.90 0.87 ARDL-LAS 1.00 0.98 0.94 0.93 1.01 0.98 0.95 0.90 1.01 0.99 0.99 0.96 1.02 0.96 0.86 0.84 1.01 0.97 0.90 0.84 1.01 1.00 0.90 0.89 FaSel 0.99 0.99 0.98 0.93 0.99 0.96 0.95 0.93 0.92 0.95 0.97 0.96 1.05 1.03 0.97 0.96 1.04 1.03 0.97 0.93 1.13 1.09 0.97 0.91 ARMAr-LAS 0.97 0.84 0.70 0.64 0.97 0.84 0.70 0.68 0.98 0.84 0.77 0.74 0.97 0.82 0.71 0.65 0.97 0.86 0.74 0.66 0.96 0.89 0.74 0.69 % TP LASSO 90.10 84.10 78.30 78.10 89.10 82.80 79.60 75.70 88.80 82.10 78.70 75.90 90.20 83.30 76.30 75.80 89.00 79.80 71.30 65.70 84.80 77.00 67.80 63.20 LASSOy 90.10 84.30 76.70 77.00 89.20 83.10 78.20 74.40 88.80 82.60 77.60 74.10 90.10 83.60 74.00 72.90 88.90 79.90 69.70 63.60 84.90 77.10 65.70 60.70 GLS-LASSO 91.70 90.50 79.60 78.50 91.00 89.70 79.90 77.90 90.40 89.00 78.70 76.50 91.90 91.00 77.90 75.60 90.70 87.90 74.10 70.10 87.00 84.70 70.50 66.90 ARDL-LAS 89.20 83.40 76.50 74.40 88.70 82.30 77.30 74.00 88.00 81.80 76.80 73.20 89.70 84.00 73.70 71.10 88.50 80.10 71.00 65.90 84.50 77.00 66.60 62.40 FaSel 40.60 50.40 66.10 71.50 60.90 69.10 75.90 76.60 81.90 78.70 78.30 77.30 66.40 64.70 71.20 70.80 65.40 61.40 69.90 67.90 77.30 70.50 67.20 65.60 ARMAr-LAS 92.10 91.30 90.40 91.00 91.50 90.70 90.50 90.20 90.30 90.50 89.40 89.00 92.60 92.20 90.90 90.30 91.00 90.10 89.00 87.20 87.30 87.20 85.90 84.80 % FP LASSO 2.50 7.30 30.90 32.00 1.30 3.80 15.20 13.70 0.60 2.50 8.50 8.30 11.10 14.00 33.50 34.40 7.00 9.50 16.40 14.50 4.50 6.10 9.50 8.60 LASSOy 2.40 6.20 24.70 27.00 1.20 3.30 12.50 11.30 0.60 2.00 7.40 6.90 10.90 13.40 28.60 29.00 7.00 9.00 14.30 12.30 4.50 5.80 8.30 7.30 GLS-LASSO 2.00 2.40 17.80 21.00 1.00 1.20 9.90 9.80 0.50 0.70 5.90 6.30 10.50 11.20 23.20 24.90 6.70 7.50 12.50 11.90 4.30 4.90 7.80 7.40 ARDL-LAS 1.10 3.10 12.30 11.00 0.60 1.70 7.20 6.60 0.30 1.10 4.20 4.00 5.20 7.50 12.50 11.20 3.40 4.70 7.20 6.50 2.20 3.10 4.20 3.90 FaSel 4.10 9.20 29.40 32.10 2.00 5.20 15.90 15.20 6.70 4.90 8.40 8.70 4.00 10.90 34.80 36.40 1.90 6.20 20.60 19.90 28.20 19.20 12.30 11.60 ARMAr-LAS 2.80 3.00 2.90 3.10 1.50 1.30 1.50 1.50 0.70 0.70 0.70 0.80 11.70 11.90 12.10 11.60 7.70 7.80 7.60 7.70 4.90 4.90 4.90 4.80 10 CoEr LASSOy 1.00 0.97 0.92 0.92 1.00 0.98 0.93 0.92 1.00 0.98 0.95 0.94 1.00 0.98 0.94 0.93 1.00 0.98 0.96 0.95 1.00 0.99 0.97 0.96 GLS-LASSO 0.94 0.78 0.82 0.85 0.94 0.79 0.87 0.87 0.94 0.81 0.92 0.90 0.94 0.79 0.85 0.86 0.95 0.81 0.87 0.85 0.95 0.85 0.90 0.88 ARDL-LAS 1.01 0.96 0.78 0.80 1.01 0.97 0.87 0.88 1.01 0.97 0.90 0.90 0.99 0.95 0.80 0.81 1.00 0.96 0.88 0.89 0.99 0.97 0.91 0.92 FaSel 1.77 1.42 1.12 1.08 1.42 1.20 1.04 1.00 1.36 1.13 1.00 0.96 1.19 1.13 1.07 1.07 1.16 1.09 1.07 1.04 1.52 1.22 1.04 1.02 ARMAr-LAS 0.94 0.77 0.51 0.50 0.95 0.77 0.59 0.57 0.95 0.79 0.63 0.60 0.94 0.75 0.52 0.50 0.95 0.75 0.59 0.56 0.96 0.78 0.63 0.60 RMSFE LASSOy 1.00 0.98 0.96 0.96 1.00 0.98 0.96 0.96 1.00 0.99 0.98 0.94 1.00 0.99 0.94 0.96 1.00 0.99 0.97 0.95 1.00 0.99 0.97 0.97 GLS-LASSO 0.95 0.83 0.87 0.89 0.95 0.83 0.87 0.82 0.96 0.85 0.89 0.86 0.94 0.86 0.87 0.87 0.96 0.88 0.84 0.81 0.96 0.88 0.88 0.85 ARDL-LAS 1.01 0.99 0.99 1.05 1.01 0.99 0.98 0.96 1.01 1.01 0.97 0.97 1.00 0.98 0.91 0.96 1.01 0.99 0.93 0.89 1.00 0.99 0.94 0.93 FaSel 0.98 0.98 0.97 0.96 1.00 0.97 0.96 0.93 0.94 0.97 0.95 0.91 1.04 1.03 0.99 0.98 1.06 1.04 0.93 0.88 1.08 1.05 0.92 0.90 ARMAr-LAS 0.97 0.83 0.70 0.69 0.99 0.83 0.70 0.65 0.98 0.84 0.74 0.69 0.96 0.85 0.68 0.64 0.97 0.86 0.66 0.65 0.96 0.86 0.70 0.69 % TP LASSO 97.20 93.10 88.00 87.70 97.10 93.60 87.30 85.30 96.70 93.00 88.30 84.90 98.20 94.70 88.40 87.20 97.40 93.00 84.40 78.70 96.60 91.80 82.80 77.50 LASSOy 97.20 93.20 87.90 87.60 97.10 93.40 87.00 84.70 96.70 93.20 87.90 84.60 98.20 94.80 87.70 86.50 97.40 93.10 84.10 77.20 96.60 91.70 82.20 76.50 GLS-LASSO 98.00 97.30 89.80 88.10 97.80 97.50 89.20 87.90 97.60 97.30 89.20 87.30 98.70 98.10 90.10 88.50 98.00 97.00 88.80 85.60 97.40 96.00 86.60 83.70 ARDL-LAS 96.80 92.70 85.70 83.10 96.70 93.10 86.70 83.80 96.50 92.70 87.40 83.50 98.10 94.60 86.30 82.30 97.20 93.00 84.50 78.80 96.50 91.40 82.60 77.80 FaSel 47.50 61.00 76.90 80.80 70.60 80.30 85.60 85.60 91.00 90.00 87.80 87.20 91.90 86.80 84.80 84.30 89.30 85.00 84.60 82.70 90.30 84.90 84.40 81.60 ARMAr-LAS 98.20 97.80 97.20 96.60 97.70 97.90 96.90 96.30 97.50 97.60 96.60 95.90 98.80 98.70 97.60 97.10 98.00 97.90 96.90 96.40 97.30 97.20 95.80 95.00 % FP LASSO 2.60 7.10 30.20 31.60 1.30 4.30 13.70 12.30 0.70 2.60 8.00 7.30 11.20 15.00 33.50 34.30 7.60 10.00 15.20 12.90 4.70 6.20 8.80 7.70 LASSOy 2.70 7.00 28.00 29.50 1.30 4.00 12.20 10.70 0.70 2.40 7.50 6.60 11.10 14.40 31.10 31.60 7.50 9.60 14.10 11.70 4.70 6.10 8.20 7.10 GLS-LASSO 2.10 2.70 18.60 21.30 1.00 1.40 8.90 9.30 0.50 0.80 5.80 5.90 11.00 11.40 24.10 26.00 7.30 7.80 12.50 11.90 4.50 5.10 7.70 7.30 ARDL-LAS 1.30 3.20 11.40 10.00 0.60 2.10 7.10 6.20 0.40 1.30 4.30 3.70 5.30 7.30 11.90 10.40 3.70 4.90 7.20 6.30 2.30 3.10 4.30 3.80 FaSel 7.10 12.30 30.70 30.70 2.80 6.20 15.10 14.10 5.30 4.70 8.00 8.20 6.50 14.70 36.10 38.00 3.00 8.90 21.40 20.40 13.50 11.40 12.80 11.90 ARMAr-LAS 2.90 3.20 3.30 3.10 1.40 1.50 1.50 1.60 0.70 0.80 0.80 0.80 12.10 12.00 12.50 12.30 8.10 7.90 7.80 7.70 5.00 5.00 5.00 4.90
D.3 Analysis of the minimum eigenvalues
In this section, we compare the minimum eigenvalues of the design matrix of LAS, GLS-LAS, and ARMAr-LAS in the case of $n=50$ and $SNR=10$. Figure S.13 shows the average of the minimum eigenvalues obtained in the experiments presented in Section D.2. Both LAS and GLS-LAS reduce their minimum eigenvalues as $\phi$ increases. This does not happen for ARMAr-LAS, which maintains the same value regardless of the degree of serial correlation. Figure S.14 shows the same results but for the experiments presented in Section 4.1.2 in the main text. In this case, we compare the minimum eigenvalues for the two DGPs reported as 0 for no common factor ($q=0$) and 1 for common factor ($q=1$). Again, ARMAr-LAS maintains larger minimum eigenvalue with respect to LAS and GLS-LAS. This analysis corroborates the statement of Remark 6 of the main text.
Figure S.13: Minimum eigenvalues for the design matrix of LAS, GLS-LAS, and ARMAr-LAS, for various degrees of serial correlation ($\phi$) under DGPs (A) and (B).
Figure S.14: Minimum eigenvalues for the design matrix of LAS, GLS-LAS, and ARMAr-LAS, for DGP in Section 4.1.2 of the main text. In this case, we compare the minimum eigenvalues for the two DGPs reported as 0 for no common factor ($q=0$) and 1 for common factor ($q=1$).
D.4 Performance in a Large $T$ Regime
Here we compare our ARMAr-LAS with the employed LASSO-based benchmarks in the case of DGP (A) (see Section D.2) with $T=1500$, $n=50$, and SNR=10. This section aims to evaluate the performances of ARMAr-LAS in a large sample size regime. Results in Table S.5 show that ARMAr-LAS performs as GLS-LAS. This result is expected since under DGP (A) these two estimators coincide. Further, both outperform the other LASSO-based methods providing more accurate coefficient estimates and forecasts, as well as a perfect variable selection accuracy.
Table S.5: DGPs (A). CoEr, RMSFE (relative to LAS), %TP and %FP for LASSO-based benchmarks and ARMAr-LASSO, under 4 values of $\phi$ with $T=1500$ and $n=50$.
0.3 0.6 0.9 0.95
CoEr
LASSOy 1.00 1.01 0.95 0.93
GLS-LAS 0.91 0.68 0.22 0.16
ARDL-LAS 1.01 0.83 0.28 0.24
FaSel 13.13 9.86 3.04 2.02
ARMAr-LAS 0.91 0.68 0.21 0.15
RMSFE
LASSOy 1.00 0.99 0.97 0.96
GLS-LAS 0.95 0.81 0.46 0.35
ARDL-LAS 1.00 0.83 0.49 0.39
FaSel 1.01 0.99 0.99 1.00
ARMAr-LAS 0.95 0.82 0.46 0.34
% TP
LASSO 100.00 100.00 100.00 100.00
LASSOy 100.00 100.00 100.00 100.00
GLS-LAS 100.00 100.00 100.00 100.00
ARDL-LAS 100.00 100.00 100.00 100.00
FaSel 65.80 67.90 80.90 86.90
ARMAr-LAS 100.00 100.00 100.00 100.00
% FP
LASSO 0.10 0.10 1.70 2.20
LASSOy 0.10 0.10 1.60 2.20
GLS-LAS 0.00 0.00 0.00 0.10
ARDL-LAS 0.00 1.40 1.40 1.40
FaSel 1.10 1.10 1.80 2.10
ARMAr-LAS 0.10 0.10 0.00 0.00
D.5 Performance with Misspecified Autoregressive Structure
In this section, we compare our ARMAr-LASSO with the LASSO-based benchmarks in the case where the former misspecifies the autoregressive model of predictors. In particular, we generated both predictors and error terms from an AR(2) model with autoregressive coefficients equal to 1.2 and -0.4, but ARMAr-LASSO filters predictors through an AR(1) model. We consider $T=150$, SNR=10, and $n=50,150,300$. Results are reported in Table S.6. Also in this case where the predictors are filtered with a misspecified autoregressive model, ARMAr-LAS outperforms LASSO-based benchmarks. This is because, despite the misspecification, the filter can remove the majority of serial correlation with the proper estimation of a single autoregressive coefficient. This is corroborated by the averages of the minimum eigenvalues of the correlation matrices for LAS, GLS-LAS, and ARMAr-LAS which are 0.00680, 0.01824, and 0.02861, respectively.
Table S.6: CoEr, RMSFE (relative to LAS), %TP and %FP for LASSO-based benchmarks and ARMAr-LASSO, under 3 values of $n$.
50 150 300
CoEr
LASSOy 0.97 0.98 0.99
GLS-LAS 0.59 0.72 0.80
ARDL-LAS 0.76 0.94 0.94
FaSel 1.67 1.33 1.28
ARMAr-LAS 0.43 0.60 0.64
RMSFE
LASSOy 0.99 0.99 0.98
GLS-LAS 0.70 0.70 0.76
ARDL-LAS 0.94 0.98 0.98
FaSel 1.00 0.91 0.80
ARMAr-LAS 0.59 0.61 0.66
% TP
LASSO 99.70 99.70 99.80
LASSOy 99.70 99.70 99.80
GLS-LAS 99.90 100.00 100.00
ARDL-LAS 99.80 99.80 99.80
FaSel 78.90 95.80 98.30
ARMAr-LAS 100.00 100.00 100.00
% FP
LASSO 51.20 12.90 8.40
LASSOy 50.10 12.40 8.10
GLS-LAS 17.70 4.80 4.30
ARDL-LAS 35.50 7.50 4.80
FaSel 54.00 19.30 12.80
ARMAr-LAS 7.00 2.50 1.40
Appendix E List of Time Series in the Euro Area Data
[…]
The acronyms for the sectors refer to:
  • (a)
    ICS: Industry & Construction Survey
  • (b)
    CCI: Consumer Confidence Indicators
  • (c)
    M&IR: Money & Interest Rates
  • (d)
    IP: Industrial Production
  • (e)
    HCPI: Harm. Consumer Price Index
  • (f)
    PPI: Producer Price Index
  • (g)
    TO: Turnover & Retail Sale
  • (h)
    HUR: Harm. Unemployment rate
  • (i)
    SI: Service Svy.
As mentioned in the main text, for the first variable of each group we report in brackets its autocorrelation function to show that predictors are serially correlated.
Table S.7: Euro Area macroeconomic variables from Ale2021
ID Description Area Sector Tcode
1 Ind Svy: Employment Expectations ($\boldsymbol{acf:\ 0.97}$) EA ICS 1
2 Ind Svy: Export Order-Book Levels EA ICS 1
3 Ind Svy: Order-Book Levels EA ICS 1
4 Ind Svy: Mfg - Selling Price Expectations EA ICS 1
5 Ind Svy: Production Expectations EA ICS 1
6 Ind Svy: Production Trend EA ICS 1
7 Ind Svy: Mfg - Stocks Of Finished Products EA ICS 1
8 Constr. Svy: Price Expectations EA ICS 1
9 Ind Svy: Export Order Book Position EA ICS 1
10 Ind Svy: Production Trends In Recent Mth. EA ICS 1
11 Ind Svy: Selling Prc. Expect. Mth. Ahead EA ICS 1
12 Ret. Svy: Employment EA ICS 1
13 Ret. Svy: Orders Placed With Suppliers EA ICS 1
14 Constr. Svy: Synthetic Bus. Indicator FR ICS 1
15 Bus. Svy: Constr. Sector - Capacity Utilisation Rate FR ICS 1
16 Constr. Svy: Activity Expectations FR ICS 1
17 Constr. Svy: Price Expectations FR ICS 1
18 Constr. Svy: Unable To Increase Capacity FR ICS 1
19 Constr. Svy: Workforce Changes FR ICS 1
20 Constr. Svy: Workforce Forecast Changes FR ICS 1
21 Svy: Mfg Output - Order Book & Demand FR ICS 1
22 Svy: Mfg Output - Order Book & Foreign Demand FR ICS 1
23 Svy: Mfg Output - Personal Outlook FR ICS 1
24 Svy: Auto Ind - Order Book & Demand FR ICS 1
25 Svy: Auto Ind - Personal Outlook FR ICS 1
26 Svy: Basic & Fab Met Pdt Ex Mach & Eq - Personal Outlook FR ICS 1
27 Svy: Ele & Elec Eq, Mach Eq - Order Book & Demand FR ICS 1
28 Svy: Ele & Elec Eq, Mach Eq - Order Book & Foreign Demand FR ICS 1
29 Svy: Ele & Elec Eq, Mach Eq - Personal Outlook FR ICS 1
30 Svy: Mfg Output - Price Outlook FR ICS 1
31 Svy: Mfg Of Chemicals & Chemical Pdt - Order Book & Demand FR ICS 1
32 Svy: Mfg Of Chemicals & Chemical Pdt - Personal Outlook FR ICS 1
33 Svy: Mfg Of Food Pr & Beverages - Order Book & Demand FR ICS 1
34 Svy: Mfg Of Food Pr & Beverages - Order Book & Foreign Demand FR ICS 1
35 Svy: Mfg Of Trsp Eq - Finished Goods Inventories FR ICS 1
36 Svy: Mfg Of Trsp Eq - Order Book & Demand FR ICS 1
37 Svy: Mfg Of Trsp Eq - Order Book & Foreign Demand FR ICS 1
38 Svy: Mfg Of Trsp Eq - Personal Outlook FR ICS 1
39 Svy: Oth Mfg, Mach & Eq Rpr & Instal - Ord Book & Demand FR ICS 1
40 Svy: Oth Mfg, Mach & Eq Rpr & Instal - Ord Book & Fgn Demand FR ICS 1
41 Svy: Oth Mfg, Mach & Eq Rpr & Instal - Personal Outlook FR ICS 1
42 Svy: Other Mfg - Order Book & Demand FR ICS 1
43 Svy: Rubber, Plastic & Non Met Pdt - Order Book & Demand FR ICS 1
44 Svy: Rubber, Plastic & Non Met Pdt - Order Book & Fgn Demand FR ICS 1
45 Svy: Rubber, Plastic & Non Met Pdt - Personal Outlook FR ICS 1
46 Svy: Total Ind - Order Book & Demand FR ICS 1
47 Svy: Total Ind - Order Book & Foreign Demand FR ICS 1
48 Svy: Total Ind - Personal Outlook FR ICS 1
49 Svy: Total Ind - Price Outlook FR ICS 1
50 Svy: Wood & Paper, Print & Media - Ord Book & Fgn Demand FR ICS 1
51 Trd. & Ind: Bus Sit DE ICS 1
52 Trd. & Ind: Bus Expect In 6Mo DE ICS 1
53 Trd. & Ind: Bus Sit DE ICS 1
54 Trd. & Ind: Bus Climate DE ICS 1
55 Cnstr Ind: Bus Climate DE ICS 1
56 Mfg: Bus Climate DE ICS 1
57 Mfg: Bus Climate DE ICS 1
58 Mfg Cons Gds: Bus Climate DE ICS 1
59 Mfg (Excl Fbt): Bus Climate DE ICS 1
60 Whsle (Incl Mv): Bus Climate DE ICS 1
61 Mfg: Bus Sit DE ICS 1
62 Mfg: Bus Sit DE ICS 1
63 Mfg (Excl Fbt): Bus Sit DE ICS 1
64 Mfg (Excl Fbt): Bus Sit DE ICS 1
65 Cnstr Ind: Bus Expect In 6Mo DE ICS 1
66 Cnstr Ind: Bus Expect In 6Mo DE ICS 1
67 Mfg: Bus Expect In 6Mo DE ICS 1
68 Mfg: Bus Expect In 6Mo DE ICS 1
69 Mfg Cons Gds: Bus Expect In 6Mo DE ICS 1
70 Mfg (Excl Fbt): Bus Expect In 6Mo DE ICS 1
71 Mfg (Excl Fbt): Bus Expect In 6Mo DE ICS 1
72 Rt (Incl Mv): Bus Expect In 6Mo DE ICS 1
73 Whsle (Incl Mv): Bus Expect In 6Mo DE ICS 1
74 Bus. Conf. Indicator IT ICS 1
75 Order Book Level: Ind ES ICS 1
76 Order Book Level: Foreign - Ind ES ICS 1
77 Order Book Level: Investment Goods ES ICS 1
78 Order Book Level: Int. Goods ES ICS 1
79 Production Level - Ind ES ICS 1
80 Cons. Confidence Indicator ($\boldsymbol{acf:\ 0.98}$) EA CCI 1
81 Cons. Svy: Economic Situation Last 12 Mth. - Emu 11/12 EA CCI 1
82 Cons. Svy: Possible Savings Opinion FR CCI 1
83 Cons. Svy: Future Financial Situation FR CCI 1
84 Svy - Households, Economic Situation Next 12M FR CCI 1
85 Cons. Confidence Indicator - DE DE CCI 1
86 Cons. Confidence Index DE CCI 5
87 Gfk Cons. Climate Svy - Bus. Cycle Expectations DE CCI 1
88 Cons.S Confidence Index DE CCI 5
89 Cons. Confidence Climate (Balance) DE CCI 1
90 Cons. Svy: Economic Climate Index (N.West It) IT CCI 5
91 Cons. Svy: Economic Climate Index (Southern It) IT CCI 5
92 Cons. Svy: General Economic Situation (Balance) IT CCI 1
93 Cons. Svy: Prices In Next 12 Mths. - Lower IT CCI 5
94 Cons. Svy: Unemployment Expectations (Balance) IT CCI 1
95 Cons. Svy: Unemployment Expectations - Approx. Same IT CCI 5
96 Cons. Svy: Unemployment Expectations - Large Increase IT CCI 5
97 Cons. Svy: Unemployment Expectations - Small Increase IT CCI 5
98 Cons. Svy: General Economic Situation (Balance) IT CCI 1
99 Cons. Svy: Household Budget - Deposits To/Withdrawals ES CCI 5
100 Cons. Svy: Household Economy (Cpy) - Much Worse FR CCI 5
101 Cons. Svy: Italian Econ.In Next 12 Mths.- Much Worse FR CCI 5
102 Cons. Svy: Major Purchase Intentions - Balance FR CCI 1
103 Cons. Svy: Major Purchase Intentions - Much Less FR CCI 5
104 Cons. Svy: Households Fin Situation - Balance FR CCI 1
105 Indl. Prod. - Excluding Constr. ($\boldsymbol{acf:\ -0.21}$) EA IP 5
106 Indl. Prod. - Cap. Goods EA IP 5
107 Indl. Prod. - Cons. Non-Durables EA IP 5
108 Indl. Prod. - Cons. Durables EA IP 5
109 Indl. Prod. - Cons. Goods EA IP 5
110 Indl. Prod. FR IP 5
111 Indl. Prod. - Mfg FR IP 5
112 Indl. Prod. - Mfg (2010=100) FR IP 5
113 Indl. Prod. - Manuf. Of Motor Vehicles, Trailers, Semitrailers FR IP 5
114 Indl. Prod. - Int. Goods FR IP 5
115 Indl. Prod. - Indl. Prod. - Constr. FR IP 5
116 Indl. Prod. - Manuf. Of Wood And Paper Products FR IP 5
117 Indl. Prod. - Manuf. Of Computer, Electronic And Optical Prod FR IP 5
118 Indl. Prod. - Manuf. Of Electrical Equipment FR IP 5
119 Indl. Prod. - Manuf. Of Machinery And Equipment FR IP 5
120 Indl. Prod. - Manuf. Of Transport Equipment FR IP 5
121 Indl. Prod. - Other Mfg FR IP 5
122 Indl. Prod. - Manuf. Of Chemicals And Chemical Products FR IP 5
123 Indl. Prod. - Manuf. Of Rubber And Plastics Products FR IP 5
124 Indl. Prod. - Investment Goods IT IP 5
125 Indl. Prod. IT IP 5
126 Indl. Prod. IT IP 5
127 Indl. Prod. - Cons. Goods - Durable IT IP 5
128 Indl. Prod. - Investment Goods IT IP 5
129 Indl. Prod. - Int. Goods IT IP 5
130 Indl. Prod. - Chemical Products & Synthetic Fibres IT IP 5
131 Indl. Prod. - Machines & Mechanical Apparatus IT IP 5
132 Indl. Prod. - Means Of Transport IT IP 5
133 Indl. Prod. - Metal & Metal Products IT IP 5
134 Indl. Prod. - Rubber Items & Plastic Materials IT IP 5
135 Indl. Prod. - Wood & Wood Products IT IP 5
136 Indl. Prod. IT IP 5
137 Indl. Prod. - Computer, Electronic And Optical Products IT IP 5
138 Indl. Prod. - Basic Pharmaceutical Products IT IP 5
139 Indl. Prod. - Constr. DE IP 5
140 Indl. Prod. - Ind Incl Cnstr DE IP 5
141 Indl. Prod. - Mfg DE IP 5
142 Indl. Prod. - Rebased To 1975=100 DE IP 5
143 Indl. Prod. - Chems & Chem Prds DE IP 5
144 Indl. Prod. - Ind Excl Cnstr DE IP 5
145 Indl. Prod. - Ind Excl Energy & Cnstr DE IP 5
146 Indl. Prod. - Mining & Quar DE IP 5
147 Indl. Prod. - Cmptr, Eleccl & Opt Prds, Elecl Eqp DE IP 5
148 Indl. Prod. - Interm Goods DE IP 5
149 Indl. Prod. - Cap. Goods DE IP 5
150 Indl. Prod. - Durable Cons Goods DE IP 5
151 Indl. Prod. - Tex & Wearing Apparel DE IP 5
152 Indl. Prod. - Pulp, Paper&Prds, Pubshg&Print DE IP 5
153 Indl. Prod. - Chem Prds DE IP 5
154 Indl. Prod. - Rub&Plast Prds DE IP 5
155 Indl. Prod. - Basic Mtls DE IP 5
156 Indl. Prod. - Cmptr, Eleccl & Opt Prds, Elecl Eqp DE IP 5
157 Indl. Prod. - Motor Vehicles, Trailers&Semi Trail DE IP 5
158 Indl. Prod. - Tex & Wearing Apparel DE IP 5
159 Indl. Prod. - Paper & Prds, Print, Reprod Of Recrd Media DE IP 5
160 Indl. Prod. - Chems & Chem Prds DE IP 5
161 Indl. Prod. - Basic Mtls, Fab Mtl Prds, Excl Mach&Eqp DE IP 5
162 Indl. Prod. - Repair & Install Of Mach & Eqp DE IP 5
163 Indl. Prod. - Mfg Excl Cnstr & Fbt DE IP 5
164 Indl. Prod. - Mining & Ind Excl Fbt DE IP 5
165 Indl. Prod. - Ind Excl Fbt DE IP 5
166 Indl. Prod. - Interm & Cap. Goods DE IP 5
167 Indl. Prod. - Fab Mtl Prds Excl Mach & Eqp ES IP 5
168 Indl. Prod. ES IP 5
169 Indl. Prod. - Cons. Goods ES IP 5
170 Indl. Prod. - Cap. Goods ES IP 5
171 Indl. Prod. - Int. Goods ES IP 5
172 Indl. Prod. - Energy ES IP 5
173 Indl. Prod. - Cons. Goods, Non-Durables ES IP 5
174 Indl. Prod. - Mining ES IP 5
175 Indl. Prod. - Mfg Ind ES IP 5
176 Indl. Prod. - Other Mining & Quarrying ES IP 5
177 Indl. Prod. - Textile ES IP 5
178 Indl. Prod. - Chemicals & Chemical Products ES IP 5
179 Indl. Prod. - Plastic & Rubber Products ES IP 5
180 Indl. Prod. - Other Non-Metal Mineral Products ES IP 5
181 Indl. Prod. - Metal Processing Ind ES IP 5
182 Indl. Prod. - Metal Products Excl. Machinery ES IP 5
183 Indl. Prod. - Electrical Equipment ES IP 5
184 Indl. Prod. - Automobile ES IP 5
185 Euro Interbank Offered Rate - 3-Month (Mean) ($\boldsymbol{acf:\ 0.67}$) EA M&IR 5
186 Money Supply: Loans To Other Ea Residents Excl. Govt. EA M&IR 5
187 Money Supply: M3 EA M&IR 5
188 Euro Short Term Repo Rate FR M&IR 5
189 Datastream Euro Share Price Index (Mth. Avg.) FR M&IR 1
190 Euribor: 3-Month (Mth. Avg.) FR M&IR 5
191 Mfi Loans To Resident Private Sector FR M&IR 5
192 Money Supply - M1 FR M&IR 5
193 Money Supply - M3 FR M&IR 5
194 Share Price Index - Sbf 250 DE M&IR 1
195 Fibor - 3 Month (Mth.Avg.) DE M&IR 5
196 Money Supply - M3 DE M&IR 5
197 Money Supply - M2 DE M&IR 5
198 Bank Prime Lending Rate / Ecb Marginal Lending Facility DE M&IR 5
199 Dax Share Price Index, Ep IT M&IR 1
200 Interbank Deposit Rate-Average On 3-Months Deposits IT M&IR 5
201 Official Reserve Assets ES M&IR 5
202 Money Supply: M3 - Spanish ES M&IR 5
203 Madrid S.E - General Index ES M&IR 5
204 Hicp - Overall Index ($\boldsymbol{acf:\ -0.54}$) EA HCPI 6
205 Hicp - All-Items Excluding Energy, Index EA HCPI 6
206 Hicp - Food Incl. Alcohol And Tobacco, Index EA HCPI 6
207 Hicp - Processed Food Incl. Alcohol And Tobacco, Index EA HCPI 6
208 Hicp - Unprocessed Food, Index EA HCPI 6
209 Hicp - Goods, Index EA HCPI 6
210 Hicp - Industrial Goods, Index EA HCPI 6
211 Hicp - Industrial Goods Excluding Energy, Index EA HCPI 6
212 Hicp - Services, Index EA HCPI 6
213 Hicp - All-Items Excluding Tobacco, Index EA HCPI 6
214 Hicp - All-Items Excluding Energy And Food, Index EA HCPI 6
215 Hicp - All-Items Excluding Energy And Unprocessed Food, Index EA HCPI 6
216 All-Items Hicp DE HCPI 6
217 All-Items Hicp ES HCPI 6
218 All-Items Hicp FR HCPI 6
219 All-Items Hicp IT HCPI 6
220 Goods (Overall Index Excluding Services) DE HCPI 6
221 Goods (Overall Index Excluding Services) FR HCPI 6
222 Processed Food Including Alcohol And Tobacco DE HCPI 6
223 Processed Food Including Alcohol And Tobacco ES HCPI 6
224 Processed Food Including Alcohol And Tobacco FR HCPI 6
225 Processed Food Including Alcohol And Tobacco IT HCPI 6
226 Unprocessed Food DE HCPI 6
227 Unprocessed Food ES HCPI 6
228 Unprocessed Food FR HCPI 6
229 Unprocessed Food IT HCPI 6
230 Non-Energy Industrial Goods DE HCPI 6
231 Non-Energy Industrial Goods FR HCPI 6
232 Services (Overall Index Excluding Goods) DE HCPI 6
233 Services (Overall Index Excluding Goods) FR HCPI 6
234 Overall Index Excluding Tobacco DE HCPI 6
235 Overall Index Excluding Tobacco FR HCPI 6
236 Overall Index Excluding Energy DE HCPI 6
237 Overall Index Excluding Energy FR HCPI 6
238 Overall Index Excluding Energy And Unprocessed Food DE HCPI 6
239 Overall Index Excluding Energy And Unprocessed Food FR HCPI 6
240 Ppi: Ind Excluding Constr. ($\boldsymbol{acf:\ -0.62}$) & Energy EA PPI 6
241 Ppi: Cap. Goods EA PPI 6
242 Ppi: Non-Durable Cons. Goods EA PPI 6
243 Ppi: Int. Goods EA PPI 6
244 Ppi: Non Dom. - Mining, Mfg & Quarrying EA PPI 6
245 Ppi: Non Dom. Mfg DE PPI 6
246 Ppi: Int. Goods Excluding Energy DE PPI 6
247 Ppi: Cap. Goods DE PPI 6
248 Ppi: Cons. Goods DE PPI 6
249 Ppi: Fuel DE PPI 6
250 Ppi: Indl. Products (Excl. Energy) DE PPI 6
251 Ppi: Machinery DE PPI 6
252 Deflated T/O: Ret. Sale In Non-Spcld Str With Food, Bev & Tob ($\boldsymbol{acf:\ -0.47}$) DE T/O 5
253 Deflated T/O: Oth Ret. Sale In Non-Spcld Str DE T/O 5
254 Deflated T/O: Sale Of Motor Vehicle Pts & Acces DE T/O 5
255 Deflated T/O: Wholesale Of Agl Raw Matls & Live Animals DE T/O 5
256 Deflated T/O: Wholesale Of Household Goods IT T/O 5
257 T/O: Ret. Trd, Exc Of Mv , Motorcyles & Fuel ES T/O 5
258 T/O: Ret. Sale Of Clth & Leath Gds In Spcld Str ES T/O 5
259 T/O: Ret. Sale Of Non-Food Prds (Exc Fuel) ES T/O 5
260 T/O: Ret. Sale Of Info, Househld & Rec Eqp In Spcld Str ES T/O 5
261 Ek Unemployment: All ($\boldsymbol{acf:\ 0.76}$) EA HUR 5
262 Ek Unemployment: Persons Over 25 Years Old EA HUR 5
263 Ek Unemployment: Women Under 25 Years Old EA HUR 5
264 Ek Unemployment: Women Over 25 Years Old EA HUR 5
265 Ek Unemployment: Men Over 25 Years Old EA HUR 5
266 Fr Hur All Persons (All Ages) FR HUR 5
267 Fr Hur Femmes (Ages 15-24) FR HUR 5
268 Fr Hur Femmes (All Ages) FR HUR 5
269 Fr Hur Hommes (Ages 15-24) FR HUR 5
270 Fr Hur Hommes (All Ages) FR HUR 5
271 Fr Hur All Persons (Ages 15-24) FR HUR 5
272 Fr Hurall Persons(Ages 25 And Over) FR HUR 5
273 Fr Hur Females (Ages 25 And Over) FR HUR 5
274 Fr Hur Males (Ages 25 And Over) FR HUR 5
275 Bd Hur All Persons (All Ages) DE HUR 5
276 Bd Hur Femmes (Ages 15-24) DE HUR 5
277 Bd Hur Femmes (All Ages) DE HUR 5
278 Bd Hur Hommes (Ages 15-24) DE HUR 5
279 Bd Hur Hommes (All Ages) DE HUR 5
280 Bd Hur All Persons (Ages 15-24) DE HUR 5
281 Bd Hurall Persons(Ages 25 And Over) DE HUR 5
282 Bd Hur Females (Ages 25 And Over) DE HUR 5
283 Bd Hur Males (Ages 25 And Over) DE HUR 5
284 It Hur All Persons (All Ages) IT HUR 5
285 It Hur Femmes (All Ages) IT HUR 5
286 It Hur Hommes (All Ages) IT HUR 5
287 It Hur All Persons (Ages 15-24) IT HUR 5
288 It Hurall Persons(Ages 25 And Over) IT HUR 5
289 Es Hur All Persons (All Ages) ES HUR 5
290 Es Hur Femmes (Ages 16-24) ES HUR 5
291 Es Hur Femmes (All Ages) ES HUR 5
292 Es Hur Hommes (Ages 16-24) ES HUR 5
293 Es Hur Hommes (All Ages) ES HUR 5
294 Es Hur All Persons (Ages 16-24) ES HUR 5
295 Es Hurall Persons(Ages 25 And Over) ES HUR 5
296 Es Hur Females (Ages 25 And Over) ES HUR 5
297 Es Hur Males (Ages 25 And Over) ES HUR 5
298 De - Service Confidence Indicator ($\boldsymbol{acf:\ 0.96}$) DE SI 1
299 De Services - Buss. Dev. Past 3 Months DE SI 1
300 De Services - Evol. Demand Past 3 Months DE SI 1
301 De Services - Exp. Demand Next 3 Months DE SI 1
302 De Services - Evol. Employ. Past 3 Months DE SI 1
303 Fr - Service Confidence Indicator FR SI 1
304 Fr Services - Buss. Dev. Past 3 Months FR SI 1
305 Fr Services - Evol. Demand Past 3 Months FR SI 1
306 Fr Services - Exp. Demand Next 3 Months FR SI 1
307 Fr Services - Evol. Employ. Past 3 Months FR SI 1
308 Fr Services - Exp. Employ. Next 3 Months FR SI 1
309 Fr Services - Exp. Prices Next 3 Months FR SI 1
Appendix F Five most frequently selected predictors
Table S.8 reports the list of the top 5 predictors in terms of selection frequency across forecasting samples obtained from the empirical application in Section 4.2 of the main text. Regardless of the forecasting horizon $h$, the top predictor for ARMAr-LAS is the Goods Index. The other top predictors, also in the HCPI domain, include EA measurements (e.g., Services Index), or are specific to France and Germany (e.g., All-Items).
Table S.8: Five most frequently selected predictors. Selection percentages are ratios between the times a predictor appears in a forecast and the total number of forecasts (120 for hh=12 and 96 for $h=24$).
Rank Selected Variables
hh=12 $h=24$
Goods, Index Goods, Index
85.8% 85.4%
II° Industrial Goods, Index Services, Index
47.5% 43.8%
III° Services, Index All-Items (De)
40.8% 35.4%
IV° All-Items Excluding Tobacco, Index All-Items Excluding Tobacco, Index
32.5% 32.3%
All-Items (Fr) Industrial Goods, Index
24.2% 30.2%
Appendix G Distribution of $\widehat{Cov}(u_{1},u_{2})$
In Figure S.15 we report the density of $\widehat{Cov}(u_{1},u_{2})$ when $u_{1}$ and $u_{2}$ are standard Normal in the cases of $T=30$ and 100. Red line shows the density of $N\mathopen{}\mathclose{{\left(0,\frac{1}{T-1}}}\right)$. Observations are obtained on 5000 Monte Carlo replications. We observe that the approximation of $\widehat{Cov}(u_{1},u_{2})$ to $N(0,\frac{1}{T-1})$ holds also when $T$ is small (see Figure S.15 (a) relative to TT=30). In particular, for $T=30$, the p-value of the Shapiro test is 0.89, the skewness is 0.031 and the kurtosis is 3.001. For $T=100$, the values for the same statistics are 0.200, -0.016, and 3.146, respectively. This analysis corroborate numerically the results in Glen2004, which show that if $x$ and $y$ are $N\mathopen{}\mathclose{{\left(0,1}}\right)$, then the probability density function of $xy$ is $\frac{K_{0}(|xy|)}{pi}$, where $K_{0}(|xy|)$ is the Bessel function of the second kind.
Figure S.15: Density of $\widehat{Cov}(u_{1},u_{2})$ between two uncorrelated standard Normal variables for $T=30$ (a) and $T=100$ (b).